A Complete Bibliography of Publications in Annals of Probability (1970–1979)

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
      WWW URL: http://www.math.utah.edu/~beebe/

14 March 2019 Version 1.04

Title word cross-reference

(−∞, ∞) [Cox77b]. (V_{λ})^{1} [Tay75b]. 2 [Kun74]. 3 [BG79]. α, 1 < α < 2 [Bas76, Owe73]. d̄ [GNS75, Sch76c, ST75]. mod1 [Dia77]. C[0, 1] [Man76]. d [Eri77, Gor73, Tra79, Yod75]. d_{q} [MR75]. D[0, 1] [DT79]. E(|S_{n}|) [Owe73]. F [Jam76]. f(n) ± cg(n) [LW79]. G/G/∞ [Bri74]. GI/G/∞ [Kap75]. I [Csi75]. K [BR77, HN75, Jof74, Ros76, Sch76a, Sch76c, WN73]. L [Kum73, KS78, Yam78]. L_{1} [Eri74]. L_{p} [Eri73b, Eri73c, HC78a, Mae78]. lim_{λ} ↓ 0V_{λ} [Tay75b]. m [Ber73a, Eri73c, Eri74]. N [Ker74, Mye74, OP73, Tra77, Tra79, Yod75, Dav78]. p [Bru75, Jos75, Jos77, LR76, Mou79]. P (ϕ)_{1} [RS76]. P (R_{i} ≤ b_{i}, 1 ≤ i ≤ m | m, n, F = G^{k}) [Ste74, Ste75]. ϕ [GB77, Num79, Sen74b]. Π [dHR79a]. R [NA76, Twe74a, Twe74b, DS75a, Gut74a, Lai76c, Lai77a, LR79b, Smy73]. R^{d} [Eri76, Ste77b]. R^{k} [GS75]. s [LR79a]. S(−∞, ∞) [Cox79b]. S_{n}/a_{n} [Kla74]. S_{n}/n^{1} [Kla73]. Σ [DV74, Kud74, Bru75, Isa79, LR76, Ros79c, Sim75]. sup |S_{n}/n| [Gut79]. t [Gro76, Ism77]. × [ST75]. U [CW77b, Sen74a, Sen77]. |S_{n}|/a_{n} [Kla74].

-Algebras [Kud74]. -Automorphisms [BR77]. -Convergence [Sch76c]. -Dependent [Ber73a, Eri73c, Eri74]. -Dimensional [BG79, Eri77, Gor73, Ros76, Sch76a, Smy73]. -Distribution [Gro76, Ism77]. -Divergence [Csi75]. -Divisibility [Jam76]. -Field [Ros79c]. -Fields [Isa79, DV74]. -Functions [Jos75, Jos77, Mou79]. -Independent [Jof74]. -Invariant [NA76]. -Lattice [LR76]. -Lattices [Bru75]. -Lower [Sim75]. -Mean [DS75a, Gut74a]. -Means [Bru75, LR76, LR79b]. -Metric [ST75]. -Mixing [GB77, Sen74b]. -Parameter [OP73, Tra77, Tra79, Yod75]. -Predictors [LR79a]. -Quick [Lai76c, Lai77a]. -Recurrent [Num79, Twe74a]. -Space [Tra79, Yod75]. -Statistics [CW77b, Sen74a, Sen77]. -Step [Sch76c, Mye74]. -Subinvariant [Twe74b]. -Theory [Twe74a, Twe74b]. -Transient [Twe74b]. -Variation [dHR79a].

2 [Chu73].

42 [Mar74b].

Abelian [DMW74, FP75]. above [Ber73b]. Absolute [Cam75, Gil77, Kuo73, PS79]. Absolutely [DS79, Mar77a]. Absorbing [Bla76, Sie76b]. Absorption [Eth79]. Accuracy [Rei74]. Acknowledgement [Mar74b]. Acknowledgment [Ste75]. Action [Kie75]. Addendum [Ano74m]. Additive [Bil74c, Dub74, Dub75, Dyn77, Fel77, GS73, Har78, PS76]. Adequate [DS75b]. Adjoint [Dun77]. Admissible [Hud76]. Aftereffects [AT78]. Age [AK76, BdH74, Est75, Fea76, Kin75, KW73, Sav76, Wei77]. Age-Dependent [AK76, Est75, Sav76, Wei77, Kin75]. Algebras [Kud74]. Algorithm [Pho78]. Algorithms [Sal79]. Allocations [Sha79]. Almost [GH76, HGS79, Jof74, Ker77, Lin74c, Mon76, Sen73a, Sen76a, Sen77, WW79, de 74b]. Alternate [Cox79a]. Amarts [Ast78]. Amenable [Kie75]. Amount [BF73]. Analogue [HR79]. Analogues [SvH79]. Analysis [Bri74, Smy75]. Analytic [Dav79a, De 73]. Ann [Mar74b]. Ann. [Ano73a, Ano73b, Ano73c, Ano73d, Ano73e, Ano73f, Ano73g, Ano73h, Ano73i, Ano73j, Ano73k, Ano73l, Ano74a, Ano74b, Ano74c, Ano74d, Ano74e, Ano74f, Ano74g, Ano74h, Ano74i, Ano74j, Ano74k, Ano74l, Ano75a, Ano75b, Ano75c, Ano75d, Ano75e, Ano75f, Ano75g, Ano75h, Ano75i, Ano75j, Ano75k, Ano75l, Ano76a, Ano76b, Ano76c, Ano76d, Ano76e, Ano76f, Ano76g, Ano76h, Ano76i, Ano76j, Ano76k, Ano76l, Ano77a, Ano77b, Ano77c, Ano77d, Ano77e, Ano77f, Ano77g, Ano77h, Ano77i, Ano77j, Ano77k, Ano77l, Ano78a, Ano78b, Ano78c, Ano78d, Ano78e, Ano78f, Ano78g, Ano78h, Ano78i, Ano78j, Ano78k, Ano78l, Ano79a, Ano79b, Ano79c, Ano79d, Ano79e, Ano79f, Ano79g, Ano79h, Ano79i, Ano79j, Ano79k, Ano79l]. Annals [Ano74m]. Annihilating [EN74, Eri73a, Sch78b]. Anscombe [CHL79].

Anticipative [KO73]. Application [Afg75, Ben75, Hwa77, Ker78, KK74, Les75, Lyn79, Nea78, Orn73b, Sav75, Wei75b]. Applications [AK73, Bha77, Chu73, Gab77b, GNS75, Kap77, Ker74, Kha73, Lai76a, Lai76d, LW79, Mic76, Sch77b, Sim74, SS78, Wri79, Mod78]. Approach [Bro78, Hal79, HS76, PS75]. Approximation [BP79, Bha75, Che75a, Che75b, Fre73, Fre74a, Hel77, JP75b, Kap77, Ker77, Kie74, Kos74, Kur75, LR77b, Nor75, Rei74, Rev76, Ser75]. Approximations [BCCR79, LR77a, Sch76c]. Arbitrarily [WPR77]. Arbitrary [Qua77]. Arithmetic [Bil74c]. armed [Rod78]. Arrays [Smy73, Tom76]. Associated [Kan75b, PS77, Ste77b, Tei79, Wic73b]. Asymmetric [KT77, Lig77, Mon78a, PT77, Wic74a]. Asymptotic [DM78, De 77, Eri73b, Eri74, Hol77, Hol79, HT79, JT73, Lai76a, Lin74b, Mar74a, Qui76, Rei76, Ros73, SC74, Sie79, Qui77]. Asymptotics [Gre73]. Athreya [Wil74]. Atomic [Lan76]. Attraction [Bas76, Owe73, SS78, Wic74a]. Attractive [Lig78]. Autocorrelations [Hey74]. Automorphisms [BR77]. Autonomous [Kni79]. Average [EM79, Kan79, MS79]. Averages [BE73, Lai73, Roo78, Tom75a, WPR77]. Axioms [Rob73, Rob74].

B [Wil74]. back [Ano73a, Ano73b, Ano73c, Ano73d, Ano73e, Ano73f, Ano74a, Ano74b, Ano74c, Ano74d, Ano74e, Ano74f, Ano75a, Ano75b, Ano75c, Ano75d, Ano75e, Ano75f, Ano76a, Ano76b, Ano76c, Ano76d, Ano76e, Ano76f, Ano77a, Ano77b, Ano77c, Ano77d, Ano77e, Ano77f, Ano78a, Ano78b, Ano78c, Ano78d, Ano78e, Ano78f, Ano79a, Ano79b, Ano79c, Ano79d, Ano79e, Ano79f]. Balayage [KP79]. Banach [AM78, BW74, BR79, Gar76, HJP76, Jai76, Kue75, Kue76a, Kue76b, Kum73, Kuo73, Pho78, SW76, Zin77]. Bandit [Rod78]. Barely [KT77]. Barrier [Bla76, Sie76b]. Barriers [Ber73b]. Based [Leh77, Sin75]. Basic [Gri79]. Baxter [Kaw79]. be [Dym78, Mon78b]. Become [Rub77]. Behavior [CL73, De 77, Fri79, Kue75, MC73, MI73, Mal74b, Mal74a, Mit74, Wol73, Wol78]. Behaviour [DM78, Wes79]. Bernoulli [Mar74c, ST75]. Berry [CW77b, Eri73c, KK74]. Bessel [HW74, Ism77, Ken78]. Bets [GS77]. Between [Bla76, Sus78, LR79b, Ver79]. Big [CR79]. Billiards [BKM78]. binary [Mar74b]. Binomial [Che74, Slu77]. Biological [Mod78]. Birth [JP77, Kin75, Tor78]. Births [Sch76b]. Blackwell [BS77a, Lin77]. Blum [Sen75]. Bonferroni [Gal77]. Borel [Bla78, Che78, Fre73]. Borel-Programmable [Bla78]. Both [VP76]. Bound [Dal76, Hey75, HC78a, Jos75, Jos77, Wri73]. Boundaries [Lai77b, Por78, Woo76]. Boundary [Cha77, Lai74b, Lai76b, Wic73b]. Bounded [Gru75]. Bounds [Dal78, Eri74, Kan79, LW79, Mal73, Mic76, Por78, SW78, Ste79]. Branching [Asm76, AN74, AK76, CNW73, Eri73a, Est75, Fea76, Gri74, Hop76, Kap74a, Kap74b, KN73, Kin75, KW73, Lin74c, Qui76, Qui77, Sav75, Tan77, Tur73a, Tur73b, Wei77, Mod78, Wil74]. Bridge [Ver79]. Brillinger [CMSB75].

Brownian [Adl80, Adl78a, Dav78, Dav79a, Dro73, DI77, DIM77, Fra73, Get79a, JT73, KP79, Kha77, Kue75, Lan78, Mon78b, Qua77, Ste77c, Tay75a, Tra79, Ver79, Wal74a, Wan77, Wan79, Wen77, Yod75, de 74b]. Building [Let75]. Burkholder [Pra75]. By-product [Fre73].

Can [Rub77, Mon78b]. Cancellation [Jon76]. Cantelli [Che78, Fre73, Sin75]. Capacity [Ahl74]. Carrying [DH79]. Case [KN79, Sim74, Sim75, Ste78b]. Cauchy [Mon78a, PT77]. Cells [SC74]. Central [Mic76, Mye76, Nor74, Afg75, Bas74, Ber73a, Bha77, Bol76, Cuz76, Deo75, Dud78, Dud79, Gar76, Gor73, Hah77, HC78a, HJP76, HT79, Igl74, M.74, Mae78, Mat75, McL74, Sen76c, Swe77, Won79, Zin77]. Century [Cra76]. Certain [Dav76, Dav77, Hos78a, KMPQ74, Man76, Res74, Sen76a, Tom76]. Chain [Cha74a, Cha77, Cog75, DO75, Kes74, NA76, O’B74b, Tor78]. Chain-Dependent [DO75]. Chains [AMN78, Cha76a, CI78, Cox77a, Gri76, JP77, Kai75b, KS75, Kus74, NT78, Num79, Sil74, Twe74a, Twe74b, CI80]. Change [Kan75c, VW74]. Changed [Gzy77]. Changes [Lyn79]. Channel [Ahl74]. Characteristic [Ben75, Mal74b, Mal74a, Wol73, Wol75, Wol78]. Characterization [EG73, KS78, PR74, Tay75b]. Characterizations [Wan75]. Chebyshev [Eri73b]. Chernoff [PS75]. Chung [Cin76, Cin77]. Circle [Sie79]. Class [Boo75, Dym78, Hal75, Har76, Kos74, Kum73, KS78, LS78, Muc73, Ros73, Sch77b, Teu75, Wei75b, Yam78]. Classes [Fre74c, KP78]. Classical [Tei73]. Closed [ST75]. Closure [BS76]. CLT [Jai76]. CLTs [PSW75]. Clustering [Hwa77]. Coding [GNO75, PG77]. Coefficient [Bar75]. Coin [DS78]. Coincidence [HN75, Hwa77]. Collision [Gis73]. Combinations [Hel77]. Combinatorial [HC78a]. Compacitè [BR74]. Compact [AV75, DMW74, Lin74a]. Compactification [Glo79]. Compactness [BR74, Man76, dHR79b]. Comparative [FG76, Fis75, KF77]. Compared [Cin76]. Comparison [Mar73a, O’B75, Pro79]. Complements [Ram79a]. Complete [Nai77]. Completely [Wic74a]. Component [Ros79d]. Components [Hen73, LP79, Lin73, NR75]. Computing [Ros73]. Concept [HL77]. Concernant [Pra75]. Concerning [Bas76, Cox79a, GS75, Jai76, Pra75]. Condenser [CG77]. Condition [Mar74c, Mit79, Ros79a, Ros79b]. Conditional [Bil74a, BD75, Bru75, Che78, Dub75, Isa79, Kai75a, LR76, Pfa74, Pfa79, Ram79a, Rog74, Tom75b, Zab79, dJ79]. Conditioned [Bol76, Dur78, Igl74, Kai76, Kur75, LR77a]. Conditioning [JP77]. Conditions [Cuz75, Dav76, Dym78, Hah77, Kal73, Mar73a, Mat77, VP76, Kal76]. Confidence [Lai76b]. Conglomerability [Dub75]. Conjecture [Lai76c, Van78b]. Conjectures [WR78]. Conjugate [dHR79a]. Consolidation [Saw77]. Constant [PR74, Ste77c]. Construction [Hal75, NA76]. Constructive [Cha74c, Cha74b, Cha76b]. Contact [Gri79, Har74, HL78]. Contents [Ano74m]. Contiguity [HL77]. Continous [Zab79]. Continuity [Cam75, Hah77, HK77, Kal73, Kal76, Kuo73, Mar73a, Mar73b, WZ77b]. Continuous [Cha74a, Cha77, DS79, Hal75, Imh76, Kni75, M.74, Mar74a, Mar77a, Mar73c, PR73, PG77, Win73, Win78]. Continuous-time [PG77]. Continuum [Gis73, Ste73]. Contractions [Dun77]. Contributions [KH73]. Controlled [GS75, Gra78, KS75, Put74]. Converge [Roo76, Ste77a]. Convergence [Adl78b, Afg75, Asm76, AK76, BS77a, Bol78, Boo75, Bro78, Cha74c, Che74, CHL79, Deo73c, DS75a, Dro73, Dun77, DIM77, Eag75, EF75, Fra73, Gab77b, GH76, GPW73, Gri74, Gut74a, Gut78, Hey75, Ker78, KN79, Kha73, Kud74, Kue76b, Kum73, Kus74, Lai76c, Lai77a, LR76, Les75, Lig74, LdMR75, Lin74c, Mac73, MR75, Mon75, Mon76, Nai77, O’R74a, O’R74b, Pro77, PS75, Res75, Res74, Rob76, RR73, Sch76c, Sen74a, Sen74b, Sen77, Ser73, Sie76a, Ste78b, Stu77, Swe77, SW76, Wei76, Wel73, Wic73a, de 74b]. Converse [Lyn78]. Conversion [Cin77]. Converting [LPQ78]. Convex [Gar73, Kie79, Lyn79, Mat75, Pit77]. Convexity [Pfa74, Rin76]. Convolutions [Che75a, Sie76a]. Corput [dJS79]. Correction [Adl80, Ber80, Bha80, CI80, Cuz87, Dar76, Dud73a, Kal76, Kan75a, Kar79, Kie76a, PS75, Qui77, Rob74, Rob83, Ros79b, She77, Wei77]. Corrections [Dud79, Rob83]. Correlation [Har77, Pit77]. Corresponding [Kuo73]. Cotype [AM78]. Countable [FP75, Gab77a, Kes76]. Counterexample [Dud73a, Kie73, Kie76a, Kue76a, LR77b, Mon76]. Coupling [Gri76, Lig76]. Coverage [GN79, Sie79]. Cox [GS75]. Cramér [Eri73b, Ivk75]. Critère [BR74]. Criteria [Bha78, Bha80]. Criterion [BR74, GNO75, Loy76, RR73]. Critical [AN74, Est75]. Cross [Bri74]. Cross-Spectral [Bri74]. Crossing [Lai74b, Lai76b, Ven79, Wic73b]. Crossings [AH76, Cuz75, Lin74b, LdMR75, Mar77a]. Curious [Lyn78]. Curved [Woo76].

D [Wic74a]. D. [Smy75]. Darling [Sho79]. d’Arrêt [Bis79]. Death [JP77, Tor78]. Deaths [Sch76b]. Decomposability [SvH79]. Decomposition [Del77, Gru75, Mil78, SG75, Sza77b]. Decompositions [Eis76]. Deduction [Ivk75]. Defined [O’B74b]. Degeneracy [CNW73]. Degrees [Gro76]. Delayed [Lai74a]. Densities [JM78, Kan75c, Kan76]. Density [Bas74, JM78, Ros76]. Density-Preserving [JM78]. Denumerable [O’B74b]. Dependence [Fea76]. Dependent [Adl78b, AK76, Ber73a, BP79, Che75b, DO75, DR78, Eri73c, Eri74, Est75, Fea76, Fre74a, Isa79, Kap77, KW73, LPQ78, LP79, Lin74b, McL74, McL75, Mey73, Phi77, Sav76, Vil74, Wei77, Kin75, Wan79]. Derivatives [Pfa79, SK75, She77, Wol75]. Derived [LM73]. Detection [Lai73]. Determinism [Ber74a]. Deterministic [Jof74, Sus78]. Deux [Pra75]. Deviation [Boo75, GOR79, Lyn79, Rob77, Rob83, Ros76, Ste78b]. Deviations [BZ79, GB77, KD79, Kos73, McD79, Mic74, Mic76, Nag79]. Dichotomous [CP76]. Difference [Lyn79, Tom75a]. Different [LR79b]. Differential [McK73, Sus78]. Diffusion [DH79, Kni73, Kuo73, Kus74, Leh77, Nor75, Por75]. Diffusions [Bha78, Cox77b, Cox79b, Put74, Bha80]. Diffussion [Ros79c]. digits [Dia77, Mar74b]. Dimension [Adl77, Adl78a, Adl80, DH79, Hen73, Tra77]. Dimensional [AK76, BG79, Cox77b, Dav78, Eri77, Ete76, Fla76, Gab77b, Gor73, Kes76, Kiu75, Put74, Ros76, Ros79c, Sch76a, Sen76c, Smy73, Cox79b, She74]. Dimensions [Afg75, Yod75]. Direct [NA76]. Directed [Ast78]. Dirichlet [KH73]. Discontinuities [Kal73, Kal74, Kal76]. Discontinuity [Ber74a]. Discontinuous [Ker74, SK75]. Discount [Put74]. Discrepancies [Ste78a]. Discrepancy [Stu77]. Discrete [Hos78a, Mey75, SvH79, BE73]. Discrete-Time [Hos78a]. Discussion [BDK^{+}73, CMSB75, Orn73a]. Disintegrations [Dub75]. Dissonant [Mat77]. Distance [GNS75]. Distinguishing [Kan75a]. Distortion [SNDL78]. Distortion-Rate [SNDL78]. Distributed [GS79, McD79, Van78a, VP76, Wei75a, Wic74b, WPR77]. Distribution [AK76, Ben75, BS77a, BS77b, Bur73, Gem73, GY75, Gle75, Goo76, Gro76, HM77, HW74, Ism77, JP75b, Let75, Mal73, MT75, MS75, O’B74b, Phi77, PS77, Roo76, SW78, SB75, Sim74, Sim75, Slu77, Van78a, Yam78, Dia77]. Distributions [Bil74a, BD75, Che74, Csi75, De 74a, Dok74, EM79, GM77, Gut74b, HT75, HM75, Hud76, Jam75, KS74, KS78, LMRS73, MS79, Mey75, Mil75, Nai77, O’C79, Pfa79, PR74, SC74, Sch77a, Sch78a, Sch74, SG75, Sie79, Teu75, Wes79, Wri73, Zab79]. Divergence [Csi75, KMPQ74]. Divisibility [Bon79, Ism77, Jam76, Mil79]. Divisible [Bro77, Eag75, Gro76, HT75, HM75, O’C79, SG75, Yam78]. Domain [Bas76, Owe73, Wic74a]. Domains [SS78]. Down [Eri77]. Downcrossings [Fra76]. Drift [Ste77c]. Dual [GS78]. Dualité [Alv76, GM73]. Duality [KD79, Nev76, Sch77b, Alv76, GM73]. Dym [Mar77b]. Dynamic [BFO74, Fre74c, KN79]. Dynkin [AK73].

E. [Sen73b]. Edgeworth [Hip77]. Elements [BR79, TW79, Wac78]. Embedded [Mon78b]. Empirical [BCCR79, Deo73a, Dud78, Dud79, GS79, GOR79, Jam75, MR75, O’R74b, Phi77, PS75, Rev76, Sen74b, SW78, Ste78a, Van78a, Wel77]. Empiricals [Sin75]. Ensemble [Alv76, GM73]. Ensembles [Air78]. Entrance [Cox77a, Kal77]. Entropy [PR73, ST75]. Enumeration [FG76]. Envelopes [Ven79]. Environment [Sol75, Tan77, Tor78]. Environments [Kap74a, Kap74b, KN73, Lin74c]. Epsilon [PR73]. Equation [Por75, Wei77]. Equations [Pro77, Sus78]. Equilibrium [Kal77, McD77]. Equivalence [Bak73, HT75, HM75, Sie76b]. Equivalent [KO73]. Erdos [Sho79]. Ergodic [BR77, Der73, Eis76, HL75, Jon76, Kie74, Kin73, Lig77, MC73, MI73, Orn73b, Rei77, Sch76b, Wri76]. Ergodicity [Gri79, Mat77, NT78]. Erlang [Oak76]. Error [Kan79, Sal79]. Errors [Bha75]. Escape [BF73, Get79a]. Esseen [CW77b, Eri73c, KK74]. Essential [LPQ78]. Estimate [Ben75, Kos73, Owe73]. Estimated [BCCR79]. Estimates [Ben75, Goo76, KK74, Ros76]. Estimation [BS77b, Kie79]. Estimator [BS77a]. Estimators [PKG79, Qui76, Qui77]. Euclidean [Gor73]. Events [AP79, Fre74a, Kap77, Mey73]. Every [Gil77]. Evolutinos [Ker74]. Evolution [Gue74]. Exact [Dav77]. Example [Ber76, DS75b, Jai76]. exceptional [Air78]. Exceptionnels [Air78]. Excessive [Cha74a]. Exchangeability [Kin78]. Exchangeable [AP79]. Exclusion [Lig76, Lig77]. Excursion [DI77, DIM77, Ver79]. Excursions [Ber73b, Get79b]. Existence [Bha78, Bha80, BD75, Kal75, Kes76, MS75, Pro77, Ram79a]. Exists [Ste73]. Exit [Cha77, GS73, Lai77b, LW79, Mai75, Mon77]. Exits [Por78]. Expansions [Eri73b, Hip77, Rei76]. Expectation [NR75, Sim77, dJ79]. Expectations [Isa79, Mal73, Pfa74, Rog74]. Expected [Kla74, Lin73, Nai77, Rob76, Ros73]. Explicit [Hal75]. Explosions [Pro77]. Exponential [EG73, MT75]. Exponentiality [KS74]. Expression [HN75]. Extended [CHL79, Kla73, McL78]. Extended-Valued [Kla73]. Extension [Mac73, Sho79, WZ77a]. Extensions [Dub74]. Extent [Mey75]. Extinction [Tor78]. Extrapolations [Hos78b]. Extrema [Sen76a]. Extremal [Adl78b, Fra76, Res75, Wei75a, Wei76]. Extreme [DO75, Dyn78, Gre75, Wei75b, Wel73]. Extremes [Ber74b, Ber80, Roo78, dHR79b].

F [Smy75]. Failure [PR74]. Falling [SC74]. Families [Ste77a]. Family [DV74, Mon75, PR74]. Features [LPQ78]. Feller [Chu73, Sen73b]. Fidelity [GNO75]. Field [Ros79c]. Fields [AH76, Adl77, Cox79a, Cuz78b, DV74, Deo75, Dob79, GZ78, Isa79, Kaw79, Kes76, Lan78, Nea78, PT79, Spi75]. Filtering [Bre73, Gab77b, Gab77a]. Fine [Rob74, Rob73]. Finite [Cuz75, Gab77b, Gia77, KT77, Kun74, Rip76b, Rob77, Rob83, RS77, Sim75]. Finite-Memory [RS77]. Finitely [Dub74, Dub75, PS76]. First [Gut74a, Gut74b, Kin75, Lai76d, Lai77b, LW79, Mil75, Por78, Rob76, Ros73, WR78, Woo76]. Fixed [Kie76b]. Flip [Gra78]. Fluctuations [Roo76, RS76]. Form [Wic74b]. Forms [GPW73, WW79, Wri73]. Formula [Oak76, Ste74, Ste75, Tay75a]. Formulas [Leh77]. Foundations [Rip76a, Rip76b]. Fourier [Dal78, Mar73b]. Frank [The74]. Freedom [Gro76]. French [Air78, Alv76, Bis79, BR74, GM73, Pra75]. front [Ano73g, Ano73h, Ano73i, Ano73j, Ano73k, Ano73l, Ano74g, Ano74h, Ano74i, Ano74j, Ano74k, Ano74l, Ano75g, Ano75h, Ano75i, Ano75j, Ano75k, Ano75l, Ano76g, Ano76h, Ano76i, Ano76j, Ano76k, Ano76l, Ano77g, Ano77h, Ano77i, Ano77j, Ano77k, Ano77l, Ano78g, Ano78h, Ano78i, Ano78j, Ano78k, Ano78l, Ano79g, Ano79h, Ano79i, Ano79j, Ano79k, Ano79l]. Function [Ben75, BS77a, BS77b, Bur73, Dal76, Dal78, Gar73, Jon76, Kli77, Mal74a, Mar77b, Mon77, Pit79, SNDL78, SW78, Ste73, Van78a, VP76, WZ77b]. Functional [Bol76, Deo75, DR78, Gar76, Igl74, Jam75, LR79b, Phi77, Sen76c, Vil74, Wic74b, Wic74a]. Functionals [AN74, Dud77, DI77, Dyn77, GS73, Goo76, Kes74, Lev75, O’R74a, Wan77]. Functions [Ber74a, Bil74c, Bla78, Cha74a, Cha73, Che79, Dav79a, Dub73, Dud73b, EP77, Gru75, HW74, Hel77, Hen73, Hip77, Hol79, Ism77, JJS75, Jos75, Jos77, KP78, Ken78, KD79, LR79b, Lyn79, Mal74b, Mal73, Mon78a, Mou79, O’C79, OP73, Phi77, Sav75, Sen76a, Sim74, Sim75, Wan79, Wei75b, Wol73, Wol75, Wol78, Yam78]. Further [Cox79b].

G [MS75, Smy75]. Galton [DS76, Hop77, Kap73]. Gap [Sus78]. Gaussian [Ber80, She77, Adl77, Ber73b, Ber74a, Ber74b, CR73, Cuz75, Cuz76, Cuz78a, Cuz78b, Cuz87, Dav76, Dav77, Deo73b, Deo73c, Deo74, Dob79, Dud73b, Dym78, EGBD75, HJSD79, HC78b, Hwa76, KO73, Kaw79, KG75, Kle76, Lai73, LS78, Lin74b, LdMR75, Man76, Mar73a, Mar73b, Mar74a, Mar77b, Mar73c, Mit74, MY76, Mit78, Mit79, Ood73, Pit77, PT79, Qua77, Wel73]. Genealogical [LM73]. General [BZ79, Bon79, Kes74, NA76, NT78, Ser75, Twe74a, Twe74b]. Generalisation [Les75]. Generalization [AK73, GNS75, Hwa77, Kie73, Kie76a, Kle78, Oak76]. Generalized [Dob79, EP77, Kie75, MR75, Sch77a, Sch78a, Sen74a, Sen77, Sim74, Sim75]. Generated [Adl78b, Wei75a]. Generation [Fea76]. Generators [Gzy78, Kun74]. Genetic [Eth79, LM73]. Genetics [Saw76]. Geometric [NT78]. Geometry [Csi75, EP77, McK73, Rip76a]. Germ [Kni79]. GI [MS75]. GI/G/s [MS75]. Gibbs [Nea78]. Given [Bru75, LR76, Let75, PS79, Sch74]. Glivenko [Sin75]. Governed [Fel77]. Graph [Yod75]. Graphical [Har78]. Graphs [Lin73, NR75]. Great [BdH74]. Group [Kie75, Sie76a]. Groupes [BR74]. Groups [BE73, BR74, DMW74, EP77, FP75, Kun74]. Growing [Tei79]. Growth [KN73, LP79, Mon77].

H [Mar77b]. H. [Mar77b]. Half [Cra76]. Hammersley [dJS79]. Harding [Smy75]. Harmonic [Sav75]. Hausdorff [Adl77, Dav77, PT77, Tra77, Yod75]. Hawkins [Hey78]. Hazard [EM79, MS79]. Heads [KT75]. Heavy [WPR77]. High [Ber73b, Lin74b, LdMR75, Mar73c]. Hilbert [Bro77, KK74]. Hitting [Gem73, Sch78b]. Holder [Dav76]. Holomorphic [CW77a]. Homogeneous [Ros79c]. Houses [DS79]. Hypercontraction [AG76]. Hypothesis [Hey78, Ste73].

Ibragimov [Les75]. Identically [GS79, Wic74b, WPR77]. Identification [Bri75, CMSB75, Ste73]. Identity [AK73, De 73]. IFRA [BS76]. II [Alv76, Lig77, McD78, Sch78a, Twe74b]. Immigration [Kap73, Qui76, Qui77]. Improved [Jos77]. Improvement [Hey78, Maj79]. Including [Bri74]. Increasing [DV74, EM79, MS79]. Increment [SK75]. Increments [Bri74, CR79, GZ78, Gre73, Gre75, HS73, KG75, Mil75, Wal74b, Wic73c]. Independence [HT79, Kal78, Lin74b, Ste77c]. Independent [BP79, CL73, Eri73c, Gab77b, GS79, GY75, Gle75, Gre73, Gre75, GPW73, HGS79, Jof74, Ker77, Lai77b, LPQ78, McD79, Mil75, Nag79, Ram79a, SK75, She74, Sho78, Sin75, Smy74, Van78a, Wac78, Wal74b, Wei75a, Wic73c, Wic74b, Wri73, WPR77, Mar74b]. Independent-Increment [SK75]. Index [Bas76, Gab77a, LM73, Owe73]. Indexed [Ast78, Gab77b, Gut74a, Nea78, Wri79]. Indicator [LR79b]. Indices [De 74a, Gut78]. Individual [BR77, Rei77]. Induced [SK75, Sen76b]. Inégalités [Pra75]. Inequalities [Bru75, Bur73, Gal77, KKO77, Kan75c, KS74, LR77a, Rog74, Slu77, Sto73, Tur73a, Tur73b, Pra75]. Inequality [Bur79, Che79, Gab77b, Gar73, Har77, Kli77, McL75, Mou79, Pit77, Pit79, Sch76a, Wel77]. Infima [HGS79]. Infinite [Bon79, Fis75, GS76, Gia77, HL75, HS76, Ism77, Jam76, Lig73, Lig74, Mil79, PSW75, Sch76b, Spi75]. Infinitely [Bro77, Eag75, Gro76, HT75, HM75, O’C79, SG75, Yam78]. Information [Ano73n, Ano74n, Ano75m, Ano76m, Ano77m, Ano78m, Ano79m, GNS75, Kie79]. Initial [Ker74]. Insensitivity [Sch77a, Sch78a]. Instability [Tor78]. Instance [dJS79]. Integer [Jam76]. Integer-Valued [Jam76]. Integers [Lig78]. Integrability [Der73, Gut79]. Integrable [HK77, Lan76]. Integral [Pro77, Wei77]. Integrals [Dud77, Hip77, HC78b, Kal75, Pro79, WZ76, WZ77a, WZ77b]. Integration [Gar78]. Inter [SB75]. Inter-Record [SB75]. Interacting [HL75, HS76]. Interactions [Har74, Lig73, Lig74]. Interpolated [Kus74]. Interpolation [Hea77, Sal79]. Interpolator [Sal79]. Interval [Won79]. Intervals [EN74, Oak76, Van78b]. Introduction [Chu73]. Invariance [Hal77, Hal79, HGS79, HS73, Kai76, Maj79, McL74, McL77, McL78, Sen73a, Sen76a, Sen76b, SS78, Tei73, Wel78]. Invariant [Bha78, Bha80, BP73, Hal75, Hop77, Kun74, NA76]. Inverse [KP79, Mar77b]. Inversion [dHR79a]. Involving [LP79]. Isotone [Wri79]. Isotrope [Stu77]. Itô [Dud77]. Iterated [Ber79, Bol78, HS73, Hey74, JP75a, Jam75, KP78, KT77, KK74, LS78, Ood73, Phi77, Qua77, She74, Sto73, Tei74, Tom75a, Tom78, Wic73b, Wic73c, Wic74b, Wic74a].

Jagers [Mod78]. Jensen [Sch76a]. Joint [KF77, SB75].

K. [Wil74]. Kakutani [Van78b]. Karlin [Hwa77]. Kendall [Smy75]. Kernel [Tay75b]. Kernels [Le 73, MC73]. Kesten [Cox79a]. Khintchine [Kos73]. Kind [Kal73, Kal74, Kal76]. Kingman [BDK^{+}73]. Knight [Glo79]. Known [Lem74]. Kolmogorov [Dym78, Sen73a]. Komlos [HR79].

Lack [MT75]. Ladder [Lai76a]. Landau [Sen73b]. Large [AV75, BZ79, BW74, BE75, Boo75, BR79, DT79, GS77, GOR79, Gut78, HJP76, KD79, Kos73, Lyn79, McD79, Nag79, Rob77, Smy73, Ste78b, TW79, Wel78, Rob83]. Last [Cha77, GS73, Mon77, Mon78a]. Late [Kai76]. Lattice [Ben75, Har74, LR76]. Lattices [Bru75, SW76]. Law [AP79, AV75, Bas76, Ber79, Bol78, BR79, EG73, HS73, JP75a, Jam75, Kal77, KP78, KK74, LS78, Ood73, Owe73, Phi77, Qua77, Sen75, Slu77, Sto73, Tei74, Tom78, VW74, Wel78, Wic73b, Wic74b, Wic74a, Win73, Wri79, BE75, Fre73, Fre74b, Gut78, HJP76, Lai77a, WPR77, dJS79]. Laws [BW74, CR73, Cox77a, DT79, Eag75, Gut78, JT73, KT77, Moh76, Mor76, Smy73, TW79, Wic73c, Wic74a, McL75, MY76]. leading [Dia77]. Learning [The74]. Lebesgue [Dub74]. Lebesgue-like [Dub74]. Lemma [Che78, Fre73]. Length [Kie76b, MS75]. Level [AH76, Adl78a, Adl80, Lin74b, LdMR75, Mar77a, Mar73c]. Lévy [Ano73m, AM78, Gzy77, Kar78, Kar79, Kaw79, Loe73, O’C79, Tra79]. Life [BdH74]. Lifts [Sza77a]. like [Dub74]. LIL [Jai76]. Limit [Afg75, AE78, AMN78, Bas74, Bas76, Ben75, Ber73a, Bha77, Bol76, Cuz76, Dar75, Dar76, De 74a, DO75, Deo75, DS76, Dud78, Dud79, Dur78, DR78, EK74, Eth79, Gar76, Gia77, Gis73, Gor73, Gri78, Gue74, Gut74b, Hah77, HC78a, HJP76, HT79, Igl74, Kai75a, Kai75b, Kap75, Ker74, KW73, Lai74a, LS73, LMRS73, M.74, Mac78, Mae78, Mat75, McD79, McL74, Mey73, Mic76, Min74, Mor76, Nea78, Nor74, Num79, O’B74a, Oma77, Ros73, Sen76c, Sho78, Ste77b, Ste73, Swe77, Tan77, Tei73, Tho77, Tom76, Vil74, Won79, Zin77]. Limiting [CL73, MS75, Mit74, O’B74b, Wei75b]. Limits [Rud73, Wac78]. Line [Cha74c, GN79, Lee74, PS77, WN73, Wol78]. Linear [Hel77, Hey74, Hos78b, Lin73, Mac73, NR75, Pho78, Por75, Sal79, Sen76c, SW78, TW79]. Linearity [LR79a]. Linearly [Sim74]. Lines [Ker74]. LLN [DMW74]. Local [Bas76, Ben75, Cin75, Cuz78a, Cuz78b, Cuz87, Dav76, DS76, Fri79, JT73, Kai75a, Kle76, Kni73, Mal74b, Mal74a, McD79, Min74, PT79, Sho78, Tho77, VW74, Wol73]. Locally [DMW74, Rip76b]. Location [Wei75b]. Log [Hey78]. Logarithm [Ber79, Bol78, HS73, Hey74, JP75a, Jam75, KP78, KT77, KK74, LS78, Ood73, Phi77, Qua77, She74, Sto73, Tei74, Tom75a, Tom78, Wic73b, Wic73c, Wic74b, Wic74a]. Longest [Van78b]. Loss [Gri76, Oak76]. Lower [HJSD79, JJS75, Kan79, Kos74, LS78, Sim75, Ste79].

M. [The74]. Majorization [EP77]. Many [Rod78]. Many-armed [Rod78]. Mappings [Won79]. Marcinkiewicz [Gut78]. Market [GS74]. Markov [CI80, AG76, AMN78, Cha74a, Cha76a, Cha77, CI78, Cog75, Cox77a, Cox79a, Dur78, Dyn77, Fel77, Get79b, Gri76, Gri78, Hal75, Har76, Har77, Har78, Isa79, JP77, Kai75b, Ker74, Kes74, KS75, Kes76, Kiu75, Kle78, Kni79, Kur75, Kus74, Lai74b, Lev73, Lev75, LS73, Lin74a, MI73, McD77, McD78, Mil78, Mye76, Nev76, Nor74, NA76, NT78, Num79, O’B74b, Sch77a, Sch78a, Sch77b, Sch76c, Sie76b, Sil74, Spi75, Twe74a, Twe74b, Win78, The74]. Markov-Dependent [Isa79]. Markovian [KD79, Nor75, Tay75b, Wan79]. Martingale [Ber76, Bro78, Bur79, CW77a, Cha73, Che79, Eag75, Gil77, Hal77, Hal79, HS76, Mac73, McL78, Tom75a, Wel77]. Martingales [Bur73, Fre75, Gab77a, Gar73, Hea77, HS73, JJS75, Kan75b, Lai74b, Lan76, Lan78, Loy76, Mon76, Pra75, PS79, VW74, WZ76]. Martintote [Gre74]. Math [Mar74b]. Mathematical [KD79]. Matrices [Wac78]. Matrix [Wac78]. Maxima [Dav79b, EGBD75, LdMR75, Mar74a, Mit74, MY76, Mit78, Mor76, O’B74b, Res73, Wic74b]. Maximal [Che79, Gri73, Jon76, McL75, Ros76, Sto73]. Maximum [GY75, Kaw79, Leh77, O’B74a, Sho79]. McGregor [Hwa77]. McKean [Mar77b]. McMillan [Kie76a, Kie73, Kie75]. Mean [BZ79, DS75a, Gut74a, KT77]. Meander [DI77, DIM77]. Means [Bru75, JM78, LR76, Ste77a, LR79b]. Measurable [Dud73a]. Measure [BP73, Dav77, DH79, Hoc78, Kha77, NA76, PS77, PT77, Sza77b]. Measure-Invariant [BP73]. Measures [AT78, AM78, Bak73, Bha78, Bha80, BE73, Bro77, Cha74c, Che75a, De 77, Dub74, Dud78, GOR79, Hal75, Hop77, Kar78, KS74, Kie74, Kum73, Kun74, Kuo73, McD77, Ram74, Ram79b, Res74, Rin76, Rud73, SK75, Sie76a, Twe74b, dA75, Dud79, Kan75a, Kar79, She77]. Medians [Tom75b]. Memory [Gri76, MT75, RS77]. Method [O’B75, Ros73]. Methods [Dal78, Har78]. Metric [ST75]. Metrics [O’R74b]. Migration [Saw76, Saw77]. Mills [Ste79]. Minima [Dav79b]. Minimal [Sal79]. Minimax [Bre73]. Minimization [Csi75]. Minimum [Kie76b]. Miscellaneous [Ano73a, Ano73b, Ano73c, Ano73d, Ano73e, Ano73f, Ano73g, Ano73h, Ano73i, Ano73j, Ano73k, Ano73l, Ano74a, Ano74b, Ano74c, Ano74d, Ano74e, Ano74f, Ano74g, Ano74h, Ano74i, Ano74j, Ano74k, Ano74l, Ano75a, Ano75b, Ano75c, Ano75d, Ano75e, Ano75f, Ano75g, Ano75h, Ano75i, Ano75j, Ano75k, Ano75l, Ano76a, Ano76b, Ano76c, Ano76d, Ano76e, Ano76f, Ano76g, Ano76h, Ano76i, Ano76j, Ano76k, Ano76l, Ano77a, Ano77b, Ano77c, Ano77d, Ano77e, Ano77f, Ano77g, Ano77h, Ano77i, Ano77j, Ano77k, Ano77l, Ano78a, Ano78b, Ano78c, Ano78d, Ano78e, Ano78f, Ano78g, Ano78h, Ano78i, Ano78j, Ano78k, Ano78l, Ano79a, Ano79b, Ano79c, Ano79d, Ano79e, Ano79f, Ano79g, Ano79h, Ano79i, Ano79j, Ano79k, Ano79l]. Mixing [AE78, BS77b, Deo73a, Deo73b, Dym78, GB77, JJS75, Lai77a, Mar74c, MR75, Mey73, Mit79, Mor76, Nea78, Ros79b, Ros79a, Sen74b]. Mixingales [McL77]. Mixtures [Eag75, KS74, Ram74, Ram79b]. Model [BG79, HL75, LP79, LM73, Mat77, Sav76, Saw76, Saw77, Sch78b]. Models [EM73, Eth79, Gue74, LPQ78, MS79, The74]. Moderate [GB77, Mic74]. Modified [HW74]. Modulus [Ben75]. Moment [CHL79, KS74]. Moments [Cuz75, Gut74b, Lai76a, Mal74a, Woo76]. Monotone [Sie76b]. Monotonic [Won79]. Monro [GH76, Ker77, Ker78]. Motion [Dav78, Dav79a, Dro73, Fra73, JT73, KP79, Kha77, Kue75, Mon78b, Mye74, Qua77, Ste77c, Tay75a, Tra79, Wal74a, Wan77, Wan79, Wen77, Yod75]. Motions [Gor73]. Motoo [Wic73b]. Move [Nor74]. Moving [Ber73b, BE73, Kan79, Lai73, Lai77b, Por78, Roo78]. Multidimensional [Bha77, Bha78, Bha80, GZ78, Gut78, Kap74b, Rev76, Rya76, Sen74b, Ste78b, Swe77]. Multiparameter [Smy76, Wic73c]. Multiple [Fla76, GN79, HC78b]. Multiplicity [DV74]. Multiply [Nea78]. Multitype [AN74, Hop76, Hop77, Kap73, Tur73b]. Multivalued [Ros79d]. Multivariate [Ber79, EM79, HT79, MS79, Mat75, PR74, Ste79]. Mutivariate [Sen73a].

near [Alv76, GM73]. Nearest [HN75, Lig78, Sch74, WN73]. Nearly [Mye74].

Necessarily [Wri73]. Needed [BF73]. Neighbor [HN75, Lig78, WN73]. Neutral [Dok74, Saw77]. Ney [Wil74]. Nikodym [She77, SK75]. Non [BD75, BE73, KO73, MI73, Nor75, Por75]. Non-Anticipative [KO73]. Non-discrete [BE73]. Non-Existence [BD75]. Non-Linear [Por75]. Non-Markovian [Nor75]. Non-Stationary [MI73]. Nonblock [GNO75]. Nondeterminism [Cuz78a, Cuz87]. Nonergodic [Gri78, SNDL78]. Nonexistence [KW73]. Nonidentically [McD79, Van78a, Wei75a]. Noninteracting [Kal78]. Noninvariance [Sch76c]. Nonlattice [Sho78]. Nonlinear [Kan79]. Nonnegative [Gil77, MC73]. Nonparametric [Sim74]. Nonstable [Rub77]. Nonstationary [KN79, McL77]. Nonuniform [Hey75, Mic76]. Nonuniformly [Lan76]. Nonuniqueness [GG77]. norm [O’R74b]. Normal [Bas76, Bha75, Hel77, JP75b, LR77a, Owe73, Rei74, RR73, Ste79, HW74]. Normality [Eri74, Hey75, Hol79, KS74, Les75]. Normalized [DS75a, EK74, Sho79, LMRS73]. Norman [The74]. Normed [Eri76, TW79]. Note [Bha80, Bil74b, BE73, Bre73, Cha73, Che78, Dar75, Dar76, Deo73a, Deo73b, Deo74, Dud73a, Fra76, Fre73, Gem73, Glo79, Kal76, Kan75a, Kap74a, Kar79, Kha73, Kie76a, Kud74, Kum73, Lin74b, MT75, Pit79, PS75, Rob73, Rob74, Ros79b, Sch76a, Sen76b, Sen75, Ste75, Wei77, Wen77, Wic73a, Win73, Zin77, Dar76]. Notes [Adl80, Ano74m, Ber80, Cha74b, CI80, Qui77, She77]. Null [Dur78]. Number [Ano73a, Ano73b, Ano73c, Ano73d, Ano73e, Ano73f, Ano73g, Ano73h, Ano73i, Ano73j, Ano73k, Ano73l, Ano74a, Ano74b, Ano74c, Ano74d, Ano74e, Ano74f, Ano74g, Ano74h, Ano74i, Ano74j, Ano74k, Ano74l, Ano75a, Ano75b, Ano75c, Ano75d, Ano75e, Ano75f, Ano75g, Ano75h, Ano75i, Ano75j, Ano75k, Ano75l, Ano76a, Ano76b, Ano76c, Ano76d, Ano76e, Ano76f, Ano76g, Ano76h, Ano76i, Ano76j, Ano76k, Ano76l, Ano77a, Ano77b, Ano77c, Ano77d, Ano77e, Ano77f, Ano77g, Ano77h, Ano77i, Ano77j, Ano77k, Ano77l, Ano78a, Ano78b, Ano78c, Ano78d, Ano78e, Ano78f, Ano78g, Ano78h, Ano78i, Ano78j, Ano78k, Ano78l, Ano79a, Ano79b, Ano79c, Ano79d, Ano79e, Ano79f, Ano79g, Ano79h, Ano79i, Ano79j, Ano79k, Ano79l]. Number [Cuz75, Cuz76, Gle75, Kie76b, Lin73, NR75]. Numbers [DT79, Smy73, TW79, dJS79, AV75, BW74, BE75, BR79, Gut78, HJP76, Wel78].

Observations [PKG79]. Observed [Kai75b]. Occupancy [Hol77, Lig74, SC74]. Occupation [GH73, Mar73c]. Odd [Gro76]. One [AK76, CR73, Cox77b, Cox79b, Kes76, LdMR75, Put74, Ros79c, Sen75, AP79]. One-Dimensional [AK76, Cox77b, Kes76, Put74, Ros79c, Cox79b]. ones [LPQ78]. Opérateurs [Pra75]. Operator [AG76, BFO74, Fre74c]. Operators [Lin74a, Pra75]. Optimal [Bis79, BFO74, CP76, Fel77, Fre74c, GS74, KN79, Kla73, Wal74b, Wal74a, Bis79]. Optimality [Put74]. Optimization [DM74]. Order [Cam75, Hel77, Ivk75, JP75b, Mal73, Nai77, Sen76b, Hos78b]. Order- [Mal73]. Ordered [De 73, Har77, KKO77, Sim74, Smy74, Wri79]. Ordering [KK78, Ste77b, dJ79]. Orders [KF77]. Ordinary [Sus78]. Ornstein [GNS75, Orn73a]. Orthogonal [BW74]. Oscillation [Gri73, Kha77]. Oscillations [Imh76]. Other [JP75a]. Overlapping [Oak76].

P [Mar77b, Mod78, Wil74]. pages [Ano73a, Ano73b, Ano73c, Ano73d, Ano73e, Ano73f, Ano73g, Ano73h, Ano73i, Ano73j, Ano73k, Ano73l, Ano74a, Ano74b, Ano74c, Ano74d, Ano74e, Ano74f, Ano74g, Ano74h, Ano74i, Ano74j, Ano74k, Ano74l, Ano75a, Ano75b, Ano75c, Ano75d, Ano75e, Ano75f, Ano75g, Ano75h, Ano75i, Ano75j, Ano75k, Ano75l, Ano76a, Ano76b, Ano76c, Ano76d, Ano76e, Ano76f, Ano76g, Ano76h, Ano76i, Ano76j, Ano76k, Ano76l, Ano77a, Ano77b, Ano77c, Ano77d, Ano77e, Ano77f, Ano77g, Ano77h, Ano77i, Ano77j, Ano77k, Ano77l, Ano78a, Ano78b, Ano78c, Ano78d, Ano78e, Ano78f, Ano78g, Ano78h, Ano78i, Ano78j, Ano78k, Ano78l, Ano79a, Ano79b, Ano79c, Ano79d, Ano79e, Ano79f, Ano79g, Ano79h, Ano79i, Ano79j, Ano79k, Ano79l]. Pairs [GS78]. Paper [BDK^{+}73, CMSB75, Orn73a]. Paradox [Ber74a]. Parameter [Cox77b, Cox79b, Goo76, Kie79, OP73, Tra77, Tra79, Win73, Win78, Yod75]. Parameters [BCCR79]. Parental [Sav76]. Part [Sch77a, Sch78a]. Partial [De 73, DS75a, Gri76, Gut74a, HM77, HGS79, KK78, LS78, Mac78, Mit78, Ser73, Ste77b, Wic74b, Wic74a]. Partially [Har77, Kai75b, KKO77, Smy74, Wri79]. Particle [Lig73, Lig74, PSW75, Sch76b, Sch78b]. Particles [EN74, Kal78]. Particular [EGBD75]. Passage [Gut74a, Gut74b, Lai76d, Mil75, Rob76, Ros73, WR78, Woo76]. Passages [EGBD75]. Path [Cam75, Hoc78, Kal75, Kue75, Mil78, PT79]. Pathak [Sen75]. Paths [Le 73, Mar77a, PR73]. Paul [Ano73m, Loe73]. Percolation [WR78]. Perez [Kie76a, Kie73]. Perfect [Ram74, Ram79b]. Periodic [Kie74]. Permutational [Sen76c]. Persistently [KN79]. Personal [Cra76]. Phenomena [Gri73]. Piecewise [Won79]. Planar [Wen77]. Plane [WZ76, WZ77a, WZ77b]. Plans [KN79]. Play [GS77]. Point [Bri75, Bro78, CMSB75, GS75, Lin75, Rip76b, Rip76c, Wes76]. Points [Dyn78, EK74]. Pointwise [Dun77, Gri79]. Poisson [Bro78, Che74, Che75b, Fre73, Fre74a, Kap77, Mey73, Mon75, Ser75]. Poisson-Type [Mey73]. Polaire [Alv76, GM73]. polar [Alv76, GM73]. Pollaczek [De 73]. Polygons [BKM78]. Population [Rob77, Saw76, Rob83]. Positive [Bol76, Igl74, SC74]. positivity [NT78]. Posterior [Dok74]. Postulates [Ham74]. potential [Air78]. Potentials [Nev76]. Potentiel [Air78]. Prediction [Eis76, Kan79, Kni79, Lin75, Pho78]. Predictive [Kni75]. Predictors [LR79a]. Près [Alv76, GM73]. Preserves [GM77]. Preserving [JM78, LPQ78]. Principle [HGS79, Kai76, Maj79, McL77, McL78, Sen73a, Sen76a, SS78]. Principles [Hal77, Hal79, HS73, McL74, Sen76b]. Priority [Ste75, Mar74b]. Probab [Ano73a, Ano73b, Ano73c, Ano73d, Ano73e, Ano73f, Ano73g, Ano73h, Ano73i, Ano73j, Ano73k, Ano73l, Ano74a, Ano74b, Ano74c, Ano74d, Ano74e, Ano74f, Ano74g, Ano74h, Ano74i, Ano74j, Ano74k, Ano74l, Ano75a, Ano75b, Ano75c, Ano75d, Ano75e, Ano75f, Ano75g, Ano75h, Ano75i, Ano75j, Ano75k, Ano75l, Ano76a, Ano76b, Ano76c, Ano76d, Ano76e, Ano76f, Ano76g, Ano76h, Ano76i, Ano76j, Ano76k, Ano76l, Ano77a, Ano77b, Ano77c, Ano77d, Ano77e, Ano77f, Ano77g, Ano77h, Ano77i, Ano77j, Ano77k, Ano77l, Ano78a, Ano78b, Ano78c, Ano78d, Ano78e, Ano78f, Ano78g, Ano78h, Ano78i, Ano78j, Ano78k, Ano78l, Ano79a, Ano79b, Ano79c, Ano79d, Ano79e, Ano79f, Ano79g, Ano79h, Ano79i, Ano79j, Ano79k, Ano79l]. Probabiliste [BR74]. Probabilistic [BR74, Ken78, Lin77, RR73]. Probabilities [Boo75, Cog75, Dok74, Dub75, Fre75, GB77, HN75, Hwa77, Lai74b, Lai76b, Mic74, Mic76, Rob77, Rob83, Sch78b, Ste78b, WN73, Wic73b, Wri73]. Probability [Bak73, Bil74c, Cha74b, Che75a, Chu73, Cra76, Csi75, Dro73, FG76, Fis75, GOR79, Hud76, KF77, Kie75, Kie76b, Kum73, KS78, O’R74a, Orn73b, Por78, PS76, Ram74, Ram79a, Rob73, SC74, Sha79, Sie76a, Ste73, dJS79, Ano74m, Rob74]. Problem [BS76, CP76, CG77, GS75, GS76, Gia77, Gri73, Kin75, Rod78, RS77, Sza77a, WN73, KP79, Mar77b]. Problems [Csi75, DM74, EN74, Fre74c, Hol77, Ker74, Lai73, Lai76d, Muc73, SC74, Sie76b, Tak74]. Process [AK76, BE75, Bri75, BCCR79, Cin75, CMSB75, CR79, Cuz76, Del77, DS76, Dud73b, EGBD75, Get79a, Get79b, Gis73, Goo76, GH76, GM77, Gri79, Hey74, Kap73, Ker77, Ker78, Kin75, KW73, Lig76, Lig77, Mar77a, O’B74a, OP73, Pho78, PS75, Qui76, Qui77, Rev76, Rip76b, Ros79c, Tra77, Wal74b, Wel77, Wes79, Wri76, dHR79a]. Processes [Adl78b, Asm76, AN74, AMN78, Ber73b, Ber74b, Bil74b, Bri74, Bro78, CW77a, CR73, Cam75, CNW73, Cin76, Cin77, Cuz75, Cuz78a, Cuz87, DM78, Dav76, Dav77, DO75, Deo73a, DMW74, Dur78, Dym78, Eri73a, EF75, EM73, Est75, Fea76, Fel77, Fra76, GS75, GS79, GH73, GB77, GG77, Gre75, Gri78, Gri74, Gzy77, HK77, Hal75, Ham74, Har76, Har77, Har78, Hen73, HS73, HS76, HL78, Hop76, Hop77, Hos78a, Hos78b, HC78b, Ivk75, JJS75, Kal73, Kal74, Kal76, KO73, Kan79, Kap74a, Kap74b, KN73, Ker74, Kiu75, KG75, Kle76, Kle78, Kni75, Kni79, KH73, Kuo73, Kur75, Lai73, Lai74b, Leh77, Lev73, Lev75, LS73, Lin74b, LdMR75, Lin74c, M.74, Mac78, MI73, Mar73a, Mar73b, Mar74a, Mar77b, Mar73c, McD77, McD78, MR75, Mil73, Mil75]. Processes [Mil78, MS75, Mil79, MY76, Mit79, Mod78, Mon77, Mon78a, Mon78b, Mor76, Nev76, Nor74, Nor75, O’B75, O’R74b, Ood73, PR73, PT77, PG77, Qua77, Res75, Rip76c, Roo78, RS76, Sal79, Sav75, Sch77a, Sch78a, Sch77b, SK75, Sen74b, ST75, Sie76b, Smy76, Ste78a, Tan77, The74, Tur73a, Tur73b, Ven79, Wan75, Wei75a, Wei75b, Wei76, Wes76, Wic73a, Wic73c, Win78, Ber80, Dud73a, Wei77, Wil74]. Product [AG76, Fre73]. Products [Let77]. Professor [BDK^{+}73, CMSB75, Orn73a]. Profitable [GS77]. Programmable [Bla78]. Programming [BFO74, Fre74c, KN79, KD79]. Proof [Cox79a, De 73, Lin77, RR73, Sen75, Van78b, Wes76, Lem74]. Proper [BD75]. Properties [Bar75, CNW73, Cuz78b, Hos78a, Kal75, Ken78, Kla73, Kle76, PT79, PT77, Ste77b, Twe74a, Van78a]. Property [Hwa76, Kni79, MT75]. Proximity [GG77]. Pseudorandom [dJS79]. Publications [Ano73m]. Pure [KT75]. Purely [Lan76].

Quadratic [GPW73, KG75, Mon76, Wan77, WW79, Wri73, de 74b]. Qualitative [Rob73, Rob74]. Quantiles [Rei74, Rei76]. Quasi [Bis79, Lin74a]. Quasi-Compact [Lin74a]. quasi-stopping [Bis79]. Quasi-Temps [Bis79]. Quelques [Air78]. Queue [Bri74, Kap75, MS75]. Quick [Lai73, Lai76c, Lai77a].

R [NT78]. R-positivity [NT78]. Radial [Mil73]. Radon [SK75, She77]. Random [AH76, Adl78b, AT78, AV75, Bar75, BW74, Ber73a, BP79, Bla76, BE73, Bol76, BR79, CP76, CL73, Cox79a, DM78, De 74a, Deo75, Dob79, Dok74, Dur79, EN74, EK74, Eri76, Eri77, Ete76, FP75, Fra73, Gab77b, GS79, GZ78, GY75, Gor73, GM77, GS78, GPW73, Gut78, HGS79, Hey78, Igl74, Imh76, Jam76, Jof74, Kai75a, Kai76, Kal73, Kal74, Kal76, Kap74a, Kap74b, Kar78, Kar79, Kaw79, KN73, Ker77, Ker74, Kes76, Kha77, KT77, KMPQ74, Kue76a, Kue76b, KZ79, Lai76a, Lai77b, LW79, LR77b, Lee74, Lem74, Let75, Lev73, Lin75, Lin74c, Lin73, Man76, Mar73b, Mar74b, McD79, Mye74, Nag79, NR75, Nea78, O’B74b, Oak76, Phi77, PS77, Rip76b, Sch74, She74, Sho78, Sin75, Smy73, Smy74, Sol75, Spi75, Ste77b, Tan77, TW79, Tei79]. Random [Tom76, Tor78, Van78a, Van78b, Ven79, Vil74, Wac78, Wei75a, Wes79, Wic74b, Wic74a, Wri73, WPR77]. Randomly [Gre73, Gut74a, Ser73]. Range [CI78, CI80, Ete76, Fla76, Gar78, Lig73, Tra77, Tra79, Yod75]. Rank [FP75, JP75b, Sen76c, Tho77]. Rao [BS77a]. Rapidly [Tei79]. ’Rare [Mey73]. Rate [Afg75, EM79, Fra73, Kha73, MS79, SNDL78]. Rates [Asm76, Boo75, Gut78, Lai77a, Lyn79, O’R74a, PR74, Saw77, Ste78b, Stu77, Ven79]. Ratio [Num79, Oma77, Ste79]. Ratios [Der73, Rud73]. Ray [Glo79]. Reaches [Kal77]. Real [Cha74c, Kie79, Lee74, PS77, Wol78]. Receivers [Ahl74]. Recollections [Cra76]. Reconstructability [Cha76a]. Reconstructing [HM77]. Record [Res73, SB75, Wes79]. Record-Time [Wes79]. Recurrence [Bha78, Bha80, Eri76, GHZ76, Hwa76, Let77]. Recurrent [CI78, CI80, Dur78, FP75, Fla76, Kai75a, Num79, Twe74a]. Refinements [Bha77]. Reflecting [Bla76, Sie76b]. Reflection [EP77, Tei73]. Regenerative [Gri73]. Region [Ahl74]. Regression [Mac78, VP76, Wri79]. Regular [BD75, GM77, Mat75, Ram79a, Zab79]. Related [DS78, Hoc78]. Relatedness [LM73]. Relation [Ver79]. Relations [FG76]. Relationship [LR79b]. Relative [MR75]. Reliability [LP79]. Remainder [Ben75, HC78a]. Remark [Fre74b, Mal74b]. Remarks [Ros79b, Ros79a]. Renewal [AMN78, Cha76b, CHL79, Dal76, DM78, Dal78, Kes74, Lai76a, Lai76d, Let77, Lin77, MS75, Moh76, Rya76, Sza77b, Woo76]. Renormalizing [BG79]. Replacement [Fel77]. Representation [Fol73, Hal79, Ivk75, Kle76]. Representations [CW77a, Hop77, KO73]. Reproducing [Le 73]. Residual [BdH74]. Residuals [Mac78]. Resolvents [Tay75b]. Respect [SK75]. Result [Bon79, Hey74, Lem74, Wes76]. Results [Adl78b, Cox79b, GS79, Hol77, Kap74b, KZ79, Les75, Mic74, Qui76, Qui77, Saw76, She74, Tom75a]. Retournement [Bis79]. Return [Kai76]. reversal [Bis79]. Reversible [LS73]. Review [Chu73, Mar77b, Mod78, Smy75, The74, Wil74]. Reward [BFO74, Fre74c]. Riemann [Hey78]. Robbins [GH76, Ker77, Ker78]. Robust [Hos78b]. Robustness [PKG79]. Rosenblatt [Mar74c]. Rules [Kla73, Kla74].

s [MS75]. Sample [BZ79, Ber74a, Dud73b, Hah77, HK77, Hen73, JM78, Kan75b, Kue75, Lai76b, M.74, Mar77a, Mon77, Mon78a, OP73, PT79, Rei76, Sen76a, Ste77a, Wac78, WZ77b, dHR79b, PS75]. Sample-Continuity [Hah77, HK77]. Samples [HM77, Mit78, Rob77, Rob83]. Savage [PS75]. Sawyer [O’R74a]. Scale [Kan75c, Wei75b]. Scheme [Kos74]. Schur [EP77]. Second [Cam75, Hos78b, Ivk75, Kal73, Kal74, Kal76, Mor76]. Second-Order [Ivk75, Hos78b]. Secretary [GS76, Gia77, Muc73, RS77]. Selected [Ser73]. Selectively [Saw77]. Self [Dob79, Dun77, Eri73a, LMRS73, SvH79]. Self-Adjoint [Dun77]. Self-Decomposability [SvH79]. Self-normalized [LMRS73]. Self-similar [Dob79]. Semi [AMN78, Kiu75, McD77, McD78, Sch77a, Sch78a]. Semi-Markov [AMN78, McD77, McD78, Sch77a, Sch78a]. Semi-Stable [Kiu75]. Semigroups [Gzy78, Kur75]. Semimarkov [Cin76, Cin77]. Semimartingales [Fol73]. Seminorms [HJSD79, dA75]. Semiregenerative [McD78]. Senders [Ahl74]. sense [Wri76]. Sensitive [Put74]. Separable [Bil74b, BR79]. Sequence [GY75, Loy76, dJS79]. Sequences [Bar75, BW74, Dav79b, Deo73a, Deo73b, Deo73c, Deo74, GHZ76, Gre73, Gri74, Kie76b, KT75, Kos74, Lai76b, Lai77a, LS78, Let77, Mac78, Man76, Mey73, Mit74, Mit78, Qua77, Roo76, Tom75a, Wel73]. Sequential [Man76]. Series [Gab77b, Hwa76, Kal74, KMPQ74, Mar73b, Wic73a]. Set [Cox77b, Dav77, Gab77a, Gri78, Har76, Har78, Jof74, Wri79, Cox79b]. Set-Valued [Gri78, Har76, Har78]. Sets [Adl78a, Adl80, AG76, AV75, Ast78, BP73, Eri76, Fis75, Gab77b, Mat75, Pit77, Rip76b, Smy74, Air78]. Shannon [Kie76a, Kie73, Kie75]. Shape [Dur79]. Sharp [Bur79]. Sheet [Adl78a, Adl80]. Shifts [Kur75, Mar74c]. Shock [EM73, Fel77, MS79]. Short [Che78]. Sieve [Hey78]. Siever [Lyn78]. Sign [Lyn79]. Signed [Hoc78]. similar [Dob79]. Simple [Bro78, De 73, Lem74, Lig76, Lig77, Wes76]. Simpler [HN75]. Size [Kun74]. Skorohod [Kos74]. Skorokhod [Hal79]. SLLNs [PSW75]. Slowing [Eri77]. Slowly [Ste77a]. Small [Nor74]. Smirnov [Sen73a]. Smooth [GH73, Hel77, Hip77]. Sojourn [Tak74]. Sojourns [Ber74b, Ber80]. Solidarity [Twe74a]. Solutions [Pro77]. Solvability [Por75]. Some [AK73, CR73, Cra76, Cuz78b, DS78, Dur78, EN74, GB77, Gue74, Gut74a, Gut74b, Hol77, Kan75b, Kap74b, Ken78, Kle76, KZ79, Lan78, Les75, Mae78, O’R74a, PS76, Rob76, Ros79a, She74, Vil74, Wan75, Wic73c, Air78, Ros79b]. Source [GNO75, Kie79, PG77]. Sources [SNDL78]. Space [BW74, Bro77, Gor73, Hoc78, Jai76, Kes74, Kie75, KK74, Kue76a, Kue76b, Kuo73, McK73, NA76, Pho78, Sav75, Tra79, Twe74a, Twe74b, Yod75]. Space-time [Sav75]. Space-Valued [BW74]. Spaces [AG76, AM78, BZ79, BR79, Gar76, Har77, HJP76, KKO77, Kue75, Kum73, Ram79a, TW79, Zin77].

Spacing [PS77]. Spacings [Hol79, Kan75b]. Special [Fre74c]. Spectra [Wac78]. Spectral [Bri74, Eis76, Mar77b, O’C79]. Speed [Bol78]. Speeds [Swe77]. Spheres [HW74]. Spin [Gra78, Lig78]. Spin-Flip [Gra78]. Spreading [AG76]. Square [Cha73, Che79, HK77, Kli77, Pit79, Wri76]. Square-Integrable [HK77]. Stability [AE78, DM74, KZ79, SvH79, WW79]. Stable [Bas76, De 77, Fri79, Hen73, Hos78a, Kan75c, Kan76, Kiu75, Kum73, Min74, Moh76, Mon77, Owe73, Roo78, Rub77, Wic74a, dA75]. Standard [Jos75, Jos77]. Starting [Lev73]. State [Hal75, Har77, Kes74, NA76, Ros79d, Rub77, Sch77a, Twe74a, Twe74b, Sch78a]. Stationarity [Rip76c]. Stationary [Ber73b, BE75, Bri74, Cuz76, Dav76, Dav77, Dav79b, Deo74, Deo75, DMW74, DS75b, DS79, Dym78, GH73, GHZ76, GB77, Gre75, HS73, Hey74, Hos78b, Ivk75, Kie74, Kle76, Lai77a, LS78, MI73, Mit74, MY76, Mit79, O’B74a, PKG79, Pho78, PS77, PG77, Qua77, Sal79, Sen74b, Wri76]. Statist [Mar74b]. Statistics [CW77b, Dyn78, Hel77, JM78, JP75b, Mal73, Nai77, Sen73a, Sen74a, Sen76b, Sen76c, Sen77, Sim74]. Stay [Bol76, Igl74]. Steady [Sch77a, Sch78a]. Steady-State [Sch77a, Sch78a]. Stein [Eri74]. Step [Mye74, Sch76c]. Stepping [Saw76]. Steps [Nor74]. Stirring [Lee74]. Stochastic [Bil74b, DH79, Del77, EF75, Gar78, Gre75, Gue74, HC78b, Kal75, KKO77, KK78, Mar77a, Mil79, O’B75, Pho78, PR73, Pro77, Pro79, PG77, Rip76a, Sal79, Sus78, Wal74b, Wan75, Wic73a, Wic73c, WZ76, WZ77a, WZ77b, dHR79b, Smy75]. Stochastically [Sie76b]. Stock [GS74]. Stone [Saw76, Sza77b]. Stopped [Gre73, Leh77, Tay75a]. Stopping [Ald78, CP76, Dav78, Dub73, GM77, GS78, GS74, Kla73, Kla74, Leh77, Tei79, Wal74b, Wal74a, Bis79]. Storage [Cin75]. Strassen [Bol78, Kos74, Lai76c, Maj79, Wic73c]. Strassen-Type [Wic73c]. Strategies [DS75b, DS79, HR79]. String [Dur79]. strip [BF73]. Strong [Deo73a, Deo73b, EK74, Kud74, Kue76b, Kun74, MR75, Mor76, Num79, Rev76, Smy73, Tom76, Wac78, Wel78, Wri79, AV75, BW74, BR79, Fre73, Fre74b, Lai77a, McL75, MY76, Wel78, WPR77]. Strong-Mixing [Deo73a, Deo73b, Mor76]. Strongly [MI73]. Structure [Cin76]. Structures [Lan76]. Student [Gro76, Ism77]. Sub [Tay75b]. Sub-Markovian [Tay75b]. Subadditive [Del77, Ham74, Kin73, Smy76, Ste78a]. Subclasses [KS78]. Subcritical [CNW73, KW73, Qui76, Qui77]. Subdivision [Van78b]. Subexponential [Teu75]. Subgroups [Le 73]. Subinvariant [Twe74b]. Submartingale [Gil77]. Submartingales [SW76]. Subordinate [Gzy78]. Subordinated [Dob79]. Subordination [Rub77]. Subordinators [Fri79]. Subsequences [BE75, BR79, Gut79, Rei77]. Subterminal [Oma77]. Successes [Gle75]. Sufficient [Dyn78]. Sum [Hol79, Tho77]. Sum-Functions [Hol79]. Summation [LR77b]. Sums [CP76, CL73, Cog75, De 73, DS75a, Eri74, Gab77b, GY75, Gut74a, HM77, HGS79, Ker77, Lai74a, Lai76b, Lai77b, LS78, LMRS73, Mac78, McD79, Nag79, Ser73, She74, Sho79, Smy74, Wic74a]. Sup [O’R74b]. Sup-norm [O’R74b]. Supercritical [AK76, Fea76, Hop76, Kap74a, KN73].

Superposition [Rip76c]. Superpositions [Ser73]. Superprocesses [Lev75]. Supporting [Lan76]. Supports [Bro77]. Supremum [Der73, Kla74]. Sure [GH76, HGS79, Ker77, Lin74c, Mon76, Sen73a, Sen76a, Sen77, WW79, de 74b]. Survey [GS79]. Survival [HL78, Sav76]. Symmetric [Imh76, Mal74a, Pit77, Sil74, Wri73]. System [Kal78, Lig73, Lig74, Sch76b]. Systems [Bri75, CMSB75, Dyn77, Fel77, Gra78, Gzy77, HL75, HS76, Lig78, Mai75, PSW75, Pro77, Ros79d].

Tail [Fre75, HJSD79, Ros79c, Wes79, Wri73]. Tailfree [Dok74]. Tails [Rud73, WPR77]. Tauberian [DS78, Lai76d, Sen73b]. Taylor [Hwa76]. Technique [Eri74]. Temps [Bis79]. Term [Ben75, O’B74a]. Terms [Mal73, Mye76]. Test [JP75b]. Tests [LS78]. Teugels [Moh76]. th [HN75, WN73]. Their [Dok74, Dyn77, Hos78a, Ven79, Dob79, Lai76d, Lyn79]. Theorem [Bas76, BR77, Cha76b, CW77b, Che75a, CHL79, Cox79a, Dar75, Deo73c, DS76, Eri73c, HR79, Hen73, Hwa77, Kai75a, Kai75b, Kaw79, Ker78, Kie73, Kie75, Kle76, Kue76b, Lin77, Lyn78, Mac73, McD79, Mey73, Min74, Moh76, Oma77, Rei77, Rya76, Sen73b, Ser75, Sho79, Sho78, Sin75, Tei73, Tho77, Wel78, Wel73, Woo76, Wri76, Nor74, Afg75, Bas74, Ber73a, Bha77, Bol76, Cuz76, Dar76, Deo75, Gor73, Hah77, HC78a, HJP76, HT79, Kie76a, M.74, Mae78, Mat75, Sen76c, Swe77, Won79, Zin77]. Theorems [AE78, AMN78, Ben75, DO75, DS78, Dud79, Dun77, Dur78, DR78, Gis73, Gri78, GOR79, Gue74, HL75, Kap75, Ker74, Lai74a, Lai76d, LS73, Lig77, Nea78, Num79, O’B74a, PG77, Sch76b, Tan77, Tom76, Tor78, Vil74, Eth79, Dud78, Gar76, Igl74, McL74, PS75]. Théorie [Air78, Alv76, GM73]. Theory [AMN78, Bil74c, Cha74b, CHL79, Chu73, Cog75, Cra76, Eis76, GNS75, Kes74, Kin73, KH73, Lai76a, Lai76d, LR77b, Lem74, Mar77b, Orn73b, Rip76b, Twe74a, Twe74b, Wei75b, WR78, dJS79, Air78, Alv76, GM73]. Thinned [Wes76]. Thorems [BP79]. Tight [TW79]. Tightness [Ald78, Bil74a, Loy76]. Time [BdH74, Cha74a, Cha77, Cin75, Cox77b, Cox79b, Ete76, Gre75, Gzy77, Hos78a, Kni73, Kni75, LW79, Lev73, Lin75, Mon77, Ros79c, SC74, Tak74, Tei79, Wan79, Wes79, Bis79, PG77, Sav75]. Time-Changed [Gzy77]. Time-dependent [Wan79]. Time-Homogeneous [Ros79c]. Times [Ald78, Cha77, Dav76, Dav78, Dub73, Eth79, Gem73, GH73, GS73, GS78, Gut74a, Gut74b, Lai77b, Leh77, Mar73c, Oma77, Rob76, Ros73, SB75, Woo76]. Timid [GS77]. Topological [Sie76a]. Topology [KN79]. Tossing [DS78]. Total [Kan75c, Kie76b, Lig74]. Transformation [VW74]. Transformations [Kun74]. Transforms [Bur79]. Transience [Por75]. Transient [Twe74b]. Transition [Cog75]. Transitions [KD79]. Translated [MS75]. Translates [Hud76, Kan75a]. Tree [Spi75]. Trees [Mon75]. Trials [Che75b, Gle75]. Two [Afg75, Ahl74, Ete76, Fla76, Goo76, Kap77, Kiu75, PS75, Sen76c, She74, Sza77a, Win73, Pra75]. Two-Dimensional [Ete76, Fla76, Sen76c, She74]. Two-Parameter [Goo76]. Two-sample [PS75]. Type [AM78, Mey73, Wic73c].

Unbounded [Ber73a]. Undominated [Lan76]. Unexpected [Sim77]. Uniform [Adl78a, Adl80, Bru75, Cog75, Fri79, KP78, Lai76d, Mye74, Rog74, Dia77]. Uniformity [Sza77b]. Uniformly [VP76]. Unimodal [O’C79]. Unimodality [Kan76, Yam78]. Uniqueness [GG77, Kes76, Pro77]. Unit [Won79]. Unitary [Dun77]. Univariate [Bas74]. Unknown [Kie79]. Upper [Dal76, Dal78, JJS75, Jos77, LS78]. Upper-Lower [LS78]. Uses [Kin78]. Using [Eri74].

Vague [EF75]. Value [Gil77, Ker74, PS79, Wei75b]. Valued [BW74, Gri78, Har76, Har78, Jam76, Kla73, Kue76a, Kue76b, Pho78, Res74]. Values [DO75, KZ79, Man76, Nai77, Res73, SB75, Wel73]. Variability [GS75]. Variable [Bar75]. Variables [Adl78b, Bas76, BW74, Ber73a, CP76, CL73, DR78, Eri73c, Gab77b, GS79, GY75, GPW73, Gut78, HGS79, Jam76, Jof74, Ker77, KT77, Kue76a, Kue76b, KZ79, Lai76a, Lai77b, Let75, Man76, McD79, Nag79, Nea78, O’B74b, Owe73, Phi77, Sch74, She74, Sho78, Sin75, Smy73, Smy74, Ste77b, Tom76, Van78a, Vil74, VP76, Wal74b, Wal74a, Wei75a, Wic74b, Wic74a, Wri73, WPR77, Wri79, Mar74b]. Variance [BF73, DM78, NR75]. Variation [GM77, Gru75, Kan75c, KG75, Kni73, Mon76, Wan77, dHR79a, de 74b]. Varying [Lin74c]. Vector [BZ79, Cuz78b, KZ79, Res74, Ven79]. Vector-valued [Res74]. Vectors [BP79]. Velocities [Kal78]. Version [Eri73c]. Versions [Bas74, Lai77a, Mae78, Zab79]. Via [Dal78]. View [Kni75]. Volume [Ano73a, Ano73b, Ano73c, Ano73d, Ano73e, Ano73f, Ano73g, Ano73h, Ano73i, Ano73j, Ano73k, Ano73l, Ano73n, Ano74a, Ano74b, Ano74c, Ano74d, Ano74e, Ano74f, Ano74g, Ano74h, Ano74i, Ano74j, Ano74k, Ano74l, Ano74m, Ano74n, Ano75a, Ano75b, Ano75c, Ano75d, Ano75e, Ano75f, Ano75g, Ano75h, Ano75i, Ano75j, Ano75k, Ano75l, Ano75m, Ano76a, Ano76b, Ano76c, Ano76d, Ano76e, Ano76f, Ano76g, Ano76h, Ano76i, Ano76j, Ano76k, Ano76l, Ano76m, Ano77a, Ano77b, Ano77c, Ano77d, Ano77e, Ano77f, Ano77g, Ano77h, Ano77i, Ano77j, Ano77k, Ano77l, Ano77m, Ano78a, Ano78b, Ano78c, Ano78d, Ano78e, Ano78f, Ano78g, Ano78h, Ano78i, Ano78j, Ano78k, Ano78l, Ano78m, Ano79a, Ano79b, Ano79c, Ano79d, Ano79e, Ano79f]. Volume [Ano79g, Ano79h, Ano79i, Ano79j, Ano79k, Ano79l, Ano79m]. Voter [BG79, HL75]. Voting [Mat77].

W. [Sen73b]. Waiting [SC74]. Walk [Bla76, EK74, Eri76, Ete76, Fla76, Fra73, GS78, Imh76, Kai75a, Kai76, LW79, Lem74, Mye74]. Walks [Bol76, DM78, Eri77, FP75, Igl74, Lai76a, Mye76, Sol75, Tei79]. Watson [DS76, Hop77, Kap73]. Weak [Adl78b, Ber76, BS77a, Deo73c, DIM77, EK74, Fis75, Ker78, KN79, Kus74, LdMR75, MR75, O’R74b, Res75, Res74, Sen74a, Sen74b, Ser73, SS78, Wei76, WZ76, PS75]. Weakly [BW74, BP79, DMW74, HL75, Phi77]. Wear [EM73]. Weight [Hel77].

Weighted [CL73, Jam75, Sin75, Tom75a, WPR77]. Wendel [De 73]. Which [Dub73, DS75b, Kal77, Roo76]. Whose [Wri73]. Wide [Wri76]. Wide-sense [Wri76]. Wiener [CR79, Dud77, Goo76, Hoc78, HC78b, OP73, Tra77]. Wilcoxon [Tho77]. William [Chu73]. Without [Kal76, KT77, Tei73, Kal73, Kal74, Kun74, Nev76]. Wold [Ivk75].

Yaglom [KW73].

Zero [AP79, CR73, Cuz75, Dav77, GPW73, Kai76, Lig73, Mon78a, Sen75, ST75, Win73]. Zero-One [CR73, Sen75, AP79]. Zero-Two [Win73]. Zeros [Cuz76, Cuz78a, Cuz87].

References

Adler:1977:HDG

[Adl77] Robert J. Adler. Hausdorff dimension and Gaussian fields. Annals of Probability, 5(1):145–151, February 1977. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http:

//projecteuclid.org/euclid.aop/1176995900.

Adler:1978:UDL

[Adl78a] Robert J. Adler. The uniform dimension of the level sets of a Brownian sheet. Annals of Probability, 6(3):509–515, June 1978. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176995535. See correc-

tion [Adl80].

Adler:1978:WCR

[Adl78b] Robert J. Adler. Weak convergence results for extremal processes generated by dependent random variables. Annals of Probability, 6(4): 660–667, August 1978. CODEN APBYAE. ISSN 0091-1798 (print),

2168-894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176995486.

Adler:1980:CNC

[Adl80] Robert J. Adler. Correction notes: Correction to “The Uniform Dimension of the Level Sets of a Brownian Sheet”. Annals of Probability, 8(5):1001–1002, October 1980. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.org/

euclid.aop/1176994630. See [Adl78a].

Aldous:1978:MSL

[AE78] D. J. Aldous and G. K. Eagleson. On mixing and stability of limit theorems. Annals of Probability, 6(2):325–331, April 1978. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http:

//projecteuclid.org/euclid.aop/1176995577.

Afghahi:1975:RCC

[Afg75] M. H. Afghahi. On the rate of convergence in the Central Limit Theorem in two dimensions and its application. Annals of Probability, 3(5):802–814, October 1975. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.org/

euclid.aop/1176996267.

Ahlswede:1976:SSP

[AG76] Rudolf Ahlswede and Peter Gacs. Spreading of sets in product spaces and hypercontraction of the Markov operator. Annals of Probability, 4(6):925–939, December 1976. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.org/

euclid.aop/1176995937.

Adler:1976:LCR

[AH76] Robert J. Adler and A. M. Hasofer. Level crossings for random fields. Annals of Probability, 4(1):1–12, February 1976. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http:

//projecteuclid.org/euclid.aop/1176996176.

Ahlswede:1974:CRC

[Ahl74] R. Ahlswede. The capacity region of a channel with two senders and two receivers. Annals of Probability, 2(5):805–814, October 1974. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176996549.

Airault:1978:QEE

[Air78] Hélène Airault. Quelques ensembles exceptionnels en théorie du potentiel. (French) [Some exceptional sets in potential theory]. Annals of Probability, 6(2):316–320, April 1978. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176995575.

Athreya:1973:GDI

[AK73] Krishna B. Athreya and Thomas G. Kurtz. A generalization of Dynkin’s identity and some applications. Annals of Probability, 1(4):

570–579, August 1973. CODEN APBYAE. ISSN 0091-1798 (print),

2168-894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176996886.

Athreya:1976:CAD

[AK76] K. B. Athreya and N. Kaplan. Convergence of the age distribution in the one-dimensional supercritical age-dependent branching process. Annals of Probability, 4(1):38–50, February 1976. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http:

//projecteuclid.org/euclid.aop/1176996179.

Aldous:1978:STT

[Ald78] David Aldous. Stopping times and tightness. Annals of Probability, 6 (2):335–340, April 1978. CODEN APBYAE. ISSN 0091-1798 (print),

2168-894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176995579.

Alvarez:1976:TDE

[Alv76] M. A. Garcia Alvarez. Une théorie de la dualité à ensemble polaire près II. (French) [[A duality theory for the near polar ensemble II]. Annals of Probability, 4(6):947–976, December 1976. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http:

//projecteuclid.org/euclid.aop/1176995939.

Araujo:1978:TCL

[AM78] Aloisio Araujo and Evarist Gine M. Type, cotype and Lévy measures in Banach spaces. Annals of Probability, 6(4):637–643, August 1978. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176995483.

Athreya:1978:LTS

[AMN78] K. B. Athreya, D. McDonald, and P. Ney. Limit theorems for semi-Markov processes and renewal theory for Markov chains. Annals of Probability, 6(5):788–797, October 1978. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176995429.

Athreya:1974:FCM

[AN74] K. Athreya and P. Ney. Functionals of critical multitype branching processes. Annals of Probability, 2(2):339–343, April 1974. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176996716.

Anonymous:1973:MBPa

[Ano73a] Anonymous. Miscellaneous back pages, Ann. Probab., Volume 1, Number 1 (1973). Annals of Probability, 1(1):??, February 1973. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176997040.

Anonymous:1973:MBPb

[Ano73b] Anonymous. Miscellaneous back pages, Ann. Probab., Volume 1, Number 2 (1973). Annals of Probability, 1(2):??, April 1973. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176996995.

Anonymous:1973:MBPc

[Ano73c] Anonymous. Miscellaneous back pages, Ann. Probab., Volume 1, Number 3 (1973). Annals of Probability, 1(3):??, June 1973. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176996949.

Anonymous:1973:MBPd

[Ano73d] Anonymous. Miscellaneous back pages, Ann. Probab., Volume 1, Number 4 (1973). Annals of Probability, 1(4):??, August 1973. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176996901.

Anonymous:1973:MBPe

[Ano73e] Anonymous. Miscellaneous back pages, Ann. Probab., Volume 1, Number 5 (1973). Annals of Probability, 1(5):??, October 1973. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176996857.

Anonymous:1973:MBPf

[Ano73f] Anonymous. Miscellaneous back pages, Ann. Probab., Volume 1, Number 6 (1973). Annals of Probability, 1(6):??, December 1973. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176996817.

Anonymous:1973:MFPa

[Ano73g] Anonymous. Miscellaneous front pages, Ann. Probab., Volume 1, Number 1 (1973). Annals of Probability, 1(1):??, February 1973. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176997020.

Anonymous:1973:MFPb

[Ano73h] Anonymous. Miscellaneous front pages, Ann. Probab., Volume 1, Number 2 (1973). Annals of Probability, 1(2):??, April 1973. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176996972.

Anonymous:1973:MFPc

[Ano73i] Anonymous. Miscellaneous front pages, Ann. Probab., Volume 1, Number 3 (1973). Annals of Probability, 1(3):??, June 1973. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176996931.

Anonymous:1973:MFPd

[Ano73j] Anonymous. Miscellaneous front pages, Ann. Probab., Volume 1, Number 4 (1973). Annals of Probability, 1(4):??, August 1973. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176996883.

Anonymous:1973:MFPe

[Ano73k] Anonymous. Miscellaneous front pages, Ann. Probab., Volume 1, Number 5 (1973). Annals of Probability, 1(5):??, October 1973. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176996842.

Anonymous:1973:MFPf

[Ano73l] Anonymous. Miscellaneous front pages, Ann. Probab., Volume 1, Number 6 (1973). Annals of Probability, 1(6):??, December 1973. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176996797.

Anonymous:1973:PPL

[Ano73m] Anonymous. Publications of Paul Lévy. Annals of Probability, 1(1): 9–18, February 1973. CODEN APBYAE. ISSN 0091-1798 (print),

2168-894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176997022.

Anonymous:1973:VI

[Ano73n] Anonymous. Volume information. Annals of Probability, 1(6):??, December 1973. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.aop/
1176996796.

Anonymous:1974:MBPa

[Ano74a] Anonymous. Miscellaneous back pages, Ann. Probab., Volume 2, Number 1 (1974). Annals of Probability, 2(1):??, February 1974. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176996769.

Anonymous:1974:MBPb

[Ano74b] Anonymous. Miscellaneous back pages, Ann. Probab., Volume 2, Number 2 (1974). Annals of Probability, 2(2):??, April 1974. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176996720.

Anonymous:1974:MBPc

[Ano74c] Anonymous. Miscellaneous back pages, Ann. Probab., Volume 2, Number 3 (1974). Annals of Probability, 2(3):??, June 1974. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176996676.

Anonymous:1974:MBPd

[Ano74d] Anonymous. Miscellaneous back pages, Ann. Probab., Volume 2, Number 4 (1974). Annals of Probability, 2(4):??, August 1974. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176996621.

Anonymous:1974:MBPe

[Ano74e] Anonymous. Miscellaneous back pages, Ann. Probab., Volume 2, Number 5 (1974). Annals of Probability, 2(5):??, October 1974. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176996565.

Anonymous:1974:MBPf

[Ano74f] Anonymous. Miscellaneous back pages, Ann. Probab., Volume 2, Number 6 (1974). Annals of Probability, 2(6):??, December 1974. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176996508.

Anonymous:1974:MFPa

[Ano74g] Anonymous. Miscellaneous front pages, Ann. Probab., Volume 2, Number 1 (1974). Annals of Probability, 2(1):??, February 1974. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176996746.

Anonymous:1974:MFPb

[Ano74h] Anonymous. Miscellaneous front pages, Ann. Probab., Volume 2, Number 2 (1974). Annals of Probability, 2(2):??, April 1974. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176996702.

Anonymous:1974:MFPc

[Ano74i] Anonymous. Miscellaneous front pages, Ann. Probab., Volume 2, Number 3 (1974). Annals of Probability, 2(3):??, June 1974. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176996653.

Anonymous:1974:MFPd

[Ano74j] Anonymous. Miscellaneous front pages, Ann. Probab., Volume 2, Number 4 (1974). Annals of Probability, 2(4):??, August 1974. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176996603.

Anonymous:1974:MFPe

[Ano74k] Anonymous. Miscellaneous front pages, Ann. Probab., Volume 2, Number 5 (1974). Annals of Probability, 2(5):??, October 1974. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176996546.

Anonymous:1974:MFPf

[Ano74l] Anonymous. Miscellaneous front pages, Ann. Probab., Volume 2, Number 6 (1974). Annals of Probability, 2(6):??, December 1974. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176996492.

Anonymous:1974:NAB

[Ano74m] Anonymous. Notes: Addendum to “The Annals of Probability, Contents of Volume 1 1973.”. Annals of Probability, 2(4):747, August 1974. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176996620.

Anonymous:1974:VI

[Ano74n] Anonymous. Volume information. Annals of Probability, 2(6):??, December 1974. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.aop/
1176996491.

Anonymous:1975:MBPa

[Ano75a] Anonymous. Miscellaneous back pages, Ann. Probab., Volume 3, Number 1 (1975). Annals of Probability, 3(1):??, February 1975. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176996463.

Anonymous:1975:MBPb

[Ano75b] Anonymous. Miscellaneous back pages, Ann. Probab., Volume 3, Number 2 (1975). Annals of Probability, 3(2):??, April 1975. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176996413.

Anonymous:1975:MBPc

[Ano75c] Anonymous. Miscellaneous back pages, Ann. Probab., Volume 3, Number 3 (1975). Annals of Probability, 3(3):??, June 1975. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176996365.

Anonymous:1975:MBPd

[Ano75d] Anonymous. Miscellaneous back pages, Ann. Probab., Volume 3, Number 4 (1975). Annals of Probability, 3(4):??, August 1975. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176996317.

Anonymous:1975:MBPe

[Ano75e] Anonymous. Miscellaneous back pages, Ann. Probab., Volume 3, Number 5 (1975). Annals of Probability, 3(5):??, October 1975. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176996281.

Anonymous:1975:MBPf

[Ano75f] Anonymous. Miscellaneous back pages, Ann. Probab., Volume 3, Number 6 (1975). Annals of Probability, 3(6):??, December 1975. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176996238.

Anonymous:1975:MFPa

[Ano75g] Anonymous. Miscellaneous front pages, Ann. Probab., Volume 3, Number 1 (1975). Annals of Probability, 3(1):??, February 1975. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176996443.

Anonymous:1975:MFPb

[Ano75h] Anonymous. Miscellaneous front pages, Ann. Probab., Volume 3, Number 2 (1975). Annals of Probability, 3(2):??, April 1975. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176996392.

Anonymous:1975:MFPc

[Ano75i] Anonymous. Miscellaneous front pages, Ann. Probab., Volume 3, Number 3 (1975). Annals of Probability, 3(3):??, June 1975. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176996346.

Anonymous:1975:MFPd

[Ano75j] Anonymous. Miscellaneous front pages, Ann. Probab., Volume 3, Number 4 (1975). Annals of Probability, 3(4):??, August 1975. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176996301.

Anonymous:1975:MFPe

[Ano75k] Anonymous. Miscellaneous front pages, Ann. Probab., Volume 3, Number 5 (1975). Annals of Probability, 3(5):??, October 1975. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176996260.

Anonymous:1975:MFPf

[Ano75l] Anonymous. Miscellaneous front pages, Ann. Probab., Volume 3, Number 6 (1975). Annals of Probability, 3(6):??, December 1975. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176996217.

Anonymous:1975:VI

[Ano75m] Anonymous. Volume information. Annals of Probability, 3(6):??, December 1975. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.aop/
1176996216.

Anonymous:1976:MBPa

[Ano76a] Anonymous. Miscellaneous back pages, Ann. Probab., Volume 4, Number 1 (1976). Annals of Probability, 4(1):??, February 1976. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176996197.

Anonymous:1976:MBPb

[Ano76b] Anonymous. Miscellaneous back pages, Ann. Probab., Volume 4, Number 2 (1976). Annals of Probability, 4(2):??, April 1976. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176996141.

Anonymous:1976:MBPc

[Ano76c] Anonymous. Miscellaneous back pages, Ann. Probab., Volume 4, Number 3 (1976). Annals of Probability, 4(3):??, June 1976. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176996104.

Anonymous:1976:MBPd

[Ano76d] Anonymous. Miscellaneous back pages, Ann. Probab., Volume 4, Number 4 (1976). Annals of Probability, 4(4):??, August 1976. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176996043.

Anonymous:1976:MBPe

[Ano76e] Anonymous. Miscellaneous back pages, Ann. Probab., Volume 4, Number 5 (1976). Annals of Probability, 4(5):??, October 1976. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176995993.

Anonymous:1976:MBPf

[Ano76f] Anonymous. Miscellaneous back pages, Ann. Probab., Volume 4, Number 6 (1976). Annals of Probability, 4(6):??, December 1976. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176995951.

Anonymous:1976:MFPa

[Ano76g] Anonymous. Miscellaneous front pages, Ann. Probab., Volume 4, Number 1 (1976). Annals of Probability, 4(1):??, February 1976. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176996175.

Anonymous:1976:MFPb

[Ano76h] Anonymous. Miscellaneous front pages, Ann. Probab., Volume 4, Number 2 (1976). Annals of Probability, 4(2):??, April 1976. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176996127.

Anonymous:1976:MFPc

[Ano76i] Anonymous. Miscellaneous front pages, Ann. Probab., Volume 4, Number 3 (1976). Annals of Probability, 4(3):??, June 1976. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176996083.

Anonymous:1976:MFPd

[Ano76j] Anonymous. Miscellaneous front pages, Ann. Probab., Volume 4, Number 4 (1976). Annals of Probability, 4(4):??, August 1976. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176996024.

Anonymous:1976:MFPe

[Ano76k] Anonymous. Miscellaneous front pages, Ann. Probab., Volume 4, Number 5 (1976). Annals of Probability, 4(5):??, October 1976. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176995979.

Anonymous:1976:MFPf

[Ano76l] Anonymous. Miscellaneous front pages, Ann. Probab., Volume 4, Number 6 (1976). Annals of Probability, 4(6):??, December 1976. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176995932.

Anonymous:1976:VI

[Ano76m] Anonymous. Volume information. Annals of Probability, 4(6):??, December 1976. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.aop/
1176995931.

Anonymous:1977:MBPa

[Ano77a] Anonymous. Miscellaneous back pages, Ann. Probab., Volume 5, Number 1 (1977). Annals of Probability, 5(1):??, February 1977. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176995904.

Anonymous:1977:MBPb

[Ano77b] Anonymous. Miscellaneous back pages, Ann. Probab., Volume 5, Number 2 (1977). Annals of Probability, 5(2):??, April 1977. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176995860.

Anonymous:1977:MBPc

[Ano77c] Anonymous. Miscellaneous back pages, Ann. Probab., Volume 5, Number 3 (1977). Annals of Probability, 5(3):??, June 1977. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176995814.

Anonymous:1977:MBPd

[Ano77d] Anonymous. Miscellaneous back pages, Ann. Probab., Volume 5, Number 4 (1977). Annals of Probability, 5(4):??, August 1977. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176995777.

Anonymous:1977:MBPe

[Ano77e] Anonymous. Miscellaneous back pages, Ann. Probab., Volume 5, Number 5 (1977). Annals of Probability, 5(5):??, October 1977. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176995729.

Anonymous:1977:MBPf

[Ano77f] Anonymous. Miscellaneous back pages, Ann. Probab., Volume 5, Number 6 (1977). Annals of Probability, 5(6):??, December 1977. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176995675.

Anonymous:1977:MFPa

[Ano77g] Anonymous. Miscellaneous front pages, Ann. Probab., Volume 5, Number 1 (1977). Annals of Probability, 5(1):??, February 1977. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176995886.

Anonymous:1977:MFPb

[Ano77h] Anonymous. Miscellaneous front pages, Ann. Probab., Volume 5, Number 2 (1977). Annals of Probability, 5(2):??, April 1977. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176995842.

Anonymous:1977:MFPc

[Ano77i] Anonymous. Miscellaneous front pages, Ann. Probab., Volume 5, Number 3 (1977). Annals of Probability, 5(3):??, June 1977. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176995794.

Anonymous:1977:MFPd

[Ano77j] Anonymous. Miscellaneous front pages, Ann. Probab., Volume 5, Number 4 (1977). Annals of Probability, 5(4):??, August 1977. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176995756.

Anonymous:1977:MFPe

[Ano77k] Anonymous. Miscellaneous front pages, Ann. Probab., Volume 5, Number 5 (1977). Annals of Probability, 5(5):??, October 1977. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176995710.

Anonymous:1977:MFPf

[Ano77l] Anonymous. Miscellaneous front pages, Ann. Probab., Volume 5, Number 6 (1977). Annals of Probability, 5(6):??, December 1977. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176995654.

Anonymous:1977:VI

[Ano77m] Anonymous. Volume information. Annals of Probability, 5(6):??, December 1977. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.aop/
1176995653.

Anonymous:1978:MBPa

[Ano78a] Anonymous. Miscellaneous back pages, Ann. Probab., Volume 6, Number 1 (1978). Annals of Probability, 6(1):??, February 1978. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176995625.

Anonymous:1978:MBPb

[Ano78b] Anonymous. Miscellaneous back pages, Ann. Probab., Volume 6, Number 2 (1978). Annals of Probability, 6(2):??, April 1978. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176995583.

Anonymous:1978:MBPc

[Ano78c] Anonymous. Miscellaneous back pages, Ann. Probab., Volume 6, Number 3 (1978). Annals of Probability, 6(3):??, June 1978. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176995537.

Anonymous:1978:MBPd

[Ano78d] Anonymous. Miscellaneous back pages, Ann. Probab., Volume 6, Number 4 (1978). Annals of Probability, 6(4):??, August 1978. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176995494.

Anonymous:1978:MBPe

[Ano78e] Anonymous. Miscellaneous back pages, Ann. Probab., Volume 6, Number 5 (1978). Annals of Probability, 6(5):??, October 1978. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176995437.

Anonymous:1978:MBPf

[Ano78f] Anonymous. Miscellaneous back pages, Ann. Probab., Volume 6, Number 6 (1978). Annals of Probability, 6(6):??, December 1978. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176995393.

Anonymous:1978:MFPa

[Ano78g] Anonymous. Miscellaneous front pages, Ann. Probab., Volume 6, Number 1 (1978). Annals of Probability, 6(1):??, February 1978. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176995606.

Anonymous:1978:MFPb

[Ano78h] Anonymous. Miscellaneous front pages, Ann. Probab., Volume 6, Number 2 (1978). Annals of Probability, 6(2):??, April 1978. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176995565.

Anonymous:1978:MFPc

[Ano78i] Anonymous. Miscellaneous front pages, Ann. Probab., Volume 6, Number 3 (1978). Annals of Probability, 6(3):??, June 1978. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176995522.

Anonymous:1978:MFPd

[Ano78j] Anonymous. Miscellaneous front pages, Ann. Probab., Volume 6, Number 4 (1978). Annals of Probability, 6(4):??, August 1978. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176995473.

Anonymous:1978:MFPe

[Ano78k] Anonymous. Miscellaneous front pages, Ann. Probab., Volume 6, Number 5 (1978). Annals of Probability, 6(5):??, October 1978. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176995423.

Anonymous:1978:MFPf

[Ano78l] Anonymous. Miscellaneous front pages, Ann. Probab., Volume 6, Number 6 (1978). Annals of Probability, 6(6):??, December 1978. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176995383.

Anonymous:1978:VI

[Ano78m] Anonymous. Volume information. Annals of Probability, 6(6):??, December 1978. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.aop/
1176995382.

Anonymous:1979:MBPa

[Ano79a] Anonymous. Miscellaneous back pages, Ann. Probab., Volume 7, Number 1 (1979). Annals of Probability, 7(1):??, February 1979. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176995165.

Anonymous:1979:MBPb

[Ano79b] Anonymous. Miscellaneous back pages, Ann. Probab., Volume 7, Number 2 (1979). Annals of Probability, 7(2):??, April 1979. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176995097.

Anonymous:1979:MBPc

[Ano79c] Anonymous. Miscellaneous back pages, Ann. Probab., Volume 7, Number 3 (1979). Annals of Probability, 7(3):??, June 1979. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176995057.

Anonymous:1979:MBPd

[Ano79d] Anonymous. Miscellaneous back pages, Ann. Probab., Volume 7, Number 4 (1979). Annals of Probability, 7(4):??, August 1979. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176994996.

Anonymous:1979:MBPe

[Ano79e] Anonymous. Miscellaneous back pages, Ann. Probab., Volume 7, Number 5 (1979). Annals of Probability, 7(5):??, October 1979. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176994954.

Anonymous:1979:MBPf

[Ano79f] Anonymous. Miscellaneous back pages, Ann. Probab., Volume 7, Number 6 (1979). Annals of Probability, 7(6):??, December 1979. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176994909.

Anonymous:1979:MFPa

[Ano79g] Anonymous. Miscellaneous front pages, Ann. Probab., Volume 7, Number 1 (1979). Annals of Probability, 7(1):??, February 1979. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176995144.

Anonymous:1979:MFPb

[Ano79h] Anonymous. Miscellaneous front pages, Ann. Probab., Volume 7, Number 2 (1979). Annals of Probability, 7(2):??, April 1979. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176995084.

Anonymous:1979:MFPc

[Ano79i] Anonymous. Miscellaneous front pages, Ann. Probab., Volume 7, Number 3 (1979). Annals of Probability, 7(3):??, June 1979. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176995040.

Anonymous:1979:MFPd

[Ano79j] Anonymous. Miscellaneous front pages, Ann. Probab., Volume 7, Number 4 (1979). Annals of Probability, 7(4):??, August 1979. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176994983.

Anonymous:1979:MFPe

[Ano79k] Anonymous. Miscellaneous front pages, Ann. Probab., Volume 7, Number 5 (1979). Annals of Probability, 7(5):??, October 1979. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176994937.

Anonymous:1979:MFPf

[Ano79l] Anonymous. Miscellaneous front pages, Ann. Probab., Volume 7, Number 6 (1979). Annals of Probability, 7(6):??, December 1979. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176994887.

Anonymous:1979:VI

[Ano79m] Anonymous. Volume information. Annals of Probability, 7(6):??, December 1979. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.aop/
1176994886.

Aldous:1979:ZOL

[AP79] David Aldous and Jim Pitman. On the zero-one law for exchangeable events. Annals of Probability, 7(4):704–723, August 1979. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176994992.

Asmussen:1976:CRB

[Asm76] Soren Asmussen. Convergence rates for branching processes. Annals of Probability, 4(1):139–146, February 1976. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http:

//projecteuclid.org/euclid.aop/1176996193.

Astbury:1978:AID

[Ast78] Kenneth A. Astbury. Amarts indexed by directed sets. Annals of Probability, 6(2):267–278, April 1978. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176995572.

Ammann:1978:RMA

[AT78] Larry P. Ammann and Peter F. Thall. Random measures with aftereffects. Annals of Probability, 6(2):216–230, April 1978. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http:

//projecteuclid.org/euclid.aop/1176995569.

Artstein:1975:SLL

[AV75] Zvi Artstein and Richard A. Vitale. A Strong Law of Large Numbers for random compact sets. Annals of Probability, 3(5):879–882, October 1975. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.aop/
1176996275.

Baker:1973:EPM

[Bak73] Charles R. Baker. On equivalence of probability measures. Annals of Probability, 1(4):690–698, August 1973. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176996895.

Barlow:1975:CPR

[Bar75] William J. Barlow. Coefficient properties of random variable sequences. Annals of Probability, 3(5):840–848, October 1975. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176996270.

Basu:1974:DVU

[Bas74] Sujit K. Basu. Density versions of the univariate Central Limit Theorem. Annals of Probability, 2(2):270–276, April 1974. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http:

//projecteuclid.org/euclid.aop/1176996708.

Basu:1976:LLT

[Bas76] Sujit K. Basu. On a local limit theorem concerning variables in the domain of normal attraction of a stable law of index α, 1 < α < 2. Annals of Probability, 4(3):486–489, June 1976. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http:

//projecteuclid.org/euclid.aop/1176996099.

Burke:1979:AEP

[BCCR79] M. D. Burke, M. Csörgő, S. Csörgő, and P. Revesz. Approximations of the empirical process when parameters are estimated. Annals of Probability, 7(5):790–810, October 1979. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http:

//projecteuclid.org/euclid.aop/1176994939.

Blackwell:1975:ENE

[BD75] David Blackwell and Lester E. Dubins. On existence and non-existence of proper, regular, conditional distributions. Annals of Probability, 3(5):741–752, October 1975. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.org/

euclid.aop/1176996261.

Balkema:1974:RLT

[BdH74] A. A. Balkema and L. de Haan. Residual life time at great age. Annals of Probability, 2(5):792–804, October 1974. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http:

//projecteuclid.org/euclid.aop/1176996548.

Burkholder:1973:DPK

[BDK^{+}73] D. L. Burkholder, Daryl Daley, H. Kesten, P. Ney, Frank Spitzer, J. M. Hammersley, and J. F. C. Kingman. Discussion on Professor Kingman’s paper. Annals of Probability, 1(6):900–909, December 1973. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.aop/
1176996799.

Blum:1973:NRM

[BE73] Julius Blum and Bennett Eisenberg. A note on random measures and moving averages on non-discrete groups. Annals of Probability, 1(2): 336–337, April 1973. CODEN APBYAE. ISSN 0091-1798 (print),

2168-894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176996987.

Blum:1975:LLN

[BE75] Julius Blum and Bennett Eisenberg. The Law of Large Numbers for subsequences of a stationary process. Annals of Probability, 3(2): 281–288, April 1975. CODEN APBYAE. ISSN 0091-1798 (print),

2168-894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176996398.

Benedicks:1975:EMC

[Ben75] Michael Benedicks. An estimate of the modulus of the characteristic function of a lattice distribution with application to remainder term estimates in local limit theorems. Annals of Probability, 3(1):162– 165, February 1975. CODEN APBYAE. ISSN 0091-1798 (print),

2168-894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176996456.

Berk:1973:CLT

[Ber73a] Kenneth N. Berk. A Central Limit Theorem for m-dependent random variables with unbounded m. Annals of Probability, 1(2):352– 354, April 1973. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.aop/
1176996992.

Berman:1973:ESG

[Ber73b] Simeon M. Berman. Excursions of stationary Gaussian processes above high moving barriers. Annals of Probability, 1(3):365–387, June 1973. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.aop/
1176996932.

Berman:1974:GPD

[Ber74a] Simeon M. Berman. A Gaussian paradox: Determinism and discontinuity of sample functions. Annals of Probability, 2(5):950–953, October 1974. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.aop/
1176996560.

Berman:1974:SEG

[Ber74b] Simeon M. Berman. Sojourns and extremes of Gaussian processes. Annals of Probability, 2(6):999–1026, December 1974. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176996495. See correc-

tion [Ber80].

Berman:1976:EWM

[Ber76] Jeremy Berman. An example of a weak martingale. Annals of Probability, 4(1):107–108, February 1976. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176996187.

Berning:1979:MLI

[Ber79] John A. Berning. On the multivariate law of the iterated logarithm. Annals of Probability, 7(6):980–988, December 1979. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http:

//projecteuclid.org/euclid.aop/1176994891.

Berman:1980:CNC

[Ber80] Simeon M. Berman. Correction notes: Correction to “Sojourns and Extremes of Gaussian Processes”. Annals of Probability, 8(5):999, October 1980. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.aop/

1176994628. See [Ber74b].

Blackwell:1973:AVN

[BF73] David Blackwell and David Freedman. On the amount of variance needed to escape from a strip. Annals of Probability, 1(5):772–787, October 1973. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.aop/
1176996845.

Blackwell:1974:ORO

[BFO74] D. Blackwell, D. Freedman, and M. Orkin. The optimal reward operator in dynamic programming. Annals of Probability, 2(5):926–941, October 1974. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.aop/
1176996558.

Bramson:1979:RDV

[BG79] Maury Bramson and David Griffeath. Renormalizing the 3dimensional voter model. Annals of Probability, 7(3):418–432, June 1979. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.aop/
1176995043.

Bhattacharya:1975:ENA

[Bha75] R. N. Bhattacharya. On errors of normal approximation. Annals of Probability, 3(5):815–828, October 1975. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176996268.

Bhattacharya:1977:RMC

[Bha77] R. N. Bhattacharya. Refinements of the multidimensional Central Limit Theorem and applications. Annals of Probability, 5(1):1–27, February 1977. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.aop/
1176995887.

Bhattacharya:1978:CRE

[Bha78] R. N. Bhattacharya. Criteria for recurrence and existence of invariant measures for multidimensional diffusions. Annals of Probability, 6(4): 541–553, August 1978. CODEN APBYAE. ISSN 0091-1798 (print),

2168-894X (electronic). URL http://projecteuclid.org/euclid.

aop/1176995476. See correction [Bha80].

Bhattacharya:1980:CNC

[Bha80] R. N. Bhattacharya. Correction note: Correction to “Criteria for Recurrence and Existence of Invariant Measures for Multidimensional Diffusions”. Annals of Probability, 8(6):1194–1195, December 1980. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.aop/

1176994584. See [Bha78].

Billingsley:1974:CDT

[Bil74a] Patrick Billingsley. Conditional distributions and tightness. Annals of Probability, 2(3):480–485, June 1974. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176996663.

Billingsley:1974:NSS

[Bil74b] Patrick Billingsley. A note on separable stochastic processes. Annals of Probability, 2(3):476–479, June 1974. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176996662.

Billingsley:1974:PTA

[Bil74c] Patrick Billingsley. The probability theory of additive arithmetic functions. Annals of Probability, 2(5):749–791, October 1974. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176996547.

Bismut:1979:TAO

[Bis79] Jean-Michel Bismut. Temps d’arrêt optimal, quasi-temps d’arrêt et retournement du temps. (French) [Optimal stopping time, quasistopping time, and time reversal]. Annals of Probability, 7(6):933– 964, December 1979. CODEN APBYAE. ISSN 0091-1798 (print),

2168-894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176994889.

Boldrighini:1978:BP

[BKM78] Carlo Boldrighini, Michael Keane, and Federico Marchetti. Billiards in polygons. Annals of Probability, 6(4):532–540, August 1978. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176995475.

Blasi:1976:RWB

[Bla76] Alessandro Blasi. On a random walk between a reflecting and an absorbing barrier. Annals of Probability, 4(4):695–696, August 1976. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176996041.

Blackwell:1978:BPF

[Bla78] D. Blackwell. Borel-programmable functions. Annals of Probability, 6 (2):321–324, April 1978. CODEN APBYAE. ISSN 0091-1798 (print),

2168-894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176995576.

Bolthausen:1976:FCL

[Bol76] Erwin Bolthausen. On a functional Central Limit Theorem for random walks conditioned to stay positive. Annals of Probability, 4(3):480– 485, June 1976. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.aop/
1176996098.

Bolthausen:1978:SCS

[Bol78] E. Bolthausen. On the speed of convergence in Strassen’s law of the iterated logarithm. Annals of Probability, 6(4):668–672, August 1978. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176995487.

Bondesson:1979:GRI

[Bon79] Lennart Bondesson. A general result on infinite divisibility. Annals of Probability, 7(6):965–979, December 1979. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http:

//projecteuclid.org/euclid.aop/1176994890.

Book:1975:CRC

[Boo75] Stephen A. Book. Convergence rates for a class of large deviation probabilities. Annals of Probability, 3(3):516–525, June 1975. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176996357.

Blum:1973:MIS

[BP73] Julius R. Blum and Pramod K. Pathak. Measure-invariant sets. Annals of Probability, 1(4):590–602, August 1973. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http:

//projecteuclid.org/euclid.aop/1176996888.

Berkes:1979:ATI

[BP79] István Berkes and Walter Philipp. Approximation thorems for independent and weakly dependent random vectors. Annals of Probability, 7(1):29–54, February 1979. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.org/

euclid.aop/1176995146.

Brunel:1974:CPC

[BR74] A. Brunel and D. Revuz. Un critère probabiliste de compacitè des groupes. (French) [A probabilistic criterion for the compactness of groups]. Annals of Probability, 2(4):745–746, August 1974. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176996618.

Blum:1977:IET

[BR77] J. R. Blum and J. I. Reich. On the individual ergodic theorem for K-automorphisms. Annals of Probability, 5(2):309–314, April 1977. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176995857.

Bozorgnia:1979:SLL

[BR79] A. Bozorgnia and M. Bhaskara Rao. A Strong Law of Large Numbers for subsequences of random elements in separable Banach spaces. Annals of Probability, 7(1):156–158, February 1979. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http:

//projecteuclid.org/euclid.aop/1176995157.

Breiman:1973:NMF

[Bre73] Leo Breiman. A note on minimax filtering. Annals of Probability, 1(1):175–179, February 1973. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.org/

euclid.aop/1176997033.

Brillinger:1974:CSA

[Bri74] David R. Brillinger. Cross-spectral analysis of processes with stationary increments including the stationary G/G/∞ queue. Annals of Probability, 2(5):815–827, October 1974. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176996550.

Brillinger:1975:IPP

[Bri75] David R. Brillinger. The identification of point process systems. Annals of Probability, 3(6):909–924, December 1975. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http:

//projecteuclid.org/euclid.aop/1176996218.

Brockett:1977:SID

[Bro77] Patrick L. Brockett. Supports of infinitely divisible measures on Hilbert space. Annals of Probability, 5(6):1012–1017, December 1977.

CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176995668.

Brown:1978:MAP

[Bro78] Tim Brown. A martingale approach to the Poisson convergence of simple point processes. Annals of Probability, 6(4):615–628, August 1978. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176995481.

Brunk:1975:UIC

[Bru75] H. D. Brunk. Uniform inequalities for conditional p-means given σlattices. Annals of Probability, 3(6):1025–1030, December 1975. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176996229.

Block:1976:ICP

[BS76] Henry W. Block and T. H. Savits. The IFRA closure problem. Annals of Probability, 4(6):1030–1032, December 1976. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http:

//projecteuclid.org/euclid.aop/1176995949.

Bhattacharyya:1977:WCR

[BS77a] B. B. Bhattacharyya and P. K. Sen. Weak convergence of the Rao– Blackwell estimator of a distribution function. Annals of Probability, 5(3):500–510, June 1977. CODEN APBYAE. ISSN 0091-1798 (print),

2168-894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176995813.

Blum:1977:EMD

[BS77b] J. R. Blum and V. Susarla. Estimation of a mixing distribution function. Annals of Probability, 5(2):200–209, April 1977. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http:

//projecteuclid.org/euclid.aop/1176995845.

Burkholder:1973:DFI

[Bur73] D. L. Burkholder. Distribution function inequalities for martingales. Annals of Probability, 1(1):19–42, February 1973. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176997023.

Burkholder:1979:SIM

[Bur79] D. L. Burkholder. A sharp inequality for martingale transforms. Annals of Probability, 7(5):858–863, October 1979. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http:

//projecteuclid.org/euclid.aop/1176994944.

Beck:1974:SLL

[BW74] Anatole Beck and Peter Warren. Strong Laws of Large Numbers for weakly orthogonal sequences of Banach space-valued random variables. Annals of Probability, 2(5):918–925, October 1974. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176996557.

Bahadur:1979:LDS

[BZ79] R. R. Bahadur and S. L. Zabell. Large deviations of the sample mean in general vector spaces. Annals of Probability, 7(4):587–621, August 1979. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.aop/
1176994985.

Cambanis:1975:PAC

[Cam75] Stamatis Cambanis. On the path absolute continuity of second order processes. Annals of Probability, 3(6):1050–1054, December 1975. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176996233.

Chung:1977:CP

[CG77] K. L. Chung and R. K. Getoor. The condenser problem. Annals of Probability, 5(1):82–86, February 1977. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176995892.

Chao:1973:NMS

[Cha73] Jia-Arng Chao. A note on martingale square functions. Annals of Probability, 1(6):1059–1060, December 1973. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http:

//projecteuclid.org/euclid.aop/1176996812.

Chamberlain:1974:EFC

[Cha74a] Michael W. Chamberlain. The excessive functions of a continuous time Markov chain. Annals of Probability, 2(6):1075–1089, December 1974. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176996499.

Chan:1974:NCP

[Cha74b] Y. K. Chan. Notes on constructive probability theory. Annals of Probability, 2(1):51–75, February 1974. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176996751.

Chan:1974:CCM

[Cha74c] Y. K. Chan. On constructive convergence of measures on the real line. Annals of Probability, 2(1):131–135, February 1974. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176996757.

Chamberlain:1976:RMC

[Cha76a] Michael W. Chamberlain. The reconstructability of Markov chains. Annals of Probability, 4(1):127–132, February 1976. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http:

//projecteuclid.org/euclid.aop/1176996191.

Chan:1976:CRT

[Cha76b] Y. K. Chan. A constructive renewal theorem. Annals of Probability, 4(4):644–655, August 1976. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.org/

euclid.aop/1176996033.

Chamberlain:1977:LET

[Cha77] Michael W. Chamberlain. Last exit times from the boundary of a continuous time Markov chain. Annals of Probability, 5(1):152–156, February 1977. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.aop/
1176995901.

Chen:1974:CPB

[Che74] Louis H. Y. Chen. On the convergence of Poisson binomial to Poisson distributions. Annals of Probability, 2(1):178–180, February 1974. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176996766.

Chen:1975:ATC

[Che75a] Louis H. Y. Chen. An approximation theorem for convolutions of probability measures. Annals of Probability, 3(6):992–999, December 1975. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.aop/
1176996224.

Chen:1975:PAD

[Che75b] Louis H. Y. Chen. Poisson approximation for dependent trials. Annals of Probability, 3(3):534–545, June 1975. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176996359.

Chen:1978:SNC

[Che78] Louis H. Y. Chen. A short note on the conditional Borel–Cantelli lemma. Annals of Probability, 6(4):699–700, August 1978. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176995492.

Chen:1979:MIS

[Che79] Louis H. Y. Chen. A martingale inequality for the square and maximal functions. Annals of Probability, 7(6):1051–1055, December 1979. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176994898.

Chow:1979:ERT

[CHL79] Y. S. Chow, Chao A. Hsiung, and T. L. Lai. Extended renewal theory and moment convergence in Anscombe’s theorem. Annals of Probability, 7(2):304–318, April 1979. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.org/

euclid.aop/1176995090.

Chung:1973:RWF

[Chu73] Kai Lai Chung. Review: William Feller, An Introduction to Probability Theory and its Applications 2. Annals of Probability, 1(1):193–196, February 1973. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.aop/
1176997039.

Chosid:1978:RRM

[CI78] Leo Chosid and Richard Isaac. On the range of recurrent Markov chains. Annals of Probability, 6(4):680–687, August 1978. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176995489. See correc-

tion [CI80].

Chosid:1980:CNC

[CI80] Leo Chosid and Richard Isaac. Correction notes: Correction to “On the Range of Recurrent Markov Chains”. Annals of Probability, 8(5):

1000, October 1980. CODEN APBYAE. ISSN 0091-1798 (print),

2168-894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176994629. See [CI78].

Cinlar:1975:LTS

[Cin75] Erhan Cinlar. A local time for a storage process. Annals of Probability, 3(6):930–950, December 1975. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.org/

euclid.aop/1176996220.

Cinlar:1976:SSP

[Cin76] Erhan Cinlar. On the structure of semimarkov processes compared with Chung processes. Annals of Probability, 4(3):402–417, June 1976. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176996089.

Cinlar:1977:CSP

[Cin77] Erhan Cinlar. Conversion of semimarkov processes to Chung processes. Annals of Probability, 5(2):180–199, April 1977. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http:

//projecteuclid.org/euclid.aop/1176995844.

Chow:1973:LBW

[CL73] Y. S. Chow and T. L. Lai. Limiting behavior of weighted sums of independent random variables. Annals of Probability, 1(5):810–824, October 1973. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.aop/
1176996847.

Cox:1975:IPP

[CMSB75] D. R. Cox, P. Z. Marmarelis, J. P. Segundo, and David R. Brillinger. [the identification of point process systems]: Discussion of Professor Brillinger’s paper. Annals of Probability, 3(6):924–929, December 1975. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.aop/
1176996219.

Chover:1973:DPS

[CNW73] J. Chover, P. Ney, and S. Wainger. Degeneracy properties of subcritical branching processes. Annals of Probability, 1(4):663–673, August 1973. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.aop/
1176996893.

Cogburn:1975:UTS

[Cog75] Robert Cogburn. A uniform theory for sums of Markov chain transition probabilities. Annals of Probability, 3(2):191–214, April 1975. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176996393.

Cox:1977:ELM

[Cox77a] J. Theodore Cox. Entrance laws for Markov chains. Annals of Probability, 5(4):533–549, August 1977. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.org/

euclid.aop/1176995759.

Cox:1977:ODD

[Cox77b] J. Theodore Cox. On one-dimensional diffusions with time parameter set (−∞, ∞). Annals of Probability, 5(5):807–813, October 1977. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176995724.

Cox:1979:APT

[Cox79a] J. Theodore Cox. An alternate proof of a theorem of Kesten concerning Markov random fields. Annals of Probability, 7(2):377–378, April 1979. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.aop/
1176995095.

Cox:1979:FRO

[Cox79b] J. Theodore Cox. Further results on one-dimensional diffusions with time parameter set S(−∞, ∞). Annals of Probability, 7(3):537– 542, June 1979. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.aop/
1176995054.

Chernoff:1976:OSP

[CP76] H. Chernoff and A. J. Petkau. An optimal stopping problem for sums of dichotomous random variables. Annals of Probability, 4(6):875– 889, December 1976. CODEN APBYAE. ISSN 0091-1798 (print),

2168-894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176995933.

Cambanis:1973:SZO

[CR73] Stamatis Cambanis and Balram S. Rajput. Some zero-one laws for Gaussian processes. Annals of Probability, 1(2):304–312, April 1973.

CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176996982.

Csorgo:1979:HBI

[CR79] M. Csörgő and P. Revesz. How big are the increments of a Wiener process? Annals of Probability, 7(4):731–737, August 1979. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176994994.

Cramer:1976:HCP

[Cra76] Harald Cramér. Half a century with probability theory: Some personal recollections. Annals of Probability, 4(4):509–546, August 1976. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176996025.

Csiszar:1975:DGP

[Csi75] I. Csiszar. I-divergence geometry of probability distributions and minimization problems. Annals of Probability, 3(1):146–158, February 1975. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.aop/
1176996454.

Cuzick:1975:CFM

[Cuz75] Jack Cuzick. Conditions for finite moments of the number of zero crossings for Gaussian processes. Annals of Probability, 3(5):849–858, October 1975. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.aop/
1176996271.

Cuzick:1976:CLT

[Cuz76] Jack Cuzick. A Central Limit Theorem for the number of zeros of a stationary Gaussian process. Annals of Probability, 4(4):547–556, August 1976. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.aop/
1176996026.

Cuzick:1978:LNZ

[Cuz78a] Jack Cuzick. Local nondeterminism and the zeros of Gaussian processes. Annals of Probability, 6(1):72–84, February 1978. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176995611. See correc-

tion [Cuz87].

Cuzick:1978:SLP

[Cuz78b] Jack Cuzick. Some local properties of Gaussian vector fields. Annals of Probability, 6(6):984–994, December 1978. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http:

//projecteuclid.org/euclid.aop/1176995388.

Cuzick:1987:CLN

[Cuz87] Jack Cuzick. Correction: Local nondeterminism and the zeros of Gaussian processes. Annals of Probability, 15(3):1229, July 1987. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176992096. See

[Cuz78a].

Cairoli:1977:MRH

[CW77a] R. Cairoli and J. B. Walsh. Martingale representations and holomorphic processes. Annals of Probability, 5(4):511–521, August 1977. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176995757.

Chan:1977:BET

[CW77b] Y.-K. Chan and John Wierman. On the Berry–Esseen theorem for U-statistics. Annals of Probability, 5(1):136–139, February 1977. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176995897.

deAcosta:1975:SMS

[dA75] Alejandro de Acosta. Stable measures and seminorms. Annals of Probability, 3(5):865–875, October 1975. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176996273.

Daley:1976:AUB

[Dal76] D. J. Daley. Another upper bound for the renewal function. Annals of Probability, 4(1):109–114, February 1976. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http:

//projecteuclid.org/euclid.aop/1176996188.

Daley:1978:UBR

[Dal78] D. J. Daley. Upper bounds for the renewal function via Fourier methods. Annals of Probability, 6(5):876–884, October 1978. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http:

//projecteuclid.org/euclid.aop/1176995434.

Darling:1975:NLT

[Dar75] D. A. Darling. Note on a limit theorem. Annals of Probability, 3(5): 876–878, October 1975. CODEN APBYAE. ISSN 0091-1798 (print),

2168-894X (electronic). URL http://projecteuclid.org/euclid.

aop/1176996274. See correction [Dar76].

Darling:1976:NCB

[Dar76] D. A. Darling. Note: Correction to “Note on a Limit Theorem”. Annals of Probability, 4(4):697, August 1976. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http:

//projecteuclid.org/euclid.aop/1176996042. See [Dar75].

Davies:1976:LHC

[Dav76] Laurie Davies. Local holder conditions for the local times of certain stationary Gaussian processes. Annals of Probability, 4(2):277– 298, April 1976. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.aop/
1176996134.

Davies:1977:EHM

[Dav77] P. Laurie Davies. The exact Hausdorff measure of the zero set of certain stationary Gaussian processes. Annals of Probability, 5(5): 740–755, October 1977. CODEN APBYAE. ISSN 0091-1798 (print),

2168-894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176995716.

Davis:1978:STD

[Dav78] Burgess Davis. On stopping times for n dimensional Brownian motion. Annals of Probability, 6(4):651–659, August 1978. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http:

//projecteuclid.org/euclid.aop/1176995485.

Davis:1979:BMA

[Dav79a] Burgess Davis. Brownian motion and analytic functions. Annals of Probability, 7(6):913–932, December 1979. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http:

//projecteuclid.org/euclid.aop/1176994888.

Davis:1979:MMS

[Dav79b] Richard A. Davis. Maxima and minima of stationary sequences. Annals of Probability, 7(3):453–460, June 1979. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176995046.

DeSmit:1973:SAP

[De 73] Jos H. A. De Smit. A simple analytic proof of the Pollaczek–Wendel identity for ordered partial sums. Annals of Probability, 1(2):348– 351, April 1973. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.aop/
1176996991.

DeHaan:1974:RIL

[De 74a] Laurens De Haan. On random indices and limit distributions. Annals of Probability, 2(1):181, February 1974. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176996767.

deLaVega:1974:ASC

[de 74b] W. Fernandez de La Vega. On almost sure convergence of quadratic Brownian variation. Annals of Probability, 2(3):551–552, June 1974. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176996675.

DeAcosta:1977:ABS

[De 77] Alejandro De Acosta. Asymptotic behavior of stable measures. Annals of Probability, 5(3):494–499, June 1977. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176995812.

DelJunco:1977:DSS

[Del77] Andres Del Junco. On the decomposition of a subadditive stochastic process. Annals of Probability, 5(2):298–302, April 1977. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http:

//projecteuclid.org/euclid.aop/1176995855.

Deo:1973:NEP

[Deo73a] Chandrakant M. Deo. A note on empirical processes of strong-mixing sequences. Annals of Probability, 1(5):870–875, October 1973. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176996855.

Deo:1973:NSM

[Deo73b] Chandrakant M. Deo. A note on strong-mixing Gaussian sequences. Annals of Probability, 1(1):186–187, February 1973. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http:

//projecteuclid.org/euclid.aop/1176997036.

Deo:1973:WCT

[Deo73c] Chandrakant M. Deo. A weak convergence theorem for Gaussian sequences. Annals of Probability, 1(6):1061–1064, December 1973. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176996813.

Deo:1974:NSG

[Deo74] Chandrakant M. Deo. A note on stationary Gaussian sequences. Annals of Probability, 2(5):954–957, October 1974. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http:

//projecteuclid.org/euclid.aop/1176996561.

Deo:1975:FCL

[Deo75] Chandrakant M. Deo. A functional Central Limit Theorem for stationary random fields. Annals of Probability, 3(4):708–715, August 1975. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176996310.

Derriennic:1973:ISE

[Der73] Yves Derriennic. On the integrability of the supremum of ergodic ratios. Annals of Probability, 1(2):338–340, April 1973. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http:

//projecteuclid.org/euclid.aop/1176996988.

Dawson:1979:CDS

[DH79] Donald A. Dawson and Kenneth J. Hochberg. The carrying dimension of a stochastic measure diffusion. Annals of Probability, 7(4):693– 703, August 1979. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.aop/
1176994991.

deHaan:1979:CVP

[dHR79a] L. de Haan and S. I. Resnick. Conjugate Π-variation and process inversion. Annals of Probability, 7(6):1028–1035, December 1979. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176994895.

deHaan:1979:SCS

[dHR79b] Laurens de Haan and Geert Ridder. Stochastic compactness of sample extremes. Annals of Probability, 7(2):290–303, April 1979. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176995089.

Durrett:1977:FBM

[DI77] Richard T. Durrett and Donald L. Iglehart. Functionals of Brownian meander and Brownian excursion. Annals of Probability, 5(1):130– 135, February 1977. CODEN APBYAE. ISSN 0091-1798 (print),

2168-894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176995896.

Diaconis:1977:DLD

[Dia77] Persi Diaconis. The distribution of leading digits and uniform distribution mod1. Annals of Probability, 5(1):72–81, February 1977. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176995891.

Durrett:1977:WCB

[DIM77] Richard T. Durrett, Donald L. Iglehart, and Douglas R. Miller. Weak convergence to Brownian meander and Brownian excursion. Annals of Probability, 5(1):117–129, February 1977. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http:

//projecteuclid.org/euclid.aop/1176995895.

deJonge:1979:CEO

[dJ79] Ep de Jonge. Conditional expectation and ordering. Annals of Probability, 7(1):179–183, February 1979. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176995162.

delJunco:1979:HLV

[dJS79] A. del Junco and J. Michael Steele. Hammersley’s Law for the Van Der Corput sequence: An instance of probability theory for pseudorandom numbers. Annals of Probability, 7(2):267–275, April 1979. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (elec-

tronic). URL http://www.jstor.org/stable/pdfplus/2242879.
pdf;   http://projecteuclid.org/euclid.aop/1176995087.

Dubins:1974:SOP

[DM74] Lester E. Dubins and Isaac Meilijson. On stability for optimization problems. Annals of Probability, 2(2):243–255, April 1974. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176996706.

Daley:1978:ABV

[DM78] D. J. Daley and N. R. Mohan. Asymptotic behaviour of the variance of renewal processes and random walks. Annals of Probability, 6(3):

516–521, June 1978. CODEN APBYAE. ISSN 0091-1798 (print),

2168-894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176995536.

Driscoll:1974:LWS

[DMW74] M. F. Driscoll, J. N. McDonald, and N. A. Weiss. LLN for weakly stationary processes on locally compact Abelian groups. Annals of Probability, 2(6):1168–1171, December 1974. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http:

//projecteuclid.org/euclid.aop/1176996504.

Denzel:1975:LTE

[DO75] G. E. Denzel and G. L. O’Brien. Limit theorems for extreme values of chain-dependent processes. Annals of Probability, 3(5):773–779, October 1975. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.aop/
1176996264.

Dobrushin:1979:GTS

[Dob79] R. L. Dobrushin. Gaussian and their subordinated self-similar random generalized fields. Annals of Probability, 7(1):1–28, February 1979. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176995145.

Doksum:1974:TNR

[Dok74] Kjell Doksum. Tailfree and neutral random probabilities and their posterior distributions. Annals of Probability, 2(2):183–201, April 1974. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.aop/
1176996703.

Durrett:1978:FLT

[DR78] Richard Durrett and Sidney I. Resnick. Functional limit theorems for dependent variables. Annals of Probability, 6(5):829–846, October 1978. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.aop/
1176995431.

Drogin:1973:CPB

[Dro73] Richard Drogin. On convergence in probability to Brownian motion. Annals of Probability, 1(2):254–262, April 1973. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http:

//projecteuclid.org/euclid.aop/1176996978.

Dharmadhikari:1975:CMN

[DS75a] S. W. Dharmadhikari and M. Sreehari. On convergence in r-mean of normalized partial sums. Annals of Probability, 3(6):1023–1024, December 1975. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.aop/
1176996228.

Dubins:1975:EWS

[DS75b] Lester E. Dubins and William D. Sudderth. An example in which stationary strategies are not adequate. Annals of Probability, 3(4): 722–725, August 1975. CODEN APBYAE. ISSN 0091-1798 (print),

2168-894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176996312.

Dubuc:1976:LLT

[DS76] S. Dubuc and E. Seneta. The local limit theorem for the Galton– Watson process. Annals of Probability, 4(3):490–496, June 1976. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176996100.

Diaconis:1978:STT

[DS78] Persi Diaconis and Charles Stein. Some Tauberian theorems related to coin tossing. Annals of Probability, 6(3):483–490, June 1978. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176995532.

Dubins:1979:SSA

[DS79] Lester E. Dubins and William D. Sudderth. On stationary strategies for absolutely continuous houses. Annals of Probability, 7(3):461– 476, June 1979. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.aop/
1176995047.

Daffer:1979:LLN

[DT79] Peter Z. Daffer and Robert L. Taylor. Laws of large numbers for D[0, 1]. Annals of Probability, 7(1):85–95, February 1979. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176995150.

Dubins:1973:WFS

[Dub73] Lester E. Dubins. Which functions of stopping times are stopping times? Annals of Probability, 1(2):313–316, April 1973. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http:

//projecteuclid.org/euclid.aop/1176996983.

Dubins:1974:LLE

[Dub74] Lester E. Dubins. On Lebesgue-like extensions of finitely additive measures. Annals of Probability, 2(3):456–463, June 1974. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http:

//projecteuclid.org/euclid.aop/1176996660.

Dubins:1975:FAC

[Dub75] Lester E. Dubins. Finitely additive conditional probabilities, conglomerability and disintegrations. Annals of Probability, 3(1):89–99, February 1975. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.aop/
1176996451.

Dudley:1973:CNC

[Dud73a] R. M. Dudley. Correction note: Correction to “A Counterexample on Measurable Processes”. Annals of Probability, 1(1):191–192, February 1973. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.aop/
1176997038.

Dudley:1973:SFG

[Dud73b] R. M. Dudley. Sample functions of the Gaussian process. Annals of Probability, 1(1):66–103, February 1973. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176997026.

Dudley:1977:WFI

[Dud77] R. M. Dudley. Wiener functionals as Itô integrals. Annals of Probability, 5(1):140–141, February 1977. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.org/

euclid.aop/1176995898.

Dudley:1978:CLT

[Dud78] R. M. Dudley. Central Limit Theorems for empirical measures. Annals of Probability, 6(6):899–929, December 1978. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176995384. See correc-

tions [Dud79].

Dudley:1979:CBL

[Dud79] R. M. Dudley. Corrections to “Central Limit Theorems for Empirical Measures”. Annals of Probability, 7(5):909–911, October 1979. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176994953. See [Dud78].

Duncan:1977:PCT

[Dun77] Richard Duncan. Pointwise convergence theorems for self-adjoint and unitary contractions. Annals of Probability, 5(4):622–626, August 1977. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.aop/
1176995773.

Durrett:1978:CLT

[Dur78] Richard Durrett. Conditioned limit theorems for some null recurrent Markov processes. Annals of Probability, 6(5):798–828, October 1978. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176995430.

Durrett:1979:SRS

[Dur79] Richard Durrett. On the shape of a random string. Annals of Probability, 7(6):1014–1027, December 1979. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176994894.

Davis:1974:MIF

[DV74] M. H. A. Davis and P. Varaiya. The multiplicity of an increasing family of Σ-fields. Annals of Probability, 2(5):958–963, October 1974. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176996562.

Dym:1978:CCS

[Dym78] Harry Dym. Conditions for a class of stationary Gaussian processes to be Kolmogorov mixing. Annals of Probability, 6(1):159–161, February 1978. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.aop/
1176995621.

Dynkin:1977:MST

[Dyn77] E. B. Dynkin. Markov systems and their additive functionals. Annals of Probability, 5(5):653–677, October 1977. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http:

//projecteuclid.org/euclid.aop/1176995711.

Dynkin:1978:SSE

[Dyn78] E. B. Dynkin. Sufficient statistics and extreme points. Annals of Probability, 6(5):705–730, October 1978. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176995424.

Eagleson:1975:MCM

[Eag75] G. K. Eagleson. Martingale convergence to mixtures of infinitely divisible laws. Annals of Probability, 3(3):557–562, June 1975. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176996362.

Erickson:1975:VCS

[EF75] R. V. Erickson and Vaclav Fabian. On vague convergence of stochastic processes. Annals of Probability, 3(6):1014–1022, December 1975. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176996227.

Erickson:1973:CEL

[EG73] K. Bruce Erickson and Harry Guess. A characterization of the exponential law. Annals of Probability, 1(1):183–185, February 1973. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176997035.

Ein-Gal:1975:PMP

[EGBD75] Moshe Ein-Gal and Israel Bar-David. Passages and maxima for a particular Gaussian process. Annals of Probability, 3(3):549–556, June 1975. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.aop/
1176996361.

Eisenberg:1976:PTE

[Eis76] Bennett Eisenberg. Prediction theory and ergodic spectral decompositions. Annals of Probability, 4(1):98–101, February 1976. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176996185.

Erickson:1974:SWL

[EK74] K. Bruce Erickson and Harry Kesten. Strong and weak limit points of a normalized random walk. Annals of Probability, 2(4):553–579, August 1974. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.aop/
1176996604.

Esary:1973:SMW

[EM73] J. D. Esary and A. W. Marshall. Shock models and wear processes. Annals of Probability, 1(4):627–649, August 1973. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176996891.

Esary:1979:MDI

[EM79] J. D. Esary and A. W. Marshall. Multivariate distributions with increasing hazard rate average. Annals of Probability, 7(2):359–370, April 1979. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.aop/
1176995093.

Erdos:1974:SPR

[EN74] P. Erdős and P. Ney. Some problems on random intervals and annihilating particles. Annals of Probability, 2(5):828–839, October 1974. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176996551.

Eaton:1977:RGG

[EP77] Morris L. Eaton and Michael D. Perlman. Reflection groups, generalized Schur functions, and the geometry of majorization. Annals of Probability, 5(6):829–860, December 1977. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http:

//projecteuclid.org/euclid.aop/1176995655.

Erickson:1973:SAB

[Eri73a] K. Bruce Erickson. Self annihilating branching processes. Annals of Probability, 1(6):926–946, December 1973. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http:

//projecteuclid.org/euclid.aop/1176996801.

Erickson:1973:CCA

[Eri73b] R. V. Erickson. On L_{p} Chebyshev–Cramér asymptotic expansions. Annals of Probability, 1(2):355–361, April 1973. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176996993.

Erickson:1973:VBE

[Eri73c] R. V. Erickson. On an L_{p} version of the Berry–Esseen theorem for independent and m-dependent variables. Annals of Probability, 1(3): 497–503, June 1973. CODEN APBYAE. ISSN 0091-1798 (print),

2168-894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176996944.

Erickson:1974:BAN

[Eri74] R. V. Erickson. L_{1} bounds for asymptotic normality of m-dependent sums using Stein’s technique. Annals of Probability, 2(3):522–529, June 1974. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.aop/
1176996670.

Erickson:1976:RSN

[Eri76] K. Bruce Erickson. Recurrence sets of normed random walk in R^{d}. Annals of Probability, 4(5):802–828, October 1976. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http:

//projecteuclid.org/euclid.aop/1176995985.

Erickson:1977:SDR

[Eri77] K. Bruce Erickson. Slowing down d-dimensional random walks. Annals of Probability, 5(4):645–651, August 1977. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http:

//projecteuclid.org/euclid.aop/1176995776.

Esty:1975:CAD

[Est75] Warren W. Esty. Critical age-dependent branching processes. Annals of Probability, 3(1):49–60, February 1975. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176996447.

Etemadi:1976:RRW

[Ete76] Nasrollah Etemadi. The range of a random walk in two-dimensional time. Annals of Probability, 4(5):836–843, October 1976. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http:

//projecteuclid.org/euclid.aop/1176995987.

Ethier:1979:LTA

[Eth79] S. N. Ethier. Limit theorems for absorption times of genetic models. Annals of Probability, 7(4):622–638, August 1979. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http:

//projecteuclid.org/euclid.aop/1176994986.

Fearn:1976:SAD

[Fea76] Dean H. Fearn. Supercritical age dependent branching processes with generation dependence. Annals of Probability, 4(1):27–37, February 1976. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.aop/
1176996178.

Feldman:1977:ORS

[Fel77] Richard M. Feldman. Optimal replacement for systems governed by Markov additive shock processes. Annals of Probability, 5(3):413– 429, June 1977. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.aop/
1176995802.

Fine:1976:ECP

[FG76] Terrence Fine and John Gill. The enumeration of comparative probability relations. Annals of Probability, 4(4):667–673, August 1976. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176996036.

Fishburn:1975:WCP

[Fis75] Peter C. Fishburn. Weak comparative probability on infinite sets. Annals of Probability, 3(5):889–893, October 1975. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http:

//projecteuclid.org/euclid.aop/1176996277.

Flatto:1976:MRT

[Fla76] Leopold Flatto. The multiple range of two-dimensional recurrent walk. Annals of Probability, 4(2):229–248, April 1976. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http:

//projecteuclid.org/euclid.aop/1176996131.

Follmer:1973:RS

[Fol73] Hans Follmer. On the representation of semimartingales. Annals of Probability, 1(4):580–589, August 1973. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176996887.

Flatto:1975:RRW

[FP75] Leopold Flatto and Joel Pitt. Recurrent random walks on countable Abelian groups of rank 2. Annals of Probability, 3(2):380– 386, April 1975. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.aop/
1176996412.

Fraser:1973:RCR

[Fra73] David F. Fraser. The rate of convergence of a random walk to Brownian motion. Annals of Probability, 1(4):699–701, August 1973. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176996896.

Frankel:1976:NDE

[Fra76] Judah M. Frankel. A note on downcrossings for extremal processes. Annals of Probability, 4(1):151–152, February 1976. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http:

//projecteuclid.org/euclid.aop/1176996195.

Freedman:1973:ANB

[Fre73] David Freedman. Another note on the Borel–Cantelli lemma and the Strong Law, with the Poisson approximation as a by-product. Annals of Probability, 1(6):910–925, December 1973. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http:

//projecteuclid.org/euclid.aop/1176996800.

Freedman:1974:PAD

[Fre74a] David Freedman. The Poisson approximation for dependent events. Annals of Probability, 2(2):256–269, April 1974. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176996707.

Freedman:1974:RSL

[Fre74b] David Freedman. A remark on the Strong Law. Annals of Probability, 2(2):324–327, April 1974. CODEN APBYAE. ISSN 0091-1798 (print),

2168-894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176996713.

Freedman:1974:ORO

[Fre74c] David A. Freedman. The optimal reward operator in special classes of dynamic programming problems. Annals of Probability, 2(5):942–949, October 1974. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.aop/
1176996559.

Freedman:1975:TPM

[Fre75] David A. Freedman. On tail probabilities for martingales. Annals of Probability, 3(1):100–118, February 1975. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http:

//projecteuclid.org/euclid.aop/1176996452.

Fristedt:1979:ULB

[Fri79] Bert Fristedt. Uniform local behavior of stable subordinators. Annals of Probability, 7(6):1003–1013, December 1979. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http:

//projecteuclid.org/euclid.aop/1176994893.

Gabriel:1977:MCF

[Gab77a] J.-P. Gabriel. Martingales with a countable filtering index set. Annals of Probability, 5(6):888–898, December 1977. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http:

//projecteuclid.org/euclid.aop/1176995658.

Gabriel:1977:ISI

[Gab77b] Jean-Pierre Gabriel. An inequality for sums of independent random variables indexed by finite dimensional filtering sets and its applications to the convergence of series. Annals of Probability, 5(5):779–786, October 1977. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.aop/
1176995719.

Galambos:1977:BI

[Gal77] Janos Galambos. Bonferroni inequalities. Annals of Probability, 5(4): 577–581, August 1977. CODEN APBYAE. ISSN 0091-1798 (print),

2168-894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176995765.

Garsia:1973:CFI

[Gar73] Adriano M. Garsia. On a convex function inequality for martingales. Annals of Probability, 1(1):171–174, February 1973. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http:

//projecteuclid.org/euclid.aop/1176997032.

Garling:1976:FCL

[Gar76] D. J. H. Garling. Functional Central Limit Theorems in Banach spaces. Annals of Probability, 4(4):600–611, August 1976. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176996030.

Garling:1978:RSI

[Gar78] D. J. H. Garling. The range of stochastic integration. Annals of Probability, 6(2):332–334, April 1978. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176995578.

Ghosh:1977:PMD

[GB77] Malay Ghosh and Gutti Jogesh Babu. Probabilities of moderate deviations for some stationary ϕ-mixing processes. Annals of Probability, 5 (2):222–234, April 1977. CODEN APBYAE. ISSN 0091-1798 (print),

2168-894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176995847.

Geman:1973:NDH

[Gem73] Donald Geman. A note on the distribution of hitting times. Annals of Probability, 1(5):854–856, October 1973. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176996851.

Getoor:1979:BEP

[Get79a] R. K. Getoor. The Brownian escape process. Annals of Probability, 7(5):864–867, October 1979. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.org/

euclid.aop/1176994945.

Getoor:1979:EMP

[Get79b] R. K. Getoor. Excursions of a Markov process. Annals of Probability, 7(2):244–266, April 1979. CODEN APBYAE. ISSN 0091-1798 (print),

2168-894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176995086.

Gray:1977:UNP

[GG77] Lawrence Gray and David Griffeath. On the uniqueness and nonuniqueness of proximity processes. Annals of Probability, 5(5): 678–692, October 1977. CODEN APBYAE. ISSN 0091-1798 (print),

2168-894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176995712.

Geman:1973:OTS

[GH73] D. Geman and J. Horowitz. Occupation times for smooth stationary processes. Annals of Probability, 1(1):131–137, February 1973. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176997029.

Goodsell:1976:ASC

[GH76] C. A. Goodsell and D. L. Hanson. Almost sure convergence for the Robbins–Monro process. Annals of Probability, 4(6):890–901, December 1976. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.aop/
1176995934.

Geman:1976:RSS

[GHZ76] Donald Geman, Joseph Horowitz, and Joel Zinn. Recurrence of stationary sequences. Annals of Probability, 4(3):372–381, June 1976. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176996086.

Gianini:1977:ISP

[Gia77] Jacqueline Gianini. The infinite secretary problem as the limit of the finite problem. Annals of Probability, 5(4):636–644, August 1977. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176995775.

Gilat:1977:ENS

[Gil77] David Gilat. Every nonnegative submartingale is the absolute value of a martingale. Annals of Probability, 5(3):475–481, June 1977. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176995809.

Gisselquist:1973:CCP

[Gis73] Richard Gisselquist. A continuum of collision process limit theorems. Annals of Probability, 1(2):231–239, April 1973. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http:

//projecteuclid.org/euclid.aop/1176996976.

Gleser:1975:DNS

[Gle75] Leon Jay Gleser. On the distribution of the number of successes in independent trials. Annals of Probability, 3(1):182–188, February 1975. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176996461.

Glover:1979:NRK

[Glo79] Joseph Glover. Note on the Ray–Knight compactification. Annals of Probability, 7(3):543–546, June 1979. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176995055.

GarciaAlvarez:1973:TDE

[GM73] M. A. Garcia Alvarez and P. A. Meyer. Une théorie de la dualité à ensemble polaire près I. (French) [A duality theory for the near polar ensemble I]. Annals of Probability, 1(2):207–222, April 1973.

CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176996974.

Greenwood:1977:RSP

[GM77] Priscilla Greenwood and Itrel Monroe. Random stopping preserves regular variation of process distributions. Annals of Probability, 5(1): 42–51, February 1977. CODEN APBYAE. ISSN 0091-1798 (print),

2168-894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176995889.

Glaz:1979:MCL

[GN79] Joseph Glaz and Joseph Naus. Multiple coverage of the line. Annals of Probability, 7(5):900–906, October 1979. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176994951.

Gray:1975:NSC

[GNO75] Robert M. Gray, David L. Neuhoff, and Donald S. Ornstein. Nonblock source coding with a fidelity criterion. Annals of Probability, 3(3): 478–491, June 1975. CODEN APBYAE. ISSN 0091-1798 (print),

2168-894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176996354.

Gray:1975:GOD

[GNS75] Robert M. Gray, David L. Neuhoff, and Paul C. Shields. A generalization of Ornstein’s d̄ distance with applications to information theory. Annals of Probability, 3(2):315–328, April 1975. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http:

//projecteuclid.org/euclid.aop/1176996402.

Goodman:1976:DEF

[Goo76] Victor Goodman. Distribution estimates for functionals of the twoparameter Wiener process. Annals of Probability, 4(6):977–982, December 1976. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.aop/
1176995940.

Gorostiza:1973:CLT

[Gor73] Luis G. Gorostiza. The Central Limit Theorem for random motions of d-dimensional Euclidean space. Annals of Probability, 1(4):603– 612, August 1973. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.aop/
1176996889.

Groeneboom:1979:LDT

[GOR79] P. Groeneboom, J. Oosterhoff, and F. H. Ruymgaart. Large deviation theorems for empirical probability measures. Annals of Probability, 7(4):553–586, August 1979. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.org/

euclid.aop/1176994984.

Griffiths:1973:CZQ

[GPW73] Gary N. Griffiths, Ronald D. Platt, and F. T. Wright. Convergence to zero of quadratic forms in independent random variables. Annals of Probability, 1(5):838–848, October 1973. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http:

//projecteuclid.org/euclid.aop/1176996849.

Gray:1978:CSF

[Gra78] Lawrence Gray. Controlled spin-flip systems. Annals of Probability, 6(6):953–974, December 1978. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.org/

euclid.aop/1176995386.

Greenwood:1973:ARS

[Gre73] Priscilla Greenwood. Asymptotics of randomly stopped sequences with independent increments. Annals of Probability, 1(2):317–321, April 1973. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.aop/
1176996984.

Greenwood:1974:M

[Gre74] Priscilla Greenwood. The martintote. Annals of Probability, 2(1): 84–89, February 1974. CODEN APBYAE. ISSN 0091-1798 (print),

2168-894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176996753.

Greenwood:1975:ETS

[Gre75] Priscilla Greenwood. Extreme time of stochastic processes with stationary independent increments. Annals of Probability, 3(4):664–676, August 1975. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.aop/
1176996307.

Griffeath:1973:MOP

[Gri73] David Griffeath. The maximal oscillation problem for regenerative phenomena. Annals of Probability, 1(3):405–416, June 1973. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176996935.

Grimvall:1974:CSB

[Gri74] Anders Grimvall. On the convergence of sequences of branching processes. Annals of Probability, 2(6):1027–1045, December 1974. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176996496.

Griffeath:1976:PCL

[Gri76] David Griffeath. Partial coupling and loss of memory for Markov chains. Annals of Probability, 4(5):850–858, October 1976. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176995989.

Griffeath:1978:LTN

[Gri78] David Griffeath. Limit theorems for nonergodic set-valued Markov processes. Annals of Probability, 6(3):379–387, June 1978. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176995524.

Griffeath:1979:PEB

[Gri79] David Griffeath. Pointwise ergodicity of the basic contact process. Annals of Probability, 7(1):139–142, February 1979. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http:

//projecteuclid.org/euclid.aop/1176995154.

Grosswald:1976:SDO

[Gro76] Emil Grosswald. The Student t-distribution for odd degrees of freedom is infinitely divisible. Annals of Probability, 4(4):680–683, August 1976. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.aop/
1176996038.

Grunkemeier:1975:DFB

[Gru75] Gary L. Grunkemeier. Decomposition of functions of bounded variation. Annals of Probability, 3(2):329–337, April 1975. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http:

//projecteuclid.org/euclid.aop/1176996403.

Getoor:1973:LET

[GS73] R. K. Getoor and M. J. Sharpe. Last exit times and additive functionals. Annals of Probability, 1(4):550–569, August 1973. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http:

//projecteuclid.org/euclid.aop/1176996885.

Griffeath:1974:OSS

[GS74] David Griffeath and J. Laurie Snell. Optimal stopping in the stock market. Annals of Probability, 2(1):1–13, February 1974. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http:

//projecteuclid.org/euclid.aop/1176996747.

Gacs:1975:PCC

[GS75] P. Gacs and D. Szasz. On a problem of Cox concerning point processes in R^{k} of “controlled variability”. Annals of Probability, 3(4):597– 607, August 1975. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.aop/
1176996303.

Gianini:1976:ISP

[GS76] Jacqueline Gianini and Stephen M. Samuels. The infinite secretary problem. Annals of Probability, 4(3):418–432, June 1976. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176996090.

Gilat:1977:TPW

[GS77] David Gilat and William Sudderth. Timid play when large bets are profitable. Annals of Probability, 5(4):573–576, August 1977. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176995764.

Greenwood:1978:DPS

[GS78] Priscilla Greenwood and Moshe Shaked. Dual pairs of stopping times for random walk. Annals of Probability, 6(4):644–650, August 1978. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176995484.

Gaenssler:1979:EPS

[GS79] Peter Gaenssler and Winfried Stute. Empirical processes: A survey of results for independent and identically distributed random variables. Annals of Probability, 7(2):193–243, April 1979. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http:

//projecteuclid.org/euclid.aop/1176995085.

Guess:1974:LTS

[Gue74] Harry A. Guess. Limit theorems for some stochastic evolution models. Annals of Probability, 2(1):14–31, February 1974. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http:

//projecteuclid.org/euclid.aop/1176996748.

Gut:1974:CMS

[Gut74a] Allan Gut. On convergence in r-mean of some first passage times and randomly indexed partial sums. Annals of Probability, 2(2):321– 323, April 1974. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.aop/
1176996712.

Gut:1974:MLD

[Gut74b] Allan Gut. On the moments and limit distributions of some first passage times. Annals of Probability, 2(2):277–308, April 1974. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176996709.

Gut:1978:MLC

[Gut78] Allan Gut. Marcinkiewicz laws and convergence rates in the Law of Large Numbers for random variables with multidimensional indices. Annals of Probability, 6(3):469–482, June 1978. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http:

//projecteuclid.org/euclid.aop/1176995531.

Gut:1979:IS

[Gut79] Allan Gut. On the integrability of sup |S_{n}/n| for subsequences. Annals of Probability, 7(6):1059–1065, December 1979. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http:

//projecteuclid.org/euclid.aop/1176994900.

Gergley:1975:DMS

[GY75] T. Gergley and I. I. Yezhow. On the distribution of the maximum of the sequence of sums of independent random variables. Annals of Probability, 3(2):289–297, April 1975. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176996399.

Geman:1978:IMR

[GZ78] Donald Geman and Joel Zinn. On the increments of multidimensional random fields. Annals of Probability, 6(1):151–158, February 1978.

CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176995620.

Gzyl:1977:LST

[Gzy77] Henryk Gzyl. Lévy systems for time-changed processes. Annals of Probability, 5(4):565–570, August 1977. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176995762.

Gzyl:1978:GSS

[Gzy78] Henryk Gzyl. On generators of subordinate semigroups. Annals of Probability, 6(6):975–983, December 1978. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176995387.

Hahn:1977:CSC

[Hah77] Marjorie G. Hahn. Conditions for sample-continuity and the Central Limit Theorem. Annals of Probability, 5(3):351–360, June 1977. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176995796.

Halfin:1975:ECI

[Hal75] S. Halfin. Explicit construction of invariant measures for a class of continuous state Markov processes. Annals of Probability, 3(5):859– 864, October 1975. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.aop/
1176996272.

Hall:1977:MIP

[Hal77] Peter Hall. Martingale invariance principles. Annals of Probability, 5(6):875–887, December 1977. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.org/

euclid.aop/1176995657.

Hall:1979:SRA

[Hal79] Peter Hall. On the Skorokhod representation approach to martingale invariance principles. Annals of Probability, 7(2):371–376, April 1979. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176995094.

Hammersley:1974:PSP

[Ham74] J. M. Hammersley. Postulates for subadditive processes. Annals of Probability, 2(4):652–680, August 1974. CODEN APBYAE.

ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176996611.

Harris:1974:CIL

[Har74] T. E. Harris. Contact interactions on a lattice. Annals of Probability, 2(6):969–988, December 1974. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.org/

euclid.aop/1176996493.

Harris:1976:CSV

[Har76] T. E. Harris. On a class of set-valued Markov processes. Annals of Probability, 4(2):175–194, April 1976. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176996129.

Harris:1977:CIM

[Har77] T. E. Harris. A correlation inequality for Markov processes in partially ordered state spaces. Annals of Probability, 5(3):451–454, June 1977. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176995804.

Harris:1978:ASV

[Har78] T. E. Harris. Additive set-valued Markov processes and graphical methods. Annals of Probability, 6(3):355–378, June 1978. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176995523.

Ho:1978:BRC

[HC78a] Soo-Thong Ho and Louis H. Y. Chen. An L_{p} bound for the remainder in a combinatorial Central Limit Theorem. Annals of Probability, 6 (2):231–249, April 1978. CODEN APBYAE. ISSN 0091-1798 (print),

2168-894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176995570.

Huang:1978:SMW

[HC78b] Steel T. Huang and Stamatis Cambanis. Stochastic and multiple Wiener integrals for Gaussian processes. Annals of Probability, 6(4): 585–614, August 1978. CODEN APBYAE. ISSN 0091-1798 (print),

2168-894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176995480.

Heath:1977:IM

[Hea77] David Heath. Interpolation of martingales. Annals of Probability, 5(5): 804–806, October 1977. CODEN APBYAE. ISSN 0091-1798 (print),

2168-894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176995723.

Helmers:1977:ONA

[Hel77] R. Helmers. The order of the normal approximation for linear combinations of order statistics with smooth weight functions. Annals of Probability, 5(6):940–953, December 1977. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http:

//projecteuclid.org/euclid.aop/1176995662.

Hendricks:1973:DTS

[Hen73] W. J. Hendricks. A dimension theorem for sample functions of processes with stable components. Annals of Probability, 1(5):849–853, October 1973. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.aop/
1176996850.

Heyde:1974:ILR

[Hey74] C. C. Heyde. An iterated logarithm result for autocorrelations of a stationary linear process. Annals of Probability, 2(2):328–332, April 1974. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176996714.

Heyde:1975:NBC

[Hey75] C. C. Heyde. A nonuniform bound on convergence to normality. Annals of Probability, 3(5):903–907, October 1975. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http:

//projecteuclid.org/euclid.aop/1176996280.

Heyde:1978:LLI

[Hey78] C. C. Heyde. A log log improvement to the Riemann hypothesis for the Hawkins random sieve. Annals of Probability, 6(5):870–875, October 1978. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.aop/
1176995433.

Hebda-Grabowska:1979:ASI

[HGS79] H. Hebda-Grabowska and D. Szynal. An almost sure invariance principle for the partial sums of infima of independent random variables.

Annals of Probability, 7(6):1036–1045, December 1979. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http:

//projecteuclid.org/euclid.aop/1176994896.

Hipp:1977:EEI

[Hip77] C. Hipp. Edgeworth expansions for integrals of smooth functions. Annals of Probability, 5(6):1004–1011, December 1977. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http:

//projecteuclid.org/euclid.aop/1176995667.

Hoffmann-Jorgensen:1976:LLN

[HJP76] J. Hoffmann-Jorgensen and G. Pisier. The Law of Large Numbers and the Central Limit Theorem in Banach spaces. Annals of Probability, 4(4):587–599, August 1976. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.org/

euclid.aop/1176996029.

Hoffmann-Jorgensen:1979:LTG

[HJSD79] J. Hoffmann-Jorgensen, L. A. Shepp, and R. M. Dudley. On the lower tail of Gaussian seminorms. Annals of Probability, 7(2):319– 342, April 1979. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.aop/
1176995091.

Hahn:1977:SCS

[HK77] Marjorie G. Hahn and Michael J. Klass. Sample-continuity of squareintegrable processes. Annals of Probability, 5(3):361–370, June 1977. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176995797.

Holley:1975:ETW

[HL75] Richard A. Holley and Thomas M. Liggett. Ergodic theorems for weakly interacting infinite systems and the voter model. Annals of Probability, 3(4):643–663, August 1975. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176996306.

Hall:1977:CC

[HL77] W. J. Hall and R. M. Loynes. On the concept of contiguity. Annals of Probability, 5(2):278–282, April 1977. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176995851.

Holley:1978:SCP

[HL78] R. Holley and T. M. Liggett. The survival of contact processes. Annals of Probability, 6(2):198–206, April 1978. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176995567.

Hudson:1975:MEI

[HM75] W. N. Hudson and J. D. Mason. More on equivalence of infinitely divisible distributions. Annals of Probability, 3(3):563–568, June 1975. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176996363.

Halasz:1977:RDP

[HM77] G. Halasz and P. Major. Reconstructing the distribution from partial sums of samples. Annals of Probability, 5(6):987–998, December 1977. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176995665.

Huntington:1975:SET

[HN75] Raymond J. Huntington and Joseph I. Naus. A simpler expression for K th nearest neighbor coincidence probabilities. Annals of Probability, 3(5):894–896, October 1975. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.org/

euclid.aop/1176996278.

Hochberg:1978:SMP

[Hoc78] Kenneth J. Hochberg. A signed measure on path space related to Wiener measure. Annals of Probability, 6(3):433–458, June 1978. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176995529.

Holst:1977:SAR

[Hol77] Lars Holst. Some asymptotic results for occupancy problems. Annals of Probability, 5(6):1028–1035, December 1977. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http:

//projecteuclid.org/euclid.aop/1176995671.

Holst:1979:ANS

[Hol79] Lars Holst. Asymptotic normality of sum-functions of spacings. Annals of Probability, 7(6):1066–1072, December 1979. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http:

//projecteuclid.org/euclid.aop/1176994901.

Hoppe:1976:SMB

[Hop76] Fred M. Hoppe. Supercritical multitype branching processes. Annals of Probability, 4(3):393–401, June 1976. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176996088.

Hoppe:1977:RIM

[Hop77] Fred M. Hoppe. Representations of invariant measures on multitype Galton–Watson processes. Annals of Probability, 5(2):291–297, April 1977. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.aop/
1176995854.

Hosoya:1978:DTS

[Hos78a] Yuzo Hosoya. Discrete-time stable processes and their certain properties. Annals of Probability, 6(1):94–105, February 1978. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176995613.

Hosoya:1978:RLE

[Hos78b] Yuzo Hosoya. Robust linear extrapolations of second-order stationary processes. Annals of Probability, 6(4):574–584, August 1978. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176995479.

Halevy:1979:AKT

[HR79] Avner Halevy and M. Bhaskara Rao. On an analogue of Komlos’ theorem for strategies. Annals of Probability, 7(6):1073–1077, December 1979. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.aop/
1176994902.

Heyde:1973:IPL

[HS73] C. C. Heyde and D. J. Scott. Invariance principles for the law of the iterated logarithm for martingales and processes with stationary increments. Annals of Probability, 1(3):428–436, June 1973. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176996937.

Holley:1976:MAI

[HS76] R. A. Holley and D. W. Stroock. A martingale approach to infinite systems of interacting processes. Annals of Probability, 4(2):

195–228, April 1976. CODEN APBYAE. ISSN 0091-1798 (print),

2168-894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176996130.

Hudson:1975:EID

[HT75] William N. Hudson and Howard G. Tucker. Equivalence of infinitely divisible distributions. Annals of Probability, 3(1):70–79, February 1975. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.aop/
1176996449.

Hudson:1979:AIM

[HT79] William N. Hudson and Howard G. Tucker. Asymptotic independence in the multivariate Central Limit Theorem. Annals of Probability, 7(4):662–671, August 1979. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.org/

euclid.aop/1176994989.

Hudson:1976:ATP

[Hud76] William N. Hudson. Admissible translates for probability distributions. Annals of Probability, 4(3):505–508, June 1976. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http:

//projecteuclid.org/euclid.aop/1176996103.

Hartman:1974:NDF

[HW74] Philip Hartman and Geoffrey S. Watson. “Normal” distribution functions on spheres and the modified Bessel functions. Annals of Probability, 2(4):593–607, August 1974. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.org/

euclid.aop/1176996606.

Hwang:1976:RPG

[Hwa76] J. S. Hwang. On the recurrence property of Gaussian Taylor series. Annals of Probability, 4(3):453–455, June 1976. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http:

//projecteuclid.org/euclid.aop/1176996093.

Hwang:1977:GKM

[Hwa77] F. K. Hwang. A generalization of the Karlin–McGregor theorem on coincidence probabilities and an application to clustering. Annals of Probability, 5(5):814–817, October 1977. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http:

//projecteuclid.org/euclid.aop/1176995725.

Iglehart:1974:FCL

[Igl74] Donald L. Iglehart. Functional Central Limit Theorems for random walks conditioned to stay positive. Annals of Probability, 2(4):608– 619, August 1974. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.aop/
1176996607.

Imhof:1976:OCS

[Imh76] J. P. Imhof. Oscillations of continuous symmetric random walk. Annals of Probability, 4(4):662–666, August 1976. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http:

//projecteuclid.org/euclid.aop/1176996035.

Isaac:1979:MDF

[Isa79] Richard Isaac. Markov-dependent σ-fields and conditional expectations. Annals of Probability, 7(6):1088–1091, December 1979. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176994905.

Ismail:1977:BFI

[Ism77] Mourad E. H. Ismail. Bessel functions and the infinite divisibility of the Student t-distribution. Annals of Probability, 5(4):582–585, August 1977. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.aop/
1176995766.

Ivkovic:1975:DWR

[Ivk75] Z. Ivkovic. Deduction of Wold representation of stationary processes from Cramér representation of second-order processes. Annals of Probability, 3(3):546–548, June 1975. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176996360.

Jain:1976:ECC

[Jai76] Naresh C. Jain. An example concerning CLT and LIL in Banach space. Annals of Probability, 4(4):690–694, August 1976. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176996040.

James:1975:FLI

[Jam75] Barry R. James. A functional law of the iterated logarithm for weighted empirical distributions. Annals of Probability, 3(5):762–772, October 1975. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.aop/
1176996263.

James:1976:IDI

[Jam76] Ian R. James. Infinite F -divisibility of integer-valued random variables. Annals of Probability, 4(2):326–334, April 1976. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http:

//projecteuclid.org/euclid.aop/1176996138.

Jain:1975:ULF

[JJS75] Naresh C. Jain, Kumar Jogdeo, and William F. Stout. Upper and lower functions for martingales and mixing processes. Annals of Probability, 3(1):119–145, February 1975. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176996453.

Jupp:1978:DPS

[JM78] P. E. Jupp and K. V. Mardia. Density-preserving statistics and densities for sample means. Annals of Probability, 6(4):688–694, August 1978. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.aop/
1176995490.

Joffe:1974:SAD

[Jof74] A. Joffe. On a set of almost deterministic k-independent random variables. Annals of Probability, 2(1):161–162, February 1974. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176996762.

Jones:1976:EMF

[Jon76] Roger L. Jones. The ergodic maximal function with cancellation. Annals of Probability, 4(1):91–97, February 1976. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http:

//projecteuclid.org/euclid.aop/1176996184.

Joshi:1975:NBS

[Jos75] V. M. Joshi. A new bound for standard p-functions. Annals of Probability, 3(2):346–352, April 1975. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.org/

euclid.aop/1176996405.

Joshi:1977:IUB

[Jos77] V. M. Joshi. An improved upper bound for standard p-functions. Annals of Probability, 5(6):999–1003, December 1977. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http:

//projecteuclid.org/euclid.aop/1176995666.

Jain:1975:OLI

[JP75a] Naresh C. Jain and William E. Pruitt. The other law of the iterated logarithm. Annals of Probability, 3(6):1046–1049, December 1975. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176996232.

Jureckova:1975:ONA

[JP75b] Jana Jureckova and Madan L. Puri. Order of normal approximation for rank test statistics distribution. Annals of Probability, 3(3):526– 533, June 1975. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.aop/
1176996358.

Jacobsen:1977:BDC

[JP77] M. Jacobsen and J. W. Pitman. Birth, death and conditioning of Markov chains. Annals of Probability, 5(3):430–450, June 1977. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176995803.

Jain:1973:LAL

[JT73] N. C. Jain and S. J. Taylor. Local asymptotic laws for Brownian motion. Annals of Probability, 1(4):527–549, August 1973. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http:

//projecteuclid.org/euclid.aop/1176996884.

Kaigh:1975:CLL

[Kai75a] W. D. Kaigh. A conditional local limit theorem for recurrent random walk. Annals of Probability, 3(5):883–888, October 1975. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176996276.

Kaijser:1975:LTP

[Kai75b] Thomas Kaijser. A limit theorem for partially observed Markov chains. Annals of Probability, 3(4):677–696, August 1975. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176996308.

Kaigh:1976:IPR

[Kai76] W. D. Kaigh. An invariance principle for random walk conditioned by a late return to zero. Annals of Probability, 4(1):115–121, February 1976. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.aop/
1176996189.

Kallenberg:1973:CCR

[Kal73] Olav Kallenberg. Conditions for continuity of random processes without discontinuities of the second kind. Annals of Probability, 1(3): 519–526, June 1973. CODEN APBYAE. ISSN 0091-1798 (print),

2168-894X (electronic). URL http://projecteuclid.org/euclid.

aop/1176996948. See correction [Kal76].

Kallenberg:1974:SRP

[Kal74] Olav Kallenberg. Series of random processes without discontinuities of the second kind. Annals of Probability, 2(4):729–737, August 1974. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176996615.

Kallenberg:1975:EPP

[Kal75] Olav Kallenberg. On the existence and path properties of stochastic integrals. Annals of Probability, 3(2):262–280, April 1975. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176996397.

Kallenberg:1976:NCB

[Kal76] Olav Kallenberg. Note: Correction to “Conditions for Continuity of Random Processes Without Discontinuities of the Second Kind”. Annals of Probability, 4(2):338, April 1976. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176996140. See [Kal73].

Kalikow:1977:ELW

[Kal77] Steven Kalikow. An entrance law which reaches equilibrium. Annals of Probability, 5(3):467–469, June 1977. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176995807.

Kallenberg:1978:IVS

[Kal78] Olav Kallenberg. On the independence of velocities in a system of noninteracting particles. Annals of Probability, 6(5):885–890, October 1978. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.aop/
1176995435.

Kanter:1975:CNC

[Kan75a] Marek Kanter. Correction note: Correction to “On Distinguishing Translates of Measures”. Annals of Probability, 3(1):189–190, February 1975. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.aop/
1176996462.

Kanter:1975:SMA

[Kan75b] Marek Kanter. Some martingales associated with sample spacings. Annals of Probability, 3(3):569–572, June 1975. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http:

//projecteuclid.org/euclid.aop/1176996364.

Kanter:1975:SDU

[Kan75c] Marek Kanter. Stable densities under change of scale and total variation inequalities. Annals of Probability, 3(4):697–707, August 1975. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176996309.

Kanter:1976:USD

[Kan76] Marek Kanter. On the unimodality of stable densities. Annals of Probability, 4(6):1006–1008, December 1976. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176995944.

Kanter:1979:LBN

[Kan79] Marek Kanter. Lower bounds for nonlinear prediction error in moving average processes. Annals of Probability, 7(1):128–138, February 1979. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176995153.

Kaplan:1973:MGW

[Kap73] Norman Kaplan. The multitype Galton–Watson process with immigration. Annals of Probability, 1(6):947–953, December 1973. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176996802.

Kaplan:1974:NSB

[Kap74a] Norman Kaplan. A note on the supercritical branching processes with random environments. Annals of Probability, 2(3):509–514, June 1974.

CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176996668.

Kaplan:1974:SRA

[Kap74b] Norman Kaplan. Some results about multidimensional branching processes with random environments. Annals of Probability, 2(3): 441–455, June 1974. CODEN APBYAE. ISSN 0091-1798 (print),

2168-894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176996659.

Kaplan:1975:LTQ

[Kap75] Norman Kaplan. Limit theorems for a GI/G/∞ queue. Annals of Probability, 3(5):780–789, October 1975. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176996265.

Kaplan:1977:TAP

[Kap77] Norman Kaplan. Two applications of a Poisson approximation for dependent events. Annals of Probability, 5(5):787–794, October 1977. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176995720.

Karr:1978:LRM

[Kar78] Alan F. Karr. Lévy random measures. Annals of Probability, 6(1): 57–71, February 1978. CODEN APBYAE. ISSN 0091-1798 (print),

2168-894X (electronic). URL http://projecteuclid.org/euclid.

aop/1176995610. See correction [Kar79].

Karr:1979:NCB

[Kar79] Alan F. Karr. Note: Correction to “Lévy Random Measures”. Annals of Probability, 7(6):1098, December 1979. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http:

//projecteuclid.org/euclid.aop/1176994908. See [Kar78].

Kawada:1979:MLB

[Kaw79] Takayuki Kawada. Maximum in the Lévy–Baxter theorem for Gaussian random fields. Annals of Probability, 7(1):173–178, February 1979. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.aop/
1176995161.

Kim:1979:LDF

[KD79] Geung-Ho Kim and H. T. David. Large deviations of functions of Markovian transitions and mathematical programming duality. Annals of Probability, 7(5):874–881, October 1979. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http:

//projecteuclid.org/euclid.aop/1176994947.

Kent:1978:SPP

[Ken78] John Kent. Some probabilistic properties of Bessel functions. Annals of Probability, 6(5):760–770, October 1978. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176995427.

Kertz:1974:LTD

[Ker74] Robert P. Kertz. Limit theorems for discontinuous random evolutinos with applications to initial value problems and to Markov processes on N lines. Annals of Probability, 2(6):1046–1064, December 1974. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176996497.

Kersting:1977:ASA

[Ker77] Gotz Kersting. Almost sure approximation of the Robbins–Monro process by sums of independent random variables. Annals of Probability, 5(6):954–965, December 1977. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.org/

euclid.aop/1176995663.

Kersting:1978:WCT

[Ker78] Gotz D. Kersting. A weak convergence theorem with application to the Robbins–Monro process. Annals of Probability, 6(6):1015–1025, December 1978. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.aop/
1176995390.

Kesten:1974:RTF

[Kes74] Harry Kesten. Renewal theory for functionals of a Markov chain with general state space. Annals of Probability, 2(3):355–386, June 1974. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176996654.

Kesten:1976:EUC

[Kes76] Harry Kesten. Existence and uniqueness of countable one-dimensional Markov random fields. Annals of Probability, 4(4):557–569, August 1976. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.aop/
1176996027.

Kaplan:1977:JOC

[KF77] Michael Kaplan and Terrence L. Fine. Joint orders in comparative probability. Annals of Probability, 5(2):161–179, April 1977. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176995843.

Klein:1975:QVP

[KG75] Ruben Klein and Evarist Gine. On quadratic variation of processes with Gaussian increments. Annals of Probability, 3(4):716– 721, August 1975. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.aop/
1176996311.

Korwar:1973:CTD

[KH73] Ramesh M. Korwar and Myles Hollander. Contributions to the theory of Dirichlet processes. Annals of Probability, 1(4):705–711, August 1973. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.aop/
1176996898.

Khan:1973:NRC

[Kha73] Rasul A. Khan. A note on the rate of convergence and its applications. Annals of Probability, 1(3):504–508, June 1973. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http:

//projecteuclid.org/euclid.aop/1176996945.

Khalili:1977:OBM

[Kha77] Siamak Khalili. On the oscillation of the Brownian motion random measure. Annals of Probability, 5(1):142–144, February 1977. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176995899.

Kieffer:1973:CPG

[Kie73] J. C. Kieffer. A counterexample to Perez’s generalization of the Shannon–McMillan theorem. Annals of Probability, 1(2):362–364, April 1973. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.aop/

1176996994. See correction [Kie76a].

Kieffer:1974:ASM

[Kie74] J. C. Kieffer. On the approximation of stationary measures by periodic and ergodic measures. Annals of Probability, 2(3):530–534, June 1974. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176996671.

Kieffer:1975:GSM

[Kie75] J. C. Kieffer. A generalized Shannon–McMillan theorem for the action of an amenable group on a probability space. Annals of Probability, 3 (6):1031–1037, December 1975. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.org/

euclid.aop/1176996230.

Kieffer:1976:NCB

[Kie76a] John C. Kieffer. Note: Correction to “A Counterexample to Perez’s Generalization of the Shannon–McMillan Theorem”. Annals of Probability, 4(1):153–154, February 1976. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176996196. See [Kie73].

Kieffer:1976:MNF

[Kie76b] John C. Kieffer. On the minimum number of fixed length sequences with fixed total probability. Annals of Probability, 4(2):335–337, April 1976. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.aop/
1176996139.

Kieffer:1979:ECR

[Kie79] John C. Kieffer. Estimation of a convex real parameter of an unknown information source. Annals of Probability, 7(5):882–886, October 1979. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.aop/
1176994948.

Kingman:1973:SET

[Kin73] J. F. C. Kingman. Subadditive ergodic theory. Annals of Probability, 1(6):883–899, December 1973. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.org/

euclid.aop/1176996798.

Kingman:1975:FBP

[Kin75] J. F. C. Kingman. The first birth problem for an age-dependent branching process. Annals of Probability, 3(5):790–801, October 1975.

CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176996266.

Kingman:1978:UE

[Kin78] J. F. C. Kingman. Uses of exchangeability. Annals of Probability, 6 (2):183–197, April 1978. CODEN APBYAE. ISSN 0091-1798 (print),

2168-894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176995566.

Kiu:1975:TDS

[Kiu75] Sun-Wah Kiu. Two dimensional semi-stable Markov processes. Annals of Probability, 3(3):440–448, June 1975. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176996351.

Kuelbs:1974:BEE

[KK74] J. Kuelbs and T. Kurtz. Berry–Esseen estimates in Hilbert space and an application to the law of the iterated logarithm. Annals of Probability, 2(3):387–407, June 1974. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176996655.

Kamae:1978:SPO

[KK78] T. Kamae and U. Krengel. Stochastic partial ordering. Annals of Probability, 6(6):1044–1049, December 1978. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176995392.

Kamae:1977:SIP

[KKO77] T. Kamae, U. Krengel, and G. L. O’Brien. Stochastic inequalities on partially ordered spaces. Annals of Probability, 5(6):899–912, December 1977. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.aop/
1176995659.

Klass:1973:POE

[Kla73] Michael J. Klass. Properties of optimal extended-valued stopping rules for S_{n}/n^{1}. Annals of Probability, 1(5):719–757, October 1973. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176996843.

Klass:1974:SRE

[Kla74] Michael J. Klass. On stopping rules and the expected supremum of S_{n}/a_{n} and |S_{n}|/a_{n}. Annals of Probability, 2(5):889–905, October 1974. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.aop/
1176996555.

Klein:1976:RTS

[Kle76] Ruben Klein. A representation theorem on stationary Gaussian processes and some local properties. Annals of Probability, 4(5):844–849, October 1976. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.aop/
1176995988.

Klein:1978:GMP

[Kle78] Abel Klein. A generalization of Markov processes. Annals of Probability, 6(1):128–132, February 1978. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.org/

euclid.aop/1176995616.

Klincsek:1977:SFI

[Kli77] G. Klincsek. A square function inequality. Annals of Probability, 5(5): 823–825, October 1977. CODEN APBYAE. ISSN 0091-1798 (print),

2168-894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176995727.

Koopmans:1974:DCR

[KMPQ74] L. H. Koopmans, N. Martin, P. K. Pathak, and C. Qualls. On the divergence of a certain random series. Annals of Probability, 2 (3):546–550, June 1974. CODEN APBYAE. ISSN 0091-1798 (print),

2168-894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176996674.

Keiding:1973:GSB

[KN73] Niels Keiding and John E. Nielsen. The growth of supercritical branching processes with random environments. Annals of Probability, 1 (6):1065–1067, December 1973. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.org/

euclid.aop/1176996814.

Kertz:1979:POP

[KN79] Robert P. Kertz and David C. Nachman. Persistently optimal plans for nonstationary dynamic programming: The topology of weak convergence case. Annals of Probability, 7(5):811–826, October 1979. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176994940.

Knight:1973:LVD

[Kni73] Frank B. Knight. Local variation of diffusion in local time. Annals of Probability, 1(6):1026–1034, December 1973. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http:

//projecteuclid.org/euclid.aop/1176996808.

Knight:1975:PVC

[Kni75] Frank B. Knight. A predictive view of continuous time processes. Annals of Probability, 3(4):573–596, August 1975. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http:

//projecteuclid.org/euclid.aop/1176996302.

Knight:1979:PPA

[Kni79] Frank B. Knight. Prediction processes and an autonomous germ– Markov property. Annals of Probability, 7(3):385–405, June 1979. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176995041.

Kallianpur:1973:NAR

[KO73] G. Kallianpur and H. Oodaira. Non-anticipative representations of equivalent Gaussian processes. Annals of Probability, 1(1):104–122, February 1973. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.aop/
1176997027.

Kostka:1973:KEL

[Kos73] David G. Kostka. On Khintchine’s estimate for large deviations. Annals of Probability, 1(3):509–512, June 1973. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176996946.

Kostka:1974:LCS

[Kos74] David G. Kostka. Lower class sequences for the Skorohod–Strassen approximation scheme. Annals of Probability, 2(6):1172–1178, December 1974. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.aop/
1176996505.

Kaufman:1978:ULI

[KP78] R. Kaufman and Walter Philipp. A uniform law of the iterated logarithm for classes of functions. Annals of Probability, 6(6):930–952, December 1978. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.aop/
1176995385.

Karr:1979:IBP

[KP79] A. F. Karr and A. O. Pittenger. An inverse balayage problem for Brownian motion. Annals of Probability, 7(1):186–191, February 1979. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176995164.

Keilson:1974:MDM

[KS74] Julian Keilson and F. W. Steutel. Mixtures of distributions, moment inequalities and measures of exponentiality and normality. Annals of Probability, 2(1):112–130, February 1974. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http:

//projecteuclid.org/euclid.aop/1176996756.

Kesten:1975:CMC

[KS75] Harry Kesten and Frank Spitzer. Controlled Markov chains. Annals of Probability, 3(1):32–40, February 1975. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176996445.

Kumar:1978:CSC

[KS78] A. Kumar and B. M. Schreiber. Characterization of subclasses of class L probability distributions. Annals of Probability, 6(2):279– 293, April 1978. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.aop/
1176995573.

Komlos:1975:SPH

[KT75] J. Komlos and G. Tusnady. On sequences of “pure heads”. Annals of Probability, 3(4):608–617, August 1975. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176996304.

Klass:1977:ILL

[KT77] Michael Klass and Henry Teicher. Iterated logarithm laws for asymmetric random variables barely with or without finite mean. Annals of Probability, 5(6):861–874, December 1977. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http:

//projecteuclid.org/euclid.aop/1176995656.

Kudo:1974:NSC

[Kud74] Hirokichi Kudo. A note on the strong convergence of Σ-algebras. Annals of Probability, 2(1):76–83, February 1974. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http:

//projecteuclid.org/euclid.aop/1176996752.

Kuelbs:1975:SPB

[Kue75] J. Kuelbs. Sample path behavior for Brownian motion in Banach spaces. Annals of Probability, 3(2):247–261, April 1975. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http:

//projecteuclid.org/euclid.aop/1176996396.

Kuelbs:1976:CBS

[Kue76a] J. Kuelbs. A counterexample for Banach space valued random variables. Annals of Probability, 4(4):684–689, August 1976. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http:

//projecteuclid.org/euclid.aop/1176996039.

Kuelbs:1976:SCT

[Kue76b] J. Kuelbs. A strong convergence theorem for Banach space valued random variables. Annals of Probability, 4(5):744–771, October 1976. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176995982.

Kumar:1973:NCS

[Kum73] A. Kumar. A note on the convergence of stable and class L probability measures on Banach spaces. Annals of Probability, 1(4):716–718, August 1973. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.aop/
1176996900.

Kuntz:1974:GTF

[Kun74] Amy J. Kuntz. Groups of transformations without finite invariant measures have strong generators of size 2. Annals of Probability, 2(1): 143–146, February 1974. CODEN APBYAE. ISSN 0091-1798 (print),

2168-894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176996759.

Kuo:1973:ACM

[Kuo73] Hui-Hsiung Kuo. Absolute continuity of measures corresponding to diffusion processes in Banach space. Annals of Probability, 1(3):513– 518, June 1973. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.aop/
1176996947.

Kurtz:1975:SCS

[Kur75] Thomas G. Kurtz. Semigroups of conditioned shifts and approximation of Markov processes. Annals of Probability, 3(4):618–642, August 1975. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.aop/
1176996305.

Kushner:1974:WCI

[Kus74] Harold J. Kushner. On the weak convergence of interpolated Markov chains to a diffusion. Annals of Probability, 2(1):40–50, February 1974. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176996750.

Kurtz:1973:NYL

[KW73] Thomas G. Kurtz and Stephen Wainger. The nonexistence of the Yaglom limit for an age dependent subcritical branching process. Annals of Probability, 1(5):857–861, October 1973. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http:

//projecteuclid.org/euclid.aop/1176996852.

Kuelbs:1979:SSR

[KZ79] J. Kuelbs and Joel Zinn. Some stability results for vector values random variables. Annals of Probability, 7(1):75–84, February 1979. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176995149.

Lai:1973:GPM

[Lai73] Tze Leung Lai. Gaussian processes, moving averages and quick detection problems. Annals of Probability, 1(5):825–837, October 1973. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176996848.

Lai:1974:LTD

[Lai74a] Tze Leung Lai. Limit theorems for delayed sums. Annals of Probability, 2(3):432–440, June 1974. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.org/

euclid.aop/1176996658.

Lai:1974:MBC

[Lai74b] Tze Leung Lai. Martingales and boundary crossing probabilities for Markov processes. Annals of Probability, 2(6):1152–1167, December 1974. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.aop/
1176996503.

Lai:1976:AMR

[Lai76a] Tze Leung Lai. Asymptotic moments of random walks with applications to ladder variables and renewal theory. Annals of Probability, 4(1):51–66, February 1976. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.org/

euclid.aop/1176996180.

Lai:1976:BCP

[Lai76b] Tze Leung Lai. Boundary crossing probabilities for sample sums and confidence sequences. Annals of Probability, 4(2):299–312, April 1976. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176996135.

Lai:1976:QCC

[Lai76c] Tze Leung Lai. On r-quick convergence and a conjecture of Strassen. Annals of Probability, 4(4):612–627, August 1976. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http:

//projecteuclid.org/euclid.aop/1176996031.

Lai:1976:UTT

[Lai76d] Tze Leung Lai. Uniform Tauberian theorems and their applications to renewal theory and first passage problems. Annals of Probability, 4(4): 628–643, August 1976. CODEN APBYAE. ISSN 0091-1798 (print),

2168-894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176996032.

Lai:1977:CRQ

[Lai77a] Tze Leung Lai. Convergence rates and r-quick versions of the Strong Law for stationary mixing sequences. Annals of Probability, 5(5): 693–706, October 1977. CODEN APBYAE. ISSN 0091-1798 (print),

2168-894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176995713.

Lai:1977:FET

[Lai77b] Tze Leung Lai. First exit times from moving boundaries for sums of independent random variables. Annals of Probability, 5(2):210– 221, April 1977. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.aop/
1176995846.

Lane:1976:PAS

[Lan76] David A. Lane. Purely atomic structures supporting undominated and nonuniformly integrable martingales. Annals of Probability, 4 (6):1016–1019, December 1976. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.org/

euclid.aop/1176995946.

Lane:1978:FSB

[Lan78] David A. Lane. On the fields of some Brownian martingales. Annals of Probability, 6(3):499–508, June 1978. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176995534.

Lindgren:1975:WCH

[LdMR75] Georg Lindgren, Jacques de Mare, and Holger Rootzen. Weak convergence of high level crossings and maxima for one or more Gaussian processes. Annals of Probability, 3(6):961–978, December 1975. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176996222.

LePage:1973:SPR

[Le 73] Raoul Le Page. Subgroups of paths and reproducing kernels. Annals of Probability, 1(2):345–347, April 1973. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176996990.

Lee:1974:RSR

[Lee74] Wang Chung Lee. Random stirring of the real line. Annals of Probability, 2(4):580–592, August 1974. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.org/

euclid.aop/1176996605.

Lehoczky:1977:FSD

[Leh77] John P. Lehoczky. Formulas for stopped diffusion processes with stopping times based on the maximum. Annals of Probability, 5(4):601– 607, August 1977. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.aop/
1176995770.

Lemoine:1974:SPK

[Lem74] Austin J. Lemoine. A simple proof of a known result in random walk theory. Annals of Probability, 2(2):347–348, April 1974. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http:

//projecteuclid.org/euclid.aop/1176996718.

Leslie:1975:GAS

[Les75] Julian Leslie. Generalisation and application of some results of Ibragimov on convergence to normality. Annals of Probability, 3(5):897–902, October 1975. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.aop/
1176996279.

Letac:1975:BRV

[Let75] Gerard Letac. On building random variables of a given distribution. Annals of Probability, 3(2):298–306, April 1975. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http:

//projecteuclid.org/euclid.aop/1176996400.

Letac:1977:RPR

[Let77] Gerard Letac. Recurrence for products of renewal sequences. Annals of Probability, 5(4):591–594, August 1977. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176995768.

Leviatan:1973:MPR

[Lev73] Talma Leviatan. On Markov processes with random starting time. Annals of Probability, 1(2):223–230, April 1973. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176996975.

Leviatan:1975:FMP

[Lev75] Talma Leviatan. Functionals of Markov processes and superprocesses. Annals of Probability, 3(1):41–48, February 1975. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176996446.

Liggett:1973:IPS

[Lig73] Thomas M. Liggett. An infinite particle system with zero range interactions. Annals of Probability, 1(2):240–253, April 1973. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http:

//projecteuclid.org/euclid.aop/1176996977.

Liggett:1974:CTO

[Lig74] Thomas M. Liggett. Convergence to total occupancy in an infinite particle system with interactions. Annals of Probability, 2(6):989– 998, December 1974. CODEN APBYAE. ISSN 0091-1798 (print),

2168-894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176996494.

Liggett:1976:CSE

[Lig76] Thomas M. Liggett. Coupling the simple exclusion process. Annals of Probability, 4(3):339–356, June 1976. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176996084.

Liggett:1977:ETA

[Lig77] Thomas M. Liggett. Ergodic theorems for the asymmetric simple exclusion process II. Annals of Probability, 5(5):795–801, October 1977. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176995721.

Liggett:1978:ANN

[Lig78] Thomas M. Liggett. Attractive nearest neighbor spin systems on the integers. Annals of Probability, 6(4):629–636, August 1978. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176995482.

Ling:1973:ENC

[Lin73] Robert F. Ling. The expected number of components in random linear graphs. Annals of Probability, 1(5):876–881, October 1973. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176996856.

Lin:1974:QCM

[Lin74a] Michael Lin. On quasi-compact Markov operators. Annals of Probability, 2(3):464–475, June 1974. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.org/

euclid.aop/1176996661.

Lindgren:1974:NAI

[Lin74b] Georg Lindgren. A note on the asymptotic independence of high level crossings for dependent Gaussian processes. Annals of Probability, 2 (3):535–539, June 1974. CODEN APBYAE. ISSN 0091-1798 (print),

2168-894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176996672.

Lindvall:1974:ASC

[Lin74c] Torgny Lindvall. Almost sure convergence of branching processes in varying and random environments. Annals of Probability, 2(2):344– 346, April 1974. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.aop/
1176996717.

Lindgren:1975:PRT

[Lin75] Georg Lindgren. Prediction from a random time point. Annals of Probability, 3(3):412–423, June 1975. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176996349.

Lindvall:1977:PPB

[Lin77] Torgny Lindvall. A probabilistic proof of Blackwell’s renewal theorem. Annals of Probability, 5(3):482–485, June 1977. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http:

//projecteuclid.org/euclid.aop/1176995810.

Lloyd:1973:IGR

[LM73] Stuart P. Lloyd and Colin L. Mallows. An index of genealogical relatedness derived from a genetic model. Annals of Probability, 1(5): 758–771, October 1973. CODEN APBYAE. ISSN 0091-1798 (print),

2168-894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176996844.

Logan:1973:LDS

[LMRS73] B. F. Logan, C. L. Mallows, S. O. Rice, and L. A. Shepp. Limit distributions of self-normalized sums. Annals of Probability, 1(5):788– 809, October 1973. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.aop/
1176996846.

Loeve:1973:PL

[Loe73] Michel Loeve. Paul Lévy, 1886–1971. Annals of Probability, 1(1): 1–8, February 1973. CODEN APBYAE. ISSN 0091-1798 (print),

2168-894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176997021.

Loynes:1976:CTS

[Loy76] R. M. Loynes. A criterion for tightness for a sequence of martingales. Annals of Probability, 4(5):859–862, October 1976. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176995990.

Langberg:1979:RGM

[LP79] N. A. Langberg and F. Proschan. A reliability growth model involving dependent components. Annals of Probability, 7(6):1082–1087, December 1979. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.aop/
1176994904.

Langberg:1978:CDM

[LPQ78] N. Langberg, F. Proschan, and A. J. Quinzi. Converting dependent models into independent ones, preserving essential features. Annals of Probability, 6(1):174–181, February 1978. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http:

//projecteuclid.org/euclid.aop/1176995624.

Landers:1976:CCM

[LR76] D. Landers and L. Rogge. Convergence of conditional p-means given a σ-lattice. Annals of Probability, 4(1):147–150, February 1976. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176996194.

Landers:1977:ICN

[LR77a] D. Landers and L. Rogge. Inequalities for conditioned normal approximations. Annals of Probability, 5(4):595–600, August 1977. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176995769.

Landers:1977:CAT

[LR77b] Dieter Landers and Lothar Rogge. A counterexample in the approximation theory of random summation. Annals of Probability, 5(6): 1018–1023, December 1977. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.org/

euclid.aop/1176995669.

Lander:1979:LP

[LR79a] Dieter Lander and Lothar Rogge. On linearity of s-predictors. Annals of Probability, 7(5):887–892, October 1979. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http:

//projecteuclid.org/euclid.aop/1176994949.

Landers:1979:FRB

[LR79b] D. Landers and L. Rogge. A functional relationship between the different r-means for indicator functions. Annals of Probability, 7(1): 166–169, February 1979. CODEN APBYAE. ISSN 0091-1798 (print),

2168-894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176995159.

Levita:1973:LTR

[LS73] Michael L. Levita and Lawrence H. Smolowitz. Limit theorems for reversible Markov processes. Annals of Probability, 1(6):1014–1025, December 1973. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.aop/
1176996807.

Lai:1978:LIL

[LS78] Tze Leung Lai and William Stout. The law of the iterated logarithm and upper-lower class tests for partial sums of stationary Gaussian sequences. Annals of Probability, 6(5):731–750, October 1978. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176995425.

Lai:1979:FET

[LW79] T. L. Lai and R. A. Wijsman. First exit time of a random walk from the bounds f(n)±cg(n), with applications. Annals of Probability, 7(4): 672–692, August 1979. CODEN APBYAE. ISSN 0091-1798 (print),

2168-894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176994990.

Lynch:1978:CCS

[Lyn78] James Lynch. A curious converse of Siever’s theorem. Annals of Probability, 6(1):169–173, February 1978. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176995623.

Lynch:1979:SCD

[Lyn79] James Lynch. Sign changes of the difference of convex functions and their application to large deviation rates. Annals of Probability, 7(1): 96–108, February 1979. CODEN APBYAE. ISSN 0091-1798 (print),

2168-894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176995151.

M:1974:CLT

[M.74] Evarist Gine M. On the Central Limit Theorem for sample continuous processes. Annals of Probability, 2(4):629–641, August 1974. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176996609.

MacQueen:1973:LEM

[Mac73] James B. MacQueen. A linear extension of the martingale convergence theorem. Annals of Probability, 1(2):263–271, April 1973. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176996979.

MacNeill:1978:LPS

[Mac78] Ian B. MacNeill. Limit processes for sequences of partial sums of regression residuals. Annals of Probability, 6(4):695–698, August 1978. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176995491.

Maejima:1978:SVC

[Mae78] Makoto Maejima. Some L_{p} versions for the Central Limit Theorem. Annals of Probability, 6(2):341–344, April 1978. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http:

//projecteuclid.org/euclid.aop/1176995580.

Maisonneuve:1975:ES

[Mai75] Bernard Maisonneuve. Exit systems. Annals of Probability, 3(3): 399–411, June 1975. CODEN APBYAE. ISSN 0091-1798 (print),

2168-894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176996348.

Major:1979:ISI

[Maj79] P. Major. An improvement of Strassen’s invariance principle. Annals of Probability, 7(1):55–61, February 1979. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176995147.

Mallows:1973:BDF

[Mal73] C. L. Mallows. Bounds on distribution functions in terms of expectations of order- statistics. Annals of Probability, 1(2):297–303, April 1973. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.aop/
1176996981.

Maller:1974:SML

[Mal74a] R. A. Maller. On symmetric moments and local behavior of the characteristic function. Annals of Probability, 2(1):173–177, February 1974. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176996765.

Maller:1974:RLB

[Mal74b] R. A. Maller. A remark on local behavior of characteristic functions. Annals of Probability, 2(6):1185–1187, December 1974. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http:

//projecteuclid.org/euclid.aop/1176996507.

Mangano:1976:SCC

[Man76] Gian-Carlo Mangano. Sequential compactness of certain sequences of Gaussian random variables with values in C[0, 1]. Annals of Probability, 4(6):902–913, December 1976. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176995935.

Marcus:1973:CCC

[Mar73a] M. B. Marcus. A comparison of continuity conditions for Gaussian processes. Annals of Probability, 1(1):123–130, February 1973. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176997028.

Marcus:1973:CGP

[Mar73b] M. B. Marcus. Continuity of Gaussian processes and random Fourier series. Annals of Probability, 1(6):968–981, December 1973. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176996804.

Marlow:1973:HLO

[Mar73c] Norman A. Marlow. High level occupation times for continuous Gaussian processes. Annals of Probability, 1(3):388–397, June 1973. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176996933.

Marcus:1974:AMC

[Mar74a] M. B. Marcus. Asymptotic maxima of continuous Gaussian processes. Annals of Probability, 2(4):702–713, August 1974. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176996613.

Marsaglia:1974:APRa

[Mar74b] George Marsaglia. Acknowledgement of priority to: “Random variables with independent binary digits” (Ann. Math. Statist. 42 (1971), 1922–1929). Annals of Probability, 2(4):747, August 1974. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176996619.

Martin:1974:RMC

[Mar74c] N. F. G. Martin. The Rosenblatt mixing condition and Bernoulli shifts. Annals of Probability, 2(2):333–338, April 1974. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http:

//projecteuclid.org/euclid.aop/1176996715.

Marcus:1977:LCS

[Mar77a] Michael B. Marcus. Level crossings of a stochastic process with absolutely continuous sample paths. Annals of Probability, 5(1):52–71, February 1977. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.aop/
1176995890.

Marcus:1977:RHD

[Mar77b] Michael B. Marcus. Review: H. Dym, H. P. McKean, Gaussian Processes, Function Theory and the Inverse Spectral Problem. Annals of Probability, 5(6):1047–1051, December 1977. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http:

//projecteuclid.org/euclid.aop/1176995674.

Matthes:1975:MCL

[Mat75] T. K. Matthes. The multivariate Central Limit Theorem for regular convex sets. Annals of Probability, 3(3):503–515, June 1975. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176996356.

Matloff:1977:ECD

[Mat77] Norman S. Matloff. Ergodicity conditions for a dissonant voting model. Annals of Probability, 5(3):371–386, June 1977. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http:

//projecteuclid.org/euclid.aop/1176995798.

Madsen:1973:EBN

[MC73] Richard W. Madsen and Patricia S. Conn. Ergodic behavior for nonnegative kernels. Annals of Probability, 1(6):995–1013, December 1973. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.aop/
1176996806.

McDonald:1977:EMS

[McD77] David R. McDonald. Equilibrium measures for semi-Markov processes. Annals of Probability, 5(5):818–822, October 1977. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176995726.

McDonald:1978:SMS

[McD78] David McDonald. On semi-Markov and semiregenerative processes II. Annals of Probability, 6(6):995–1014, December 1978. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http:

//projecteuclid.org/euclid.aop/1176995389.

McDonald:1979:LLT

[McD79] David McDonald. A local limit theorem for large deviations of sums of independent, nonidentically distributed random variables. Annals of Probability, 7(3):526–531, June 1979. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176995052.

McKean:1973:GDS

[McK73] H. P. McKean. Geometry of differential space. Annals of Probability, 1 (2):197–206, April 1973. CODEN APBYAE. ISSN 0091-1798 (print),

2168-894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176996973.

McLeish:1974:DCL

[McL74] D. L. McLeish. Dependent Central Limit Theorems and invariance principles. Annals of Probability, 2(4):620–628, August 1974. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176996608.

McLeish:1975:MID

[McL75] D. L. McLeish. A maximal inequality and dependent Strong Laws. Annals of Probability, 3(5):829–839, October 1975. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http:

//projecteuclid.org/euclid.aop/1176996269.

McLeish:1977:IPN

[McL77] D. L. McLeish. On the invariance principle for nonstationary mixingales. Annals of Probability, 5(4):616–621, August 1977. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http:

//projecteuclid.org/euclid.aop/1176995772.

McLeish:1978:EMI

[McL78] D. L. McLeish. An extended martingale invariance principle. Annals of Probability, 6(1):144–150, February 1978. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http:

//projecteuclid.org/euclid.aop/1176995619.

Meyer:1973:PTL

[Mey73] R. M. Meyer. A Poisson-type limit theorem for mixing sequences of dependent ’rare’ events. Annals of Probability, 1(3):480–483, June 1973. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176996941.

Meyer:1975:EDD

[Mey75] R. M. Meyer. The extent of discrete distributions. Annals of Probability, 3(1):166–168, February 1975. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.org/

euclid.aop/1176996457.

Madsen:1973:SEB

[MI73] Richard W. Madsen and Dean L. Isaacson. Strongly ergodic behavior for non-stationary Markov processes. Annals of Probability, 1(2):329– 335, April 1973. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.aop/
1176996986.

Michel:1974:RPM

[Mic74] R. Michel. Results on probabilities of moderate deviations. Annals of Probability, 2(2):349–353, April 1974. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176996719.

Michel:1976:NCL

[Mic76] R. Michel. Nonuniform central limit bounds with applications to probabilities of deviations. Annals of Probability, 4(1):102–106, February 1976. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.aop/
1176996186.

Millar:1973:RP

[Mil73] P. W. Millar. Radial processes. Annals of Probability, 1(4):613–626, August 1973. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.aop/
1176996890.

Millar:1975:FPD

[Mil75] P. W. Millar. First passage distributions of processes with independent increments. Annals of Probability, 3(2):215–233, April 1975. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176996394.

Millar:1978:PDM

[Mil78] P. W. Millar. A path decomposition for Markov processes. Annals of Probability, 6(2):345–348, April 1978. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176995581.

Miller:1979:IDS

[Mil79] H. D. Miller. Infinite divisibility in stochastic processes. Annals of Probability, 7(3):406–417, June 1979. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176995042.

Mineka:1974:SLL

[Min74] J. Mineka. A stable local limit theorem. Annals of Probability, 2(1): 167–172, February 1974. CODEN APBYAE. ISSN 0091-1798 (print),

2168-894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176996764.

Mittal:1974:LBM

[Mit74] Yash Mittal. Limiting behavior of maxima in stationary Gaussian sequences. Annals of Probability, 2(2):231–242, April 1974. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176996705.

Mittal:1978:MPS

[Mit78] Yashaswini Mittal. Maxima of partial samples in Gaussian sequences. Annals of Probability, 6(3):421–432, June 1978. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176995528.

Mittal:1979:NMC

[Mit79] Yashaswini Mittal. A new mixing condition for stationary Gaussian processes. Annals of Probability, 7(4):724–730, August 1979. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176994993.

Mode:1978:RJP

[Mod78] Charles J. Mode. Review: Jagers P., Branching Processes with Biological Applications. Annals of Probability, 6(4):701–704, August 1978. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176995493.

Mohan:1976:TRT

[Moh76] N. R. Mohan. Teugels’ renewal theorem and stable laws. Annals of Probability, 4(5):863–868, October 1976. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176995991.

Moncayo:1975:PCF

[Mon75] A. R. Moncayo. Poisson convergence and family trees. Annals of Probability, 3(6):1059–1061, December 1975. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176996235.

Monroe:1976:ASC

[Mon76] Itrel Monroe. Almost sure convergence of the quadratic variation of martingales: A counterexample. Annals of Probability, 4(1):133–138, February 1976. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.aop/
1176996192.

Monrad:1977:SPS

[Mon77] Ditlev Monrad. Stable processes: Sample function growth at a last exit time. Annals of Probability, 5(3):455–462, June 1977. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176995805.

Monrad:1978:ACP

[Mon78a] Ditlev Monrad. Asymmetric Cauchy processes: Sample functions at last zero. Annals of Probability, 6(5):771–787, October 1978. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176995428.

Monroe:1978:PCE

[Mon78b] Itrel Monroe. Processes that can be embedded in Brownian motion. Annals of Probability, 6(1):42–56, February 1978. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http:

//projecteuclid.org/euclid.aop/1176995609.

Mori:1976:LLM

[Mor76] Toshio Mori. Limit laws for maxima and second maxima from strongmixing processes. Annals of Probability, 4(1):122–126, February 1976. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176996190.

Mountford:1979:IF

[Mou79] David Mountford. An inequality in p-functions. Annals of Probability, 7(1):184–185, February 1979. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.org/

euclid.aop/1176995163.

Mehra:1975:WCG

[MR75] K. L. Mehra and M. Sudhakara Rao. Weak convergence of generalized empirical processes relative to d_{q} under strong mixing. Annals of Probability, 3(6):979–991, December 1975. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176996223.

Miller:1975:TRP

[MS75] Douglas R. Miller and F. Dennis Sentilles. Translated renewal processes and the existence of a limiting distribution for the queue length of the GI/G/s queue. Annals of Probability, 3(3):424–439, June 1975. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176996350.

Marshall:1979:MSM

[MS79] Albert W. Marshall and Moshe Shaked. Multivariate shock models for distributions with increasing hazard rate average. Annals of Probability, 7(2):343–358, April 1979. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.org/

euclid.aop/1176995092.

Marsaglia:1975:NLM

[MT75] George Marsaglia and Alberto Tubilla. A note on the “lack of memory” property of the exponential distribution. Annals of Probability, 3 (2):353–354, April 1975. CODEN APBYAE. ISSN 0091-1798 (print),

2168-894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176996406.

Mucci:1973:CSP

[Muc73] Anthony G. Mucci. On a class of secretary problems. Annals of Probability, 1(3):417–427, June 1973. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176996936.

Mittal:1976:SLM

[MY76] Yash Mittal and Donald Ylvisaker. Strong Laws for the maxima of stationary Gaussian processes. Annals of Probability, 4(3):357– 371, June 1976. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.aop/
1176996085.

Myers:1974:RWN

[Mye74] Lawrence E. Myers. A random walk with nearly uniform N-step motion. Annals of Probability, 2(1):32–39, February 1974. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http:

//projecteuclid.org/euclid.aop/1176996749.

Myers:1976:CTM

[Mye76] L. E. Myers. Central terms of Markov walks. Annals of Probability, 4 (2):313–318, April 1976. CODEN APBYAE. ISSN 0091-1798 (print),

2168-894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176996136.

Nummelin:1976:DCI

[NA76] Esa Nummelin and Elja Arjas. A direct construction of the Rinvariant measure for a Markov chain on a general state space. Annals of Probability, 4(4):674–679, August 1976. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http:

//projecteuclid.org/euclid.aop/1176996037.

Nagaev:1979:LDS

[Nag79] S. V. Nagaev. Large deviations of sums of independent random variables. Annals of Probability, 7(5):745–789, October 1979. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176994938.

Nair:1977:CCD

[Nai77] K. Aiyappan Nair. On complete convergence of distributions and expected values of order statistics. Annals of Probability, 5(5):826– 827, October 1977. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.aop/
1176995728.

Neaderhouser:1978:LTM

[Nea78] Carla C. Neaderhouser. Limit theorems for multiply indexed mixing random variables, with application to Gibbs random fields. Annals of Probability, 6(2):207–215, April 1978. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176995568.

Nevison:1976:PMP

[Nev76] Christopher H. Nevison. Potentials of Markov processes without duality. Annals of Probability, 4(3):497–501, June 1976. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http:

//projecteuclid.org/euclid.aop/1176996101.

Norman:1974:CLT

[Nor74] M. Frank Norman. A central limit theorem for Markov processes that move by small steps. Annals of Probability, 2(6):1065–1074, December 1974. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.aop/
1176996498.

Norman:1975:DAN

[Nor75] M. Frank Norman. Diffusion approximation of non-Markovian processes. Annals of Probability, 3(2):358–364, April 1975. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http:

//projecteuclid.org/euclid.aop/1176996408.

Naus:1975:EVN

[NR75] Joseph I. Naus and Larry Rabinowitz. The expectation and variance of the number of components in random linear graphs. Annals of Probability, 3(1):159–161, February 1975. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176996455.

Nummelin:1978:GER

[NT78] E. Nummelin and R. L. Tweedie. Geometric ergodicity and rpositivity for general Markov chains. Annals of Probability, 6(3):

404–420, June 1978. CODEN APBYAE. ISSN 0091-1798 (print),

2168-894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176995527.

Nummelin:1979:SRL

[Num79] E. Nummelin. Strong ratio limit theorems for ϕ-recurrent Markov chains. Annals of Probability, 7(4):639–650, August 1979. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176994987.

Oakes:1976:ROI

[Oak76] David Oakes. Random overlapping intervals — a generalization of Erlang’s loss formula. Annals of Probability, 4(6):940–946, December 1976. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.aop/
1176995938.

OBrien:1974:LTM

[O’B74a] G. L. O’Brien. Limit theorems for the maximum term of a stationary process. Annals of Probability, 2(3):540–545, June 1974. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http:

//projecteuclid.org/euclid.aop/1176996673.

OBrien:1974:LDM

[O’B74b] George O’Brien. The limiting distribution of maxima of random variables defined on a denumerable Markov chain. Annals of Probability, 2(1):103–111, February 1974. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.org/

euclid.aop/1176996755.

OBrien:1975:CMS

[O’B75] G. L. O’Brien. The comparison method for stochastic processes. Annals of Probability, 3(1):80–88, February 1975. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http:

//projecteuclid.org/euclid.aop/1176996450.

OConnor:1979:IDD

[O’C79] Thomas A. O’Connor. Infinitely divisible distributions with unimodal Lévy spectral functions. Annals of Probability, 7(3):494–499, June 1979. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.aop/
1176995049.

Oman:1977:RLT

[Oma77] Samuel D. Oman. A ratio limit theorem for subterminal times. Annals of Probability, 5(2):262–277, April 1977. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176995850.

Oodaira:1973:LIL

[Ood73] Hiroshi Oodaira. The law of the iterated logarithm for Gaussian processes. Annals of Probability, 1(6):954–967, December 1973. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176996803.

Orey:1973:SFP

[OP73] Steven Orey and William E. Pruitt. Sample functions of the Nparameter Wiener process. Annals of Probability, 1(1):138–163, February 1973. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.aop/
1176997030.

OReilly:1974:SRC

[O’R74a] Neville E. O’Reilly. On Sawyer’s rates of convergence for some functionals in probability. Annals of Probability, 2(6):1179–1184, December 1974. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.aop/
1176996506.

OReilly:1974:WCE

[O’R74b] Neville E. O’Reilly. On the weak convergence of empirical processes in sup-norm metrics. Annals of Probability, 2(4):642–651, August 1974. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176996610.

Ornstein:1973:DPO

[Orn73a] Donald Ornstein. Discussion on Professor Ornstein’s paper. Annals of Probability, 1(1):59–65, February 1973. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176997025.

Ornstein:1973:AET

[Orn73b] Donald S. Ornstein. An application of ergodic theory to probability theory. Annals of Probability, 1(1):43–58, February 1973. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176997024.

Owen:1973:EVD

[Owe73] Willis L. Owen. An estimate for E(|S_{n}|) for variables in the domain of normal attraction of a stable law of index α, 1 < α < 2. Annals of Probability, 1(6):1071–1073, December 1973. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http:

//projecteuclid.org/euclid.aop/1176996816.

Pfanzagl:1974:CCE

[Pfa74] J. Pfanzagl. Convexity and conditional expectations. Annals of Probability, 2(3):490–494, June 1974. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.org/

euclid.aop/1176996665.

Pfanzagl:1979:CDD

[Pfa79] P. Pfanzagl. Conditional distributions as derivatives. Annals of Probability, 7(6):1046–1050, December 1979. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176994897.

Pursley:1977:SCT

[PG77] Michael B. Pursley and Robert M. Gray. Source coding theorems for stationary, continuous-time stochastic processes. Annals of Probability, 5(6):966–986, December 1977. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176995664.

Philipp:1977:FLI

[Phi77] Walter Philipp. A functional law of the iterated logarithm for empirical distribution functions of weakly dependent random variables. Annals of Probability, 5(3):319–350, June 1977. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http:

//projecteuclid.org/euclid.aop/1176995795.

Phoha:1978:ALP

[Pho78] Shashi Phoha. An algorithm for linear prediction of a Banach space valued stationary stochastic process. Annals of Probability, 6(5):891– 898, October 1978. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.aop/
1176995436.

Pitt:1977:GCI

[Pit77] Loren D. Pitt. A Gaussian correlation inequality for symmetric convex sets. Annals of Probability, 5(3):470–474, June 1977. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http:

//projecteuclid.org/euclid.aop/1176995808.

Pittenger:1979:NSF

[Pit79] A. O. Pittenger. Note on a square function inequality. Annals of Probability, 7(5):907–908, October 1979. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176994952.

Papantoni-Kazakos:1979:RES

[PKG79] P. Papantoni-Kazakos and Robert M. Gray. Robustness of estimators on stationary observations. Annals of Probability, 7(6):989–1002, December 1979. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.aop/
1176994892.

Portnoy:1975:TSN

[Por75] Stephen L. Portnoy. Transience and solvability of a non-linear diffusion equation. Annals of Probability, 3(3):465–477, June 1975. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176996353.

Portnoy:1978:PBF

[Por78] Stephen Portnoy. Probability bounds for first exits through moving boundaries. Annals of Probability, 6(1):106–117, February 1978. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176995614.

Posner:1973:EES

[PR73] Edward C. Posner and Eugene R. Rodemich. Epsilon entropy of stochastic processes with continuous paths. Annals of Probability, 1(4):674–689, August 1973. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.org/

euclid.aop/1176996894.

Puri:1974:CFD

[PR74] Prem S. Puri and Herman Rubin. On a characterization of the family of distributions with constant multivariate failure rates. Annals of Probability, 2(4):738–740, August 1974. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176996616.

Pratelli:1975:DIC

[Pra75] Maurizio Pratelli. Deux inégalités concernant les opérateurs de Burkholder sur les martingales. (French) [Two inequalities concerning Burkholder operators on martingales]. Annals of Probability, 3 (2):365–370, April 1975. CODEN APBYAE. ISSN 0091-1798 (print),

2168-894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176996409.

Protter:1977:EUC

[Pro77] Philip E. Protter. On the existence, uniqueness, convergence and explosions of solutions of systems of stochastic integral equations. Annals of Probability, 5(2):243–261, April 1977. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176995849.

Protter:1979:CSI

[Pro79] Philip Protter. A comparison of stochastic integrals. Annals of Probability, 7(2):276–289, April 1979. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.org/

euclid.aop/1176995088.

Pyke:1975:CNC

[PS75] Ronald Pyke and Galen Shorack. Correction note: Correction to “Weak Convergence of a Two-sample Empirical Process and a New Approach to Chernoff–Savage Theorems”. Annals of Probability, 3(6): 1068, December 1975. CODEN APBYAE. ISSN 0091-1798 (print),

2168-894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176996237.

Purves:1976:SFA

[PS76] Roger A. Purves and William D. Sudderth. Some finitely additive probability. Annals of Probability, 4(2):259–276, April 1976. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176996133.

Port:1977:SDA

[PS77] Sidney C. Port and Charles J. Stone. Spacing distribution associated with a stationary random measure on the real line. Annals of Probability, 5(3):387–394, June 1977. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.org/

euclid.aop/1176995799.

Protter:1979:MGA

[PS79] P. Protter and M. J. Sharpe. Martingales with given absolute value. Annals of Probability, 7(6):1056–1058, December 1979. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http:

//projecteuclid.org/euclid.aop/1176994899.

Port:1975:SCI

[PSW75] S. C. Port, C. J. Stone, and N. A. Weiss. SLLNs and CLTs for infinite particle systems. Annals of Probability, 3(5):753–761, October 1975. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176996262.

Pruitt:1977:HMP

[PT77] William E. Pruitt and S. James Taylor. Hausdorff measure properties of the asymmetric Cauchy processes. Annals of Probability, 5(4):608– 615, August 1977. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.aop/
1176995771.

Pitt:1979:LSP

[PT79] Loren D. Pitt and Lanh Tat Tran. Local sample path properties of Gaussian fields. Annals of Probability, 7(3):477–493, June 1979. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176995048.

Puterman:1974:SDO

[Put74] Martin L. Puterman. Sensitive discount optimality in controlled one-dimensional diffusions. Annals of Probability, 2(3):408–419, June 1974. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.aop/
1176996656.

Qualls:1977:LIL

[Qua77] Clifford Qualls. The law of the iterated logarithm on arbitrary sequences for stationary Gaussian processes and Brownian motion. Annals of Probability, 5(5):724–739, October 1977. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http:

//projecteuclid.org/euclid.aop/1176995715.

Quine:1976:ARE

[Qui76] M. P. Quine. Asymptotic results for estimators in a subcritical branching process with immigration. Annals of Probability, 4(2):319– 325, April 1976. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.aop/

1176996137. See correction [Qui77].

Quine:1977:NCB

[Qui77] M. P. Quine. Notes: Correction to “Asymptotic Results for Estimators in a Subcritical Branching Process with Immigration”. Annals of Probability, 5(2):318, April 1977. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.org/

euclid.aop/1176995859. See [Qui76].

Ramachandran:1974:MPP

[Ram74] D. Ramachandran. Mixtures of perfect probability measures. Annals of Probability, 2(3):495–500, June 1974. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176996666.

Ramachandran:1979:EIC

[Ram79a] D. Ramachandran. Existence of independent complements in regular conditional probability spaces. Annals of Probability, 7(3):433– 443, June 1979. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.aop/
1176995044.

Ramachandran:1979:PMP

[Ram79b] D. Ramachandran. Perfect mixtures of perfect measures. Annals of Probability, 7(3):444–452, June 1979. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176995045.

Reiss:1974:ANA

[Rei74] R.-D. Reiss. On the accuracy of the normal approximation for quantiles. Annals of Probability, 2(4):741–744, August 1974. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http:

//projecteuclid.org/euclid.aop/1176996617.

Reiss:1976:AES

[Rei76] R.-D. Reiss. Asymptotic expansions for sample quantiles. Annals of Probability, 4(2):249–258, April 1976. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176996132.

Reich:1977:IET

[Rei77] Jakob I. Reich. On the individual ergodic theorem for subsequences. Annals of Probability, 5(6):1039–1046, December 1977. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http:

//projecteuclid.org/euclid.aop/1176995673.

Resnick:1973:RVM

[Res73] Sidney I. Resnick. Record values and maxima. Annals of Probability, 1(4):650–662, August 1973. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.org/

euclid.aop/1176996892.

Ressel:1974:WCC

[Res74] Paul Ressel. Weak convergence of certain vector-valued measures. Annals of Probability, 2(1):136–142, February 1974. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http:

//projecteuclid.org/euclid.aop/1176996758.

Resnick:1975:WCE

[Res75] Sidney I. Resnick. Weak convergence to extremal processes. Annals of Probability, 3(6):951–960, December 1975. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http:

//projecteuclid.org/euclid.aop/1176996221.

Revesz:1976:SAM

[Rev76] P. Revesz. On strong approximation of the multidimensional empirical process. Annals of Probability, 4(5):729–743, October 1976. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176995981.

Rinott:1976:CM

[Rin76] Yosef Rinott. On convexity of measures. Annals of Probability, 4 (6):1020–1026, December 1976. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.org/

euclid.aop/1176995947.

Ripley:1976:FSG

[Rip76a] B. D. Ripley. The foundations of stochastic geometry. Annals of Probability, 4(6):995–998, December 1976. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176995942.

Ripley:1976:LFR

[Rip76b] B. D. Ripley. Locally finite random sets: Foundations for point process theory. Annals of Probability, 4(6):983–994, December 1976. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176995941.

Ripley:1976:SSP

[Rip76c] B. D. Ripley. On stationarity and superposition of point processes. Annals of Probability, 4(6):999–1005, December 1976. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http:

//projecteuclid.org/euclid.aop/1176995943.

Roberts:1973:NFA

[Rob73] Fred S. Roberts. A note on Fine’s axioms for qualitative probability. Annals of Probability, 1(3):484–487, June 1973. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176996942. See correc-

tion [Rob74].

Roberts:1974:CNC

[Rob74] Fred S. Roberts. Correction note: Correction to “A Note on Fine’s Axioms for Qualitative Probability”. Annals of Probability, 2(1):182, February 1974. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.aop/

1176996768. See [Rob73].

Robbins:1976:CSE

[Rob76] Naomi B. Robbins. Convergence of some expected first passage times. Annals of Probability, 4(6):1027–1029, December 1976. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http:

//projecteuclid.org/euclid.aop/1176995948.

Robinson:1977:LDP

[Rob77] J. Robinson. Large deviation probabilities for samples from a finite population. Annals of Probability, 5(6):913–925, December 1977. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176995660. See correc-

tion [Rob83].

Robinson:1983:CCB

[Rob83] J. Robinson. Corrections: Correction to “Large Deviation Probabilities for Samples from a Finite Population”. Annals of Probability, 11 (4):1055, November 1983. CODEN APBYAE. ISSN 0091-1798 (print),

2168-894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176993456. See [Rob77].

Rodman:1978:MAB

[Rod78] Leiba Rodman. On the many-armed bandit problem. Annals of Probability, 6(3):491–498, June 1978. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.org/

euclid.aop/1176995533.

Rogge:1974:UIC

[Rog74] L. Rogge. Uniform inequalities for conditional expectations. Annals of Probability, 2(3):486–489, June 1974. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176996664.

Rootzen:1976:FSW

[Roo76] Holger Rootzen. Fluctuations of sequences which converge in distribution. Annals of Probability, 4(3):456–463, June 1976. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http:

//projecteuclid.org/euclid.aop/1176996094.

Rootzen:1978:EMA

[Roo78] Holger Rootzen. Extremes of moving averages of stable processes. Annals of Probability, 6(5):847–869, October 1978. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http:

//projecteuclid.org/euclid.aop/1176995432.

Rosenkrantz:1973:MCA

[Ros73] Walter A. Rosenkrantz. A method for computing the asymptotic limit of a class of expected first passage times. Annals of Probability, 1(6):1035–1043, December 1973. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.org/

euclid.aop/1176996809.

Rosenblatt:1976:MDD

[Ros76] M. Rosenblatt. On the maximal deviation of k-dimensional density estimates. Annals of Probability, 4(6):1009–1015, December 1976. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176995945.

Rosenblatt:1979:SRM

[Ros79a] M. Rosenblatt. Some remarks on a mixing condition. Annals of Probability, 7(1):170–172, February 1979. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176995160. See correction [Ros79b].

Rosenblatt:1979:NCB

[Ros79b] Murray Rosenblatt. Note: Correction to “Some Remarks on a Mixing Condition”. Annals of Probability, 7(6):1097, December 1979. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176994907. See

[Ros79a].

Rosler:1979:TFT

[Ros79c] Uwe Rosler. The tail σ-field of time-homogeneous one-dimensional diffussion process. Annals of Probability, 7(5):847–857, October 1979. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176994943.

Ross:1979:MSC

[Ros79d] Sheldon M. Ross. Multivalued state component systems. Annals of Probability, 7(2):379–383, April 1979. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176995096.

Root:1973:PPN

[RR73] D. Root and H. Rubin. A probabilistic proof of the normal convergence criterion. Annals of Probability, 1(5):867–869, October 1973. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176996854.

Rosen:1976:FP

[RS76] Jay Rosen and Barry Simon. Fluctuations in P (ϕ)_{1} processes. Annals of Probability, 4(2):155–174, April 1976. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176996128.

Rubin:1977:FMS

[RS77] H. Rubin and S. M. Samuels. The finite-memory secretary problem. Annals of Probability, 5(4):627–635, August 1977. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http:

//projecteuclid.org/euclid.aop/1176995774.

Rubinovitch:1977:CNS

[Rub77] Michael Rubinovitch. Can a nonstable state become stable by subordination? Annals of Probability, 5(3):463–466, June 1977. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176995806.

Rudin:1973:LRT

[Rud73] Walter Rudin. Limits of ratios of tails of measures. Annals of Probability, 1(6):982–994, December 1973. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176996805.

Ryan:1976:MRT

[Rya76] Thomas A. Ryan. A multidimensional renewal theorem. Annals of Probability, 4(4):656–661, August 1976. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176996034.

Salehi:1979:ALI

[Sal79] H. Salehi. Algorithms for linear interpolator and interpolation error for minimal stationary stochastic processes. Annals of Probability, 7(5):840–846, October 1979. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.org/

euclid.aop/1176994942.

Savits:1975:AST

[Sav75] Thomas H. Savits. Application of space-time harmonic functions to branching processes. Annals of Probability, 3(1):61–69, February 1975. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176996448.

Savits:1976:ADM

[Sav76] Thomas H. Savits. An age-dependent model with parental survival. Annals of Probability, 4(3):382–392, June 1976. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http:

//projecteuclid.org/euclid.aop/1176996087.

Sawyer:1976:RSS

[Saw76] Stanley Sawyer. Results for the stepping stone model for migration in population genetics. Annals of Probability, 4(5):699–728, October 1976. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.aop/
1176995980.

Sawyer:1977:RCS

[Saw77] Stanley Sawyer. Rates of consolidation in a selectively neutral migration model. Annals of Probability, 5(3):486–493, June 1977. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176995811.

Siddiqui:1975:JDR

[SB75] M. M. Siddiqui and Ronald W. Biondini. The joint distribution of record values and inter-record times. Annals of Probability, 3(6):1012– 1013, December 1975. CODEN APBYAE. ISSN 0091-1798 (print),

2168-894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176996226.

Samuel-Cahn:1974:ADO

[SC74] Ester Samuel-Cahn. Asymptotic distributions for occupancy and waiting time problems with positive probability of falling through the cells. Annals of Probability, 2(3):515–521, June 1974. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http:

//projecteuclid.org/euclid.aop/1176996669.

Schay:1974:NRV

[Sch74] Geza Schay. Nearest random variables with given distributions. Annals of Probability, 2(1):163–166, February 1974. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http:

//projecteuclid.org/euclid.aop/1176996763.

Schaefer:1976:NDJ

[Sch76a] Martin Schaefer. Note on the k-dimensional Jensen inequality. Annals of Probability, 4(3):502–504, June 1976. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176996102.

Schwartz:1976:ETI

[Sch76b] Diane Schwartz. Ergodic theorems for an infinite particle system with births and deaths. Annals of Probability, 4(5):783–801, October 1976. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176995984.

Schwarz:1976:NCS

[Sch76c] Gideon Schwarz. Noninvariance of d̄-convergence of k-step Markov approximations. Annals of Probability, 4(6):1033–1035, December 1976.

CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176995950.

Schassberger:1977:ISS

[Sch77a] R. Schassberger. Insensitivity of steady-state distributions of generalized semi-Markov processes. Part I. Annals of Probability, 5(1): 87–99, February 1977. CODEN APBYAE. ISSN 0091-1798 (print),

2168-894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176995893.

Schwartz:1977:ADC

[Sch77b] Diane L. Schwartz. Applications of duality to a class of Markov processes. Annals of Probability, 5(4):522–532, August 1977. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176995758.

Schassberger:1978:ISS

[Sch78a] R. Schassberger. Insensitivity of steady-state distributions of generalized semi-Markov processes. Part II. Annals of Probability, 6(1): 85–93, February 1978. CODEN APBYAE. ISSN 0091-1798 (print),

2168-894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176995612.

Schwartz:1978:HPA

[Sch78b] Diane Schwartz. On hitting probabilities for an annihilating particle model. Annals of Probability, 6(3):398–403, June 1978. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http:

//projecteuclid.org/euclid.aop/1176995526.

Sen:1973:ASI

[Sen73a] Pranab Kumar Sen. An almost sure invariance principle for mutivariate Kolmogorov–Smirnov statistics. Annals of Probability, 1(3): 488–496, June 1973. CODEN APBYAE. ISSN 0091-1798 (print),

2168-894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176996943.

Seneta:1973:TTL

[Sen73b] E. Seneta. A Tauberian theorem of E. Landau and W. Feller. Annals of Probability, 1(6):1057–1058, December 1973. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http:

//projecteuclid.org/euclid.aop/1176996811.

Sen:1974:WCG

[Sen74a] Pranab Kumar Sen. Weak convergence of generalized U-statistics. Annals of Probability, 2(1):90–102, February 1974. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http:

//projecteuclid.org/euclid.aop/1176996754.

Sen:1974:WCM

[Sen74b] Pranab Kumar Sen. Weak convergence of multidimensional empirical processes for stationary ϕ-mixing processes. Annals of Probability, 2(1):147–154, February 1974. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.org/

euclid.aop/1176996760.

Sendler:1975:NPZ

[Sen75] Wolfgang Sendler. A note on the proof of the zero-one law of Blum and Pathak. Annals of Probability, 3(6):1055–1058, December 1975. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176996234.

Sen:1976:ASI

[Sen76a] Pranab Kumar Sen. An almost sure invariance principle for the extrema of certain sample functions. Annals of Probability, 4(1):81–88, February 1976. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.aop/
1176996182.

Sen:1976:NIP

[Sen76b] Pranab Kumar Sen. A note on invariance principles for induced order statistics. Annals of Probability, 4(3):474–479, June 1976. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176996097.

Sen:1976:TDF

[Sen76c] Pranab Kumar Sen. A two-dimensional functional permutational Central Limit Theorem for linear rank statistics. Annals of Probability, 4 (1):13–26, February 1976. CODEN APBYAE. ISSN 0091-1798 (print),

2168-894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176996177.

Sen:1977:ASC

[Sen77] Pranab Kumar Sen. Almost sure convergence of generalized Ustatistics. Annals of Probability, 5(2):287–290, April 1977. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176995853.

Serfozo:1973:WCS

[Ser73] Richard F. Serfozo. Weak convergence of superpositions of randomly selected partial sums. Annals of Probability, 1(6):1044–1056, December 1973. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.aop/
1176996810.

Serfling:1975:GPA

[Ser75] R. J. Serfling. A general Poisson approximation theorem. Annals of Probability, 3(4):726–731, August 1975. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176996313.

Shanbhag:1975:NDI

[SG75] D. N. Shanbhag and J. Gani. A new decomposition of infinitely divisible distributions. Annals of Probability, 3(2):338–345, April 1975. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176996404.

Shafer:1979:AP

[Sha79] Glenn Shafer. Allocations of probability. Annals of Probability, 7(5): 827–839, October 1979. CODEN APBYAE. ISSN 0091-1798 (print),

2168-894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176994941.

Sheu:1974:SIL

[She74] Shey Shiung Sheu. Some iterated logarithm results for sums of independent two-dimensional random variables. Annals of Probability, 2 (6):1139–1151, December 1974. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.org/

euclid.aop/1176996502.

Shepp:1977:NCB

[She77] L. A. Shepp. Notes: Correction to “Radon–Nikodym Derivatives of Gaussian Measures”. Annals of Probability, 5(2):315–317, April 1977. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176995858.

Shore:1978:LLT

[Sho78] Terence R. Shore. On the local limit theorem for independent nonlattice random variables. Annals of Probability, 6(4):563–573, August 1978. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.aop/
1176995478.

Shorack:1979:EDE

[Sho79] Galen R. Shorack. Extension of the Darling and Erdős theorem on the maximum of normalized sums. Annals of Probability, 7(6):1092– 1096, December 1979. CODEN APBYAE. ISSN 0091-1798 (print),

2168-894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176994906.

Siebert:1976:CCP

[Sie76a] Eberhard Siebert. Convergence and convolutions of probability measures on a topological group. Annals of Probability, 4(3):433–443, June 1976. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.aop/
1176996091.

Siegmund:1976:EAR

[Sie76b] D. Siegmund. The equivalence of absorbing and reflecting barrier problems for stochastically monotone Markov processes. Annals of Probability, 4(6):914–924, December 1976. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http:

//projecteuclid.org/euclid.aop/1176995936.

Siegel:1979:ACD

[Sie79] Andrew F. Siegel. Asymptotic coverage distributions on the circle. Annals of Probability, 7(4):651–661, August 1979. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http:

//projecteuclid.org/euclid.aop/1176994988.

Silverstein:1974:SMC

[Sil74] Martin L. Silverstein. Symmetric Markov chains. Annals of Probability, 2(4):681–701, August 1974. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.org/

euclid.aop/1176996612.

Simons:1974:GDF

[Sim74] Gordon Simons. Generalized distribution functions: The linearly ordered case with applications to nonparametric statistics. Annals of Probability, 2(3):501–508, June 1974. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176996667.

Simons:1975:GDF

[Sim75] Gordon Simons. Generalized distribution functions: The σ-lower finite case. Annals of Probability, 3(3):492–502, June 1975. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http:

//projecteuclid.org/euclid.aop/1176996355.

Simons:1977:UE

[Sim77] Gordon Simons. An unexpected expectation. Annals of Probability, 5(1):157–158, February 1977. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.org/

euclid.aop/1176995902.

Singh:1975:GCT

[Sin75] Radhey S. Singh. On the Glivenko–Cantelli theorem for weighted empiricals based on independent random variables. Annals of Probability, 3(2):371–374, April 1975. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.org/

euclid.aop/1176996410.

Segall:1975:RND

[SK75] Adrian Segall and Thomas Kailath. Radon–Nikodym derivatives with respect to measures induced by discontinuous independent-increment processes. Annals of Probability, 3(3):449–464, June 1975. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176996352.

Slud:1977:DIB

[Slu77] Eric V. Slud. Distribution inequalities for the binomial law. Annals of Probability, 5(3):404–412, June 1977. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176995801.

Smythe:1973:SLL

[Smy73] R. T. Smythe. Strong laws of large numbers for r-dimensional arrays of random variables. Annals of Probability, 1(1):164–170, February 1973. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176997031.

Smythe:1974:SIR

[Smy74] R. T. Smythe. Sums of independent random variables on partially ordered sets. Annals of Probability, 2(5):906–917, October 1974. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176996556.

Smythe:1975:RDG

[Smy75] R. T. Smythe. Review: D. G. Kendall, E. F. Harding, Stochastic Analysis. Annals of Probability, 3(6):1062–1067, December 1975. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176996236.

Smythe:1976:MSP

[Smy76] R. T. Smythe. Multiparameter subadditive processes. Annals of Probability, 4(5):772–782, October 1976. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176995983.

Shields:1978:DRF

[SNDL78] P. C. Shields, D. L. Neuhoff, L. D. Davisson, and F. Ledrappier. The distortion-rate function for nonergodic sources. Annals of Probability, 6(1):138–143, February 1978. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.org/

euclid.aop/1176995618.

Solomon:1975:RWR

[Sol75] Fred Solomon. Random walks in a random environment. Annals of Probability, 3(1):1–31, February 1975. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176996444.

Spitzer:1975:MRF

[Spi75] Frank Spitzer. Markov random fields on an infinite tree. Annals of Probability, 3(3):387–398, June 1975. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176996347.

Simons:1978:WIP

[SS78] Gordon Simons and William Stout. A weak invariance principle with applications to domains of attraction. Annals of Probability, 6(2): 294–315, April 1978. CODEN APBYAE. ISSN 0091-1798 (print),

2168-894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176995574.

Shields:1975:EZB

[ST75] Paul Shields and J.-P. Thouvenot. Entropy zero × Bernoulli processes are closed in the d̄-metric. Annals of Probability, 3(4):732–736, August 1975. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.aop/
1176996314.

Stepan:1973:PLI

[Ste73] Josef Stepan. The probability limit identification function exists under the continuum hypothesis. Annals of Probability, 1(4):712–715, August 1973. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.aop/
1176996899.

Steck:1974:NF

[Ste74] G. P. Steck. A new formula for P (R_{i} ≤ b_{i}, 1 ≤ i ≤ m | m, n, F = G^{k}). Annals of Probability, 2(1):155–160, February 1974. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http:

//projecteuclid.org/euclid.aop/1176996761. See [Ste75].

Steck:1975:NAP

[Ste75] G. P. Steck. Note: Acknowledgment of priority to “A New Formula for P (R_{i} ≤ b_{i}, 1 ≤ i ≤ m | m, n, F = G^{k})”. Annals of Probability, 3(4):739, August 1975. CODEN APBYAE. ISSN 0091-1798 (print),

2168-894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176996316. See [Ste74].

Steele:1977:FSM

[Ste77a] J. Michael Steele. Families of sample means converge slowly. Annals of Probability, 5(6):1036–1038, December 1977. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http:

//projecteuclid.org/euclid.aop/1176995672.

Steele:1977:LPR

[Ste77b] J. Michael Steele. Limit properties of random variables associated with a partial ordering of R^{d}. Annals of Probability, 5(3):395–403, June 1977. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.aop/
1176995800.

Stern:1977:IBM

[Ste77c] Frederick Stern. An independence in Brownian motion with constant drift. Annals of Probability, 5(4):571–572, August 1977. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http:

//projecteuclid.org/euclid.aop/1176995763.

Steele:1978:EDS

[Ste78a] J. Michael Steele. Empirical discrepancies and subadditive processes. Annals of Probability, 6(1):118–127, February 1978. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http:

//projecteuclid.org/euclid.aop/1176995615.

Steinebach:1978:CRL

[Ste78b] Josef Steinebach. Convergence rates of large deviation probabilities in the multidimensional case. Annals of Probability, 6(5):751–759, October 1978. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.aop/
1176995426.

Steck:1979:LBM

[Ste79] George P. Steck. Lower bounds for the multivariate normal Mills’ ratio. Annals of Probability, 7(3):547–551, June 1979. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http:

//projecteuclid.org/euclid.aop/1176995056.

Stout:1973:MIL

[Sto73] William F. Stout. Maximal inequalities and the law of the iterated logarithm. Annals of Probability, 1(2):322–328, April 1973. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http:

//projecteuclid.org/euclid.aop/1176996985.

Stute:1977:CRI

[Stu77] Winfried Stute. Convergence rates for the isotrope discrepancy. Annals of Probability, 5(5):707–723, October 1977. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http:

//projecteuclid.org/euclid.aop/1176995714.

Sussmann:1978:GBD

[Sus78] Hector J. Sussmann. On the gap between deterministic and stochastic ordinary differential equations. Annals of Probability, 6(1):19–41, February 1978. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.aop/
1176995608.

Steutel:1979:DAS

[SvH79] F. W. Steutel and K. van Harn. Discrete analogues of selfdecomposability and stability. Annals of Probability, 7(5):893–899, October 1979. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.aop/
1176994950.

Szulga:1976:CSB

[SW76] Jerzy Szulga and Wojbor A. Woyczynski. Convergence of submartingales in Banach lattices. Annals of Probability, 4(3):464– 469, June 1976. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.aop/
1176996095.

Shorack:1978:LBE

[SW78] Galen R. Shorack and Jon A. Wellner. Linear bounds on the empirical distribution function. Annals of Probability, 6(2):349–353, April 1978. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176995582.

Sweeting:1977:SCM

[Swe77] T. J. Sweeting. Speeds of convergence for the multidimensional Central Limit Theorem. Annals of Probability, 5(1):28–41, February 1977. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176995888.

Szasz:1977:PTL

[Sza77a] Domokos Szasz. A problem of two lifts. Annals of Probability, 5(4): 550–559, August 1977. CODEN APBYAE. ISSN 0091-1798 (print),

2168-894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176995760.

Szasz:1977:USD

[Sza77b] Domokos Szasz. Uniformity in Stone’s decomposition of the renewal measure. Annals of Probability, 5(4):560–564, August 1977. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176995761.

Takacs:1974:STP

[Tak74] Lajos Takacs. Sojourn time problems. Annals of Probability, 2(3): 420–431, June 1974. CODEN APBYAE. ISSN 0091-1798 (print),

2168-894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176996657.

Tanny:1977:LTB

[Tan77] David Tanny. Limit theorems for branching processes in a random environment. Annals of Probability, 5(1):100–116, February 1977. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176995894.

Taylor:1975:SBM

[Tay75a] Howard M. Taylor. A stopped Brownian motion formula. Annals of Probability, 3(2):234–246, April 1975. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176996395.

Taylor:1975:CKS

[Tay75b] J. C. Taylor. A characterization of the kernel lim_{λ} ↓ 0V_{λ} for sub-Markovian resolvents (V_{λ})^{1}. Annals of Probability, 3(2):355–357, April 1975. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.aop/
1176996407.

Teicher:1973:CLT

[Tei73] Henry Teicher. A classical limit theorem without invariance or reflection. Annals of Probability, 1(4):702–704, August 1973. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http:

//projecteuclid.org/euclid.aop/1176996897.

Teicher:1974:LIL

[Tei74] Henry Teicher. On the law of the iterated logarithm. Annals of Probability, 2(4):714–728, August 1974. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.org/

euclid.aop/1176996614.

Teicher:1979:RGR

[Tei79] Henry Teicher. Rapidly growing random walks and an associated stopping time. Annals of Probability, 7(6):1078–1081, December 1979. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176994903.

Teugels:1975:CSD

[Teu75] Jozef L. Teugels. The class of subexponential distributions. Annals of Probability, 3(6):1000–1011, December 1975. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http:

//projecteuclid.org/euclid.aop/1176996225.

Theodorescu:1974:RMF

[The74] Radu Theodorescu. Review: M. Frank Norman, Markov Processes and Learning Models. Annals of Probability, 2(5):964–966, October 1974. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176996563.

Thorburn:1977:LLT

[Tho77] Daniel Thorburn. A local limit theorem for the Wilcoxon rank sum. Annals of Probability, 5(6):926–939, December 1977. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http:

//projecteuclid.org/euclid.aop/1176995661.

Tomkins:1975:ILR

[Tom75a] R. J. Tomkins. Iterated logarithm results for weighted averages of martingale difference sequences. Annals of Probability, 3(2):307– 314, April 1975. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.aop/
1176996401.

Tomkins:1975:CM

[Tom75b] R. J. Tomkins. On conditional medians. Annals of Probability, 3(2): 375–379, April 1975. CODEN APBYAE. ISSN 0091-1798 (print),

2168-894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176996411.

Tomkins:1976:SLT

[Tom76] R. J. Tomkins. Strong limit theorems for certain arrays of random variables. Annals of Probability, 4(3):444–452, June 1976. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176996092.

Tomkins:1978:LIL

[Tom78] R. J. Tomkins. On the law of the iterated logarithm. Annals of Probability, 6(1):162–168, February 1978. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176995622.

Torrez:1978:BDC

[Tor78] William C. Torrez. The birth and death chain in a random environment: Instability and extinction theorems. Annals of Probability, 6 (6):1026–1043, December 1978. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.org/

euclid.aop/1176995391.

Tran:1977:HDR

[Tra77] Lanh Tat Tran. The Hausdorff dimension of the range of the N-parameter Wiener process. Annals of Probability, 5(2):235–242, April 1977. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.aop/
1176995848.

Tran:1979:RLP

[Tra79] Lanh Tat Tran. The range of Lévy’s N-parameter Brownian motion in d-space. Annals of Probability, 7(3):532–536, June 1979. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http:

//projecteuclid.org/euclid.aop/1176995053.

Turnbull:1973:IBP

[Tur73a] Bruce W. Turnbull. Inequalities for branching processes. Annals of Probability, 1(3):457–474, June 1973. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176996939.

Turnbull:1973:IMB

[Tur73b] Bruce W. Turnbull. Inequalities for multitype branching processes. Annals of Probability, 1(3):475–479, June 1973. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176996940.

Taylor:1979:LLN

[TW79] R. L. Taylor and Duan Wei. Laws of large numbers for tight random elements in normed linear spaces. Annals of Probability, 7(1):150– 155, February 1979. CODEN APBYAE. ISSN 0091-1798 (print),

2168-894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176995156.

Tweedie:1974:TMCa

[Twe74a] Richard L. Tweedie. R-theory for Markov chains on a general state space i: Solidarity properties and R-recurrent chains. Annals of Probability, 2(5):840–864, October 1974. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176996552.

Tweedie:1974:TMCb

[Twe74b] Richard L. Tweedie. R-theory for Markov chains on a general state space II: r-subinvariant measures for r-transient chains. Annals of Probability, 2(5):865–878, October 1974. CODEN APBYAE.

ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176996553.

VanZuijlen:1978:PED

[Van78a] Martien C. A. Van Zuijlen. Properties of the empirical distribution function for independent nonidentically distributed random variables. Annals of Probability, 6(2):250–266, April 1978. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http:

//projecteuclid.org/euclid.aop/1176995571.

VanZwet:1978:PKC

[Van78b] W. R. Van Zwet. A proof of Kakutani’s conjecture on random subdivision of longest intervals. Annals of Probability, 6(1):133–137, February 1978. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.aop/
1176995617.

Veneziano:1979:EVR

[Ven79] Daniele Veneziano. Envelopes of vector random processes and their crossing rates. Annals of Probability, 7(1):62–74, February 1979. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176995148.

Vervaat:1979:RBB

[Ver79] Wim Vervaat. A relation between Brownian bridge and Brownian excursion. Annals of Probability, 7(1):143–149, February 1979. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176995155.

Villalobos:1974:SFL

[Vil74] Alvaro Gonzalez Villalobos. Some functional limit theorems for dependent random variables. Annals of Probability, 2(6):1090–1107, December 1974. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.aop/
1176996500.

Vitale:1976:CRF

[VP76] R. A. Vitale and A. C. Pipkin. Conditions on the regression function when both variables are uniformly distributed. Annals of Probability, 4(5):869–873, October 1976. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.org/

euclid.aop/1176995992.

VanSchuppen:1974:TLM

[VW74] Jan H. Van Schuppen and Eugene Wong. Transformation of local martingales under a change of law. Annals of Probability, 2(5):879– 888, October 1974. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.aop/
1176996554.

Wachter:1978:SLR

[Wac78] Kenneth W. Wachter. The strong limits of random matrix spectra for sample matrices of independent elements. Annals of Probability, 6 (1):1–18, February 1978. CODEN APBYAE. ISSN 0091-1798 (print),

2168-894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176995607.

Walker:1974:OSVb

[Wal74a] LeRoy H. Walker. Optimal stopping variables for Brownian motion. Annals of Probability, 2(2):317–320, April 1974. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http:

//projecteuclid.org/euclid.aop/1176996711.

Walker:1974:OSVa

[Wal74b] LeRoy H. Walker. Optimal stopping variables for stochastic process with independent increments. Annals of Probability, 2(2):309– 316, April 1974. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.aop/
1176996710.

Wang:1975:CSS

[Wan75] Y. H. Wang. Characterizations of some stochastic processes. Annals of Probability, 3(6):1038–1045, December 1975. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http:

//projecteuclid.org/euclid.aop/1176996231.

Wang:1977:QVF

[Wan77] Albert T. Wang. Quadratic variation of functionals of Brownian motion. Annals of Probability, 5(5):756–769, October 1977. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http:

//projecteuclid.org/euclid.aop/1176995717.

Wang:1979:TDF

[Wan79] Albert T. Wang. Time-dependent functions of Brownian motion that are Markovian. Annals of Probability, 7(3):515–525, June 1979. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176995051.

Weissman:1975:EPG

[Wei75a] Ishay Weissman. Extremal processes generated by independent nonidentically distributed random variables. Annals of Probability, 3(1): 172–177, February 1975. CODEN APBYAE. ISSN 0091-1798 (print),

2168-894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176996459.

Weissman:1975:LSF

[Wei75b] Ishay Weissman. On location and scale functions of a class of limiting processes with application to extreme value theory. Annals of Probability, 3(1):178–181, February 1975. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176996460.

Weissman:1976:WCE

[Wei76] Ishay Weissman. On weak convergence of extremal processes. Annals of Probability, 4(3):470–473, June 1976. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176996096.

Weiner:1977:NCB

[Wei77] H. Weiner. Note: Correction to “An Integral Equation in Age-Dependent Branching Processes”. Annals of Probability, 5(1):159, February 1977. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.aop/
1176995903.

Welsch:1973:CTE

[Wel73] Roy E. Welsch. A convergence theorem for extreme values from Gaussian sequences. Annals of Probability, 1(3):398–404, June 1973. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176996934.

Wellner:1977:MIE

[Wel77] Jon A. Wellner. A martingale inequality for the empirical process. Annals of Probability, 5(2):303–308, April 1977. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http:

//projecteuclid.org/euclid.aop/1176995856.

Wellner:1978:SIT

[Wel78] Jon A. Wellner. A strong invariance theorem for the Strong Law of Large Numbers. Annals of Probability, 6(4):673–679, August 1978. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176995488.

Wendel:1977:NPB

[Wen77] J. G. Wendel. A note on planar Brownian motion. Annals of Probability, 5(5):802–803, October 1977. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.org/

euclid.aop/1176995722.

Westcott:1976:SPR

[Wes76] M. Westcott. Simple proof of a result on thinned point processes. Annals of Probability, 4(1):89–90, February 1976. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http:

//projecteuclid.org/euclid.aop/1176996183.

Westcott:1979:TBR

[Wes79] Mark Westcott. On the tail behaviour of record-time distributions in a random record process. Annals of Probability, 7(5):868–873, October 1979. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.aop/
1176994946.

Wichura:1973:NCS

[Wic73a] M. J. Wichura. A note on the convergence of series of stochastic processes. Annals of Probability, 1(1):180–182, February 1973. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176997034.

Wichura:1973:BCP

[Wic73b] Michael J. Wichura. Boundary crossing probabilities associated with Motoo’s law of the iterated logarithm. Annals of Probability, 1(3): 437–456, June 1973. CODEN APBYAE. ISSN 0091-1798 (print),

2168-894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176996938.

Wichura:1973:SST

[Wic73c] Michael J. Wichura. Some Strassen-type laws of the iterated logarithm for multiparameter stochastic processes with independent increments. Annals of Probability, 1(2):272–296, April 1973. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176996980.

Wichura:1974:FLI

[Wic74a] Michael J. Wichura. Functional laws of the iterated logarithm for the partial sums of i. i. d. random variables in the domain of attraction of a completely asymmetric stable law. Annals of Probability, 2(6):1108– 1138, December 1974. CODEN APBYAE. ISSN 0091-1798 (print),

2168-894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176996501.

Wichura:1974:FFL

[Wic74b] Michael J. Wichura. On the functional form of the law of the iterated logarithm for the partial maxima of independent identically distributed random variables. Annals of Probability, 2(2):202–230, April 1974. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.aop/
1176996704.

Williamson:1974:RKB

[Wil74] John A. Williamson. Review: K. B. Athreya, P. E. Ney, Branching Processes. Annals of Probability, 2(5):966–968, October 1974. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176996564.

Winkler:1973:NCP

[Win73] William Winkler. A note on continuous parameter zero-two law. Annals of Probability, 1(2):341–344, April 1973. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176996989.

Winkler:1978:CPM

[Win78] W. Winkler. Continuous parameter Markov processes. Annals of Probability, 6(3):459–468, June 1978. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176995530.

Wallenstein:1973:PNN

[WN73] Sylvan R. Wallenstein and Joseph I. Naus. Probabilities for a k th nearest neighbor problem on the line. Annals of Probability, 1(1): 188–190, February 1973. CODEN APBYAE. ISSN 0091-1798 (print),

2168-894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176997037.

Wolfe:1973:LBC

[Wol73] Stephen J. Wolfe. On the local behavior of characteristic functions. Annals of Probability, 1(5):862–866, October 1973. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http:

//projecteuclid.org/euclid.aop/1176996853.

Wolfe:1975:DCF

[Wol75] Stephen J. Wolfe. On derivatives of characteristic functions. Annals of Probability, 3(4):737–738, August 1975. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176996315.

Wolfe:1978:BCF

[Wol78] Stephen J. Wolfe. On the behavior of characteristic functions on the real line. Annals of Probability, 6(4):554–562, August 1978. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176995477.

Wong:1979:CLT

[Won79] Sherman Wong. A Central Limit Theorem for piecewise monotonic mappings of the unit interval. Annals of Probability, 7(3):500–514, June 1979. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.aop/
1176995050.

Woodroofe:1976:RTC

[Woo76] Michael Woodroofe. A renewal theorem for curved boundaries and moments of first passage times. Annals of Probability, 4(1):67–80, February 1976. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.aop/
1176996181.

Wright:1977:SLW

[WPR77] F. T. Wright, Ronald D. Platt, and Tim Robertson. A Strong Law for weighted averages of independent, identically distributed random variables with arbitrarily heavy tails. Annals of Probability, 5(4): 586–590, August 1977. CODEN APBYAE. ISSN 0091-1798 (print),

2168-894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176995767.

Wierman:1978:CFP

[WR78] John C. Wierman and Wolfgang Reh. On conjectures in first passage percolation theory. Annals of Probability, 6(3):388–397, June 1978.

CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176995525.

Wright:1973:BTP

[Wri73] F. T. Wright. A bound on tail probabilities for quadratic forms in independent random variables whose distributions are not necessarily symmetric. Annals of Probability, 1(6):1068–1070, December 1973. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176996815.

Wright:1976:ETS

[Wri76] A. Larry Wright. An ergodic theorem for the square of a wide-sense stationary process. Annals of Probability, 4(5):829–835, October 1976. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176995986.

Wright:1979:SLV

[Wri79] F. T. Wright. A strong law for variables indexed by a partially ordered set with applications to isotone regression. Annals of Probability, 7(1): 109–127, February 1979. CODEN APBYAE. ISSN 0091-1798 (print),

2168-894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176995152.

Wilmesmeier:1979:ASS

[WW79] James M. Wilmesmeier and F. T. Wright. The almost sure stability of quadratic forms. Annals of Probability, 7(4):738–743, August 1979. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176994995.

Wong:1976:WMS

[WZ76] Eugene Wong and Moshe Zakai. Weak martingales and stochastic integrals in the plane. Annals of Probability, 4(4):570–586, August 1976. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176996028.

Wong:1977:ESI

[WZ77a] Eugene Wong and Moshe Zakai. An extension of stochastic integrals in the plane. Annals of Probability, 5(5):770–778, October 1977. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176995718.

Wong:1977:SFC

[WZ77b] Eugene Wong and Moshe Zakai. The sample function continuity of stochastic integrals in the plane. Annals of Probability, 5(6):1024– 1027, December 1977. CODEN APBYAE. ISSN 0091-1798 (print),

2168-894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176995670.

Yamazato:1978:UID

[Yam78] Makoto Yamazato. Unimodality of infinitely divisible distribution functions of class L. Annals of Probability, 6(4):523–531, August 1978. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176995474.

Yoder:1975:HDG

[Yod75] Lane Yoder. The Hausdorff dimensions of the graph and range of Nparameter Brownian motion in d-space. Annals of Probability, 3(1): 169–171, February 1975. CODEN APBYAE. ISSN 0091-1798 (print),

2168-894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176996458.

Zabell:1979:CVR

[Zab79] Sandy Zabell. Continous versions of regular conditional distributions. Annals of Probability, 7(1):159–165, February 1979. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http:

//projecteuclid.org/euclid.aop/1176995158.

Zinn:1977:NCL

[Zin77] Joel Zinn. A note on the Central Limit Theorem in Banach spaces. Annals of Probability, 5(2):283–286, April 1977. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176995852.